### v1

# Bias–variance Tradeoff in Tensor Estimation

Abstract
We study denoising of a third-order tensor when the ground-truth tensor is not necessarily Tucker low-rank. Specifically, we observe
$Y=X^{\ast}+Z\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}},$
where $X^{\ast}$ is the ground-truth tensor, and $Z$ is the noise tensor. We propose a simple variant of the higher-order tensor SVD estimator $\widetilde{X}$. We show that uniformly over all user-specified Tucker ranks $(r_{1},r_{2},r_{3})$,
$\|\widetilde{X}-X^{*}\|_{\mathrm{F}}^{2}=O\Big(\kappa^{2}\Big\{r_{1}r_{2}r_{3}+\sum_{k=1}^{3}p_{k}r_{k}\Big\}\;+\;\xi_{(r_{1},r_{2},r_{3})}^{2}\Big)\quad\text{ with high probability.}$
Here, the bias term $\xi_{(r_{1},r_{2},r_{3})}$ corresponds to the best achievable approximation error of $X^{\ast}$ over the class of tensors with Tucker ranks $(r_{1},r_{2},r_{3})$; $\kappa^{2}$ quantifies the noise level; and the variance term $\kappa^{2}\{r_{1}r_{2}r_{3}+\sum_{k=1}^{3}p_{k}r_{k}\}$ scales with the effective number of free parameters in the estimator $\widetilde{X}$. Our analysis achieves a clean rank-adaptive bias–variance tradeoff: as we increase the ranks of estimator $\widetilde{X}$, the bias $\xi(r_{1},r_{2},r_{3})$ decreases and the variance increases. As a byproduct we also obtain a convenient bias-variance decomposition for the vanilla low-rank SVD matrix estimators.
1 Introduction
Low-rank tensor models are a cornerstone of modern machine learning. They capture essential structure in high-dimensional signals while offering substantial gains in computation and storage. Applications include recommender systems (Koren et al.,, 2009), topic modeling (Blei et al.,, 2003), community detection (Anandkumar et al.,, 2014; Abbe,, 2018), and, more recently, latent variable learning (Diakonikolas and Kane,, 2024) and generative modeling (Hur et al.,, 2023; Peng et al.,, 2023). The effectiveness of these methods demonstrate that many real-world signals embedded in high-dimensional ambient spaces can be well-approximated by low-rank structures.
The study of exactly low-rank estimation is well-developed, particularly in the matrix setting. A large body of work has introduced efficient algorithms for tasks such as matrix completion and denoising, including nuclear norm relaxation (Koltchinskii et al.,, 2011), convex optimization (Candes and Recht,, 2012), nonconvex methods (Chi et al.,, 2019), and SVD guarantees with unbalanced matrices (Cai and Zhang,, 2018).
Parallel efforts in low-rank tensor estimation are well-developed, including the study of higher-order tensor singular value decompositions in the non-random setting (De Lathauwer et al., 2000a, ; De Lathauwer et al., 2000b, ) and Riemannian optimization methods for low-rank Tucker tensor completion (Kressner et al.,, 2014). Nevertheless, the tensor setting poses additional challenges: many fundamental problems are computationally intractable, multiple inequivalent notions of rank (e.g. CP and Tucker) exist, and best low-rank approximations need not be unique (e.g. Hillar and Lim,, 2013). Most theoretical guarantees in tensor estimation (e.g. Zhang and Xia,, 2018; Han et al.,, 2022), like their matrix counterparts, rely on the assumption that the underlying signal is exactly low-rank.
In practice, however, signals are not exactly low-rank: their singular spectra typically decay gradually rather than vanishing at a finite rank. This creates an interesting gap between theory and practice, since existing analyses provide guarantees under idealized exact low-rankness and leave open the question of how to analyze estimation error in more realistic settings.
We address this gap by designing and analyzing low-rank tensor estimation when the signal is not low-rank. In such settings, estimation error is governed by a bias–variance tradeoff: truncating to a smaller rank incurs approximation bias, while larger ranks inflate variance by amplifying noise. In particular, we propose a simple variant of the higher-order singular value decomposition (HOSVD) algorithm and provide the analysis of tensor estimation without assuming low-rankness. Our main result (Theorem 1) establishes an explicit bias–variance decomposition of the estimation error:
  • The bias term quantifies the approximation error incurred by truncating a spectrum.
  • The variance term matches the optimal rate for exactly low-rank models.
The success of SVD and HOSVD has been well established by their decades of widespread use. The objective of our manuscript is to provide mathematical justification for the information-efficient bias-variance tradeoff achieved by these two classical algorithms, and to establish uniform performance guarantees that hold for any user-specified ranks.
To achieve this clean bias-variance tradeoff, we utilize classical linear algebra results such as Mirsky’s and Ky Fan’s theorems to develop non-trivial extensions of classical perturbation tools to general tensors. As a convenient side result, we also recover a clean bias–variance characterization for matrices, thereby unifying the matrix and tensor perspectives. Finally, we validate our theory with simulations that confirm the existence of tradeoffs across different regimes of spectral decay and noise.
1.1 Notation
Matrices: For positive integers $p,r$, let $\mathbb{O}^{p\times r}=\{V\in\mathbb{R}^{p\times r}:V^{\top}V=I_{r}\}$, the set of column-orthonormal matrices. Let $M\in\mathbb{R}^{p\times q}$ and suppose the full singular value decomposition (SVD) satisfies $M=U\Sigma V^{\top}$. Here $s=\operatorname*{rank}(M)$, $U\in\mathbb{O}^{p\times s}$, $V\in\mathbb{O}^{q\times s}$, and $\Sigma\in\mathbb{R}^{s\times s}$ is diagonal with singular values $\sigma_{1}(M)\geq\sigma_{2}(M)\geq\cdots\geq\sigma_{s}(M)\geq 0.$ We write the smallest singular value of $M$ as $\sigma_{\min}(M)$. The operator norm and Frobenius norm of $M$ are defined as
$\|M\|=\sigma_{1}(M),\qquad\|M\|_{\mathrm{F}}=\Big(\sum_{i=1}^{p}\sum_{j=1}^{q}M_{i,j}^{2}\Big)^{1/2}.$
For $r\leq\operatorname*{rank}(M)$, the rank-$r$ truncated SVD is $M_{(r)}=U_{(r)}\Sigma_{(r)}V_{(r)}^{\top},$ where $U_{(r)}\in\mathbb{O}^{p\times r}$ and $V_{(r)}\in\mathbb{O}^{q\times r}$ corresponds to the leading $r$ left and right singular vectors respectively, and $\Sigma_{(r)}=\operatorname*{diag}\{\sigma_{1}(M),\ldots,\sigma_{r}(M)\}$. For brevity, we use $\text{SVD}_{r}(M)$ to denote the leading $r$ left singular vectors, i.e.
$\text{SVD}_{r}(M)=U_{(r)}.$
For two matrices $M_{1}\in\mathbb{R}^{p_{1}\times q_{1}}$ and $M_{2}\in\mathbb{R}^{p_{2}\times q_{2}}$, their Kronecker product is $M_{1}\otimes M_{2}\in\mathbb{R}^{(p_{1}p_{2})\times(q_{1}q_{2})}$.
Let $U,\widehat{U}\in\mathbb{O}^{p\times r}$ be two singular subspaces. We write the principal angles between $U$ and $\widehat{U}$ as
[…]
Tensors: For any tensor $B\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}}$, the Frobenius norm is
$\|B\|_{\mathrm{F}}=\left(\sum_{i_{1}=1}^{p_{1}}\sum_{i_{2}=1}^{p_{2}}\sum_{i_{3}=1}^{p_{3}}B_{i_{1},i_{2},i_{3}}^{2}\right)^{1/2}.$
The mode-$1$ matricization $\mathcal{M}_{1}(B)$ is the unfolding of $B$ into a $p_{1}\times p_{2}p_{3}$ matrix. The mode-$2$ and mode-$3$ matricization of $B$ are defined similarly. The mode-$1$ product of $B$ with a matrix $M\in\mathbb{R}^{q\times p_{1}}$ is defined as
$(B\times_{1}M)(j,i_{2},i_{3})=\sum_{i_{1}=1}^{p_{1}}M(i_{1},j)B(i_{1},i_{2},i_{3}).$
We write the Tucker rank of $B$ as $(r_{1},r_{2},r_{3})$ if
$\operatorname*{rank}(\mathcal{M}_{j}(B))=r_{j},\text{ for }j=1,2,3.$
Random variables: For a random variable $X\in\mathbb{R}$, we denote the sub-Gaussian and sub-Exponential norms as
$\|X\|_{\psi_{2}}=\inf\{K>0:\mathbb{E}\exp(X^{2}/K^{2})\leq 2\}\quad\text{and}\quad\|X\|_{\psi_{1}}=\inf\{K>0:\mathbb{E}\exp(|X|/K)\leq 2\}.$
We write $X\sim\mathrm{subGaussian}(0,\kappa^{2})$ if $\mathbb{E}(X)=0$ and $\|X\|_{\psi_{2}}\leq\kappa$.
Universal constants: We use $C_{1},C_{2},\dots$ and $c_{1},c_{2},\dots$ to denote positive constants whose values may differ from place to place.
2 Bias–variance tradeoff in tensor estimation
We consider the noisy tensor model
$Y=X^{*}+Z\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}},$ (1)
where $X^{*}$ is an unknown signal and $Z$ is a random perturbation. Even when $X^{*}$ has a full Tucker rank $(p_{1},p_{2},p_{3})$, it is common to approximate $X^{*}$ by a Tucker rank-$(r_{1},r_{2},r_{3})$ estimator. This is because (i) low-rank structure yields dramatic gains in computation and memory, see Remark 2; and (ii) the approximation bias can decay quickly when approximating $X^{*}$ by a Tucker low-rank tensor, so balancing bias and variance may lead to a smaller estimation error. Given a target Tucker rank $(r_{1},r_{2},r_{3})$, we consider the following simple variant of the HOSVD algorithm (De Lathauwer et al., 2000b, ).
Algorithm 1 One-step HOSVD
Tensor $Y$; target Tucker rank $(r_{1},r_{2},r_{3})$.
for $k=1,2,3$ do
  $U_{k}^{(0)}\leftarrow\text{SVD}_{r_{k}}\!\left(\mathcal{M}_{k}(Y)\right)\in\mathbb{O}^{p_{k}\times r_{k}}$.
end for
$U_{1}^{(1)}\leftarrow\text{SVD}_{r_{1}}\!\Big(\mathcal{M}_{1}(Y)\cdot\{U_{2}^{(0)}\otimes U_{3}^{(0)}\}\Big)\in\mathbb{O}^{p_{1}\times r_{1}}$,
$U_{2}^{(1)}\leftarrow\text{SVD}_{r_{2}}\!\Big(\mathcal{M}_{2}(Y)\cdot\{U_{1}^{(0)}\otimes U_{3}^{(0)}\}\Big)\in\mathbb{O}^{p_{2}\times r_{2}}$,
$U_{3}^{(1)}\leftarrow\text{SVD}_{r_{3}}\!\Big(\mathcal{M}_{3}(Y)\cdot\{U_{1}^{(0)}\otimes U_{2}^{(0)}\}\Big)\in\mathbb{O}^{p_{3}\times r_{3}}$.
$\widetilde{X}\leftarrow Y\times_{1}U_{1}^{(1)}U_{1}^{(1)\top}\times_{2}U_{2}^{(1)}U_{2}^{(1)\top}\times_{3}U_{3}^{(1)}U_{3}^{(1)\top}$.
Let $\mathcal{T}_{(r_{1},r_{2},r_{3})}$ denote the class of tensors in $\mathbb{R}^{p_{1}\times p_{2}\times p_{3}}$ with Tucker rank at most $(r_{1},r_{2},r_{3})$, i.e.
$\mathcal{T}_{(r_{1},r_{2},r_{3})}=\big\{A\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}}:\ \operatorname*{rank}(\mathcal{M}_{k}(A))\leq r_{k},\ k=1,2,3\big\}.$
For any tensor $X^{*}\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}}$, let $\xi_{(r_{1},r_{2},r_{3})}$ denote the best Tucker rank-$(r_{1},r_{2},r_{3})$ approximation error:
$\xi_{(r_{1},r_{2},r_{3})}\;=\;\inf_{A\in\mathcal{T}_{(r_{1},r_{2},r_{3})}}\|A-X^{*}\|_{\mathrm{F}}.$
Theorem 1.
Suppose $Y$ follows (1) with $Z_{\mu_{1},\mu_{2},\mu_{3}}\overset{\mathrm{i.i.d.}}{\sim}\mathrm{subGaussian}(0,\kappa^{2})$. Let $\widetilde{X}$ be the output of Algorithm 1 with target Tucker rank $(r_{1},r_{2},r_{3})$. Define $r_{\max}=\max_{k}r_{k}$ and $p_{\min}=\min_{k}p_{k}$. Assume the singular gaps satisfy
$\big\{\sigma_{r_{k}}(\mathcal{M}_{k}(X^{*}))-\sigma_{r_{k}+1}(\mathcal{M}_{k}(X^{*}))\big\}^{2}\;\geq\;C_{\mathrm{gap}}\kappa^{2}\Big(\sqrt{p_{1}p_{2}p_{3}\,r_{\max}}+\sum_{k=1}^{3}p_{k}r_{\max}\Big),$ (2)
for all $k=1,2,3$ and for a sufficiently large constant $C_{\mathrm{gap}}>0$. Then, with probability at least $1-C_{1}p_{1}p_{2}p_{3}\exp(-C_{2}p_{\min})$,
$\|\widetilde{X}-X^{*}\|_{\mathrm{F}}\;\leq\;C_{3}\Big\{\sqrt{\kappa^{2}\big(\textstyle\sum_{k=1}^{3}p_{k}r_{k}+r_{1}r_{2}r_{3}\big)}+\xi_{(r_{1},r_{2},r_{3})}\Big\}.$ (3)
Adaptivity in Theorem 1. For any user specified target Tucker rank $(r_{1},r_{2},r_{3})$, the estimator $\widetilde{X}$ satisfies (3). In other words, the result holds uniformly over all target Tucker ranks.
Bias–variance tradeoff. The error bound (3) exhibits a clear bias–variance decomposition for Tucker rank-$(r_{1},r_{2},r_{3})$ estimators. The bias term $\xi_{(r_{1},r_{2},r_{3})}$ is the best achievable approximation error of $X^{*}$ at the chosen ranks and decreases as the $r_{k}$ increase. Therefore, as the ranks grow, variance increases while bias shrinks. Balancing these two leads to improved accuracy.
Tightness. Our upper bound in (3) attains the minimax optimal rate (up to constants) over the class of tensors with Tucker rank at most $(r_{1},r_{2},r_{3})$. Indeed, by definition, $\xi_{(r_{1},r_{2},r_{3})}$ is the smallest possible approximation error of $X^{*}$ over $\mathcal{T}_{(r_{1},r_{2},r_{3})}$. In addition, Zhang and Xia, (2018) prove that, when $X^{*}$ is exactly low-rank with Tucker rank $(r_{1},r_{2},r_{3})$, the term $\kappa^{2}\Bigl(\sum_{k=1}^{3}p_{k}r_{k}+r_{1}r_{2}r_{3}\Bigr)$ is minimax optimal up to a universal constant. For arbitrary (e.g. full-rank) tensors, our bound should be interpreted as an upper bound; a matching minimax lower bound for that broader class is not known.
Proof techniques. For the bias, we leverage classical linear algebra tools, such as Mirsky’s and Ky Fan’s theorems, to characterize the optimal approximation error appearing in (3). To handle the variance of the estimator $\widetilde{X}$, we adopt existing techniques from Zhang and Xia, (2018).
Remark 1 (Assumptions in Theorem 1).
In Theorem 1, we assume that the entries of the noise tensor $Z$ is i.i.d. sub-Gaussian. This is a commonly seen condition in the tensor literature (e.g. Zhang and Xia,, 2018; Han et al.,, 2022).
The signal-to-noise ratio (SNR) condition (2) in Theorem 1 is also commonly seen in literature. In fact, if the ground truth $X^{*}$ is exactly Tucker low-rank with Tucker rank $(r_{1},r_{2},r_{3})$, then $\sigma_{r_{k}+1}(\mathcal{M}_{k}(X^{*}))=0$ and (2) reduces to the SNR condition used in Zhang and Xia, (2018), when $r_{\max}$ is a bounded constant.
Remark 2.
Suppose $Y\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}}$ and let
$Y_{(r_{1},r_{2},r_{3})}=S\times_{1}U_{1}\times_{2}U_{2}\times_{3}U_{3},$
be a Tucker rank-$(r_{1},r_{2},r_{3})$ approximation of $Y$, where $U_{k}\in\mathbb{R}^{p_{k}\times r_{k}}$ satisfying $U_{k}^{\top}U_{k}=I_{r_{k}}$ for $k=1,2,3$, and $S\in\mathbb{R}^{r_{1}\times r_{2}\times r_{3}}$ is the core tensor.
For dense multiplications with $v_{k}\in\mathbb{R}^{p_{k}}$ along each mode $k$, computing $Y\times_{1}v_{1}\times_{2}v_{2}\times_{3}v_{3}$ costs $O(p_{1}p_{2}p_{3})$ operations. In contrast, using the Tucker decomposition, we compute $Y_{(r_{1},r_{2},r_{3})}\times_{1}v_{1}\times_{2}v_{2}\times_{3}v_{3}$ successively as
$t_{k}=U_{k}^{\top}v_{k}\in\mathbb{R}^{r_{k}}\ (\text{cost }\approx 2\sum_{k=1}^{3}p_{k}r_{k}),\quad S^{\prime}=S\times_{1}t_{1}\in\mathbb{R}^{r_{2}\times r_{3}}\ (\text{cost }\approx 2r_{1}r_{2}r_{3}),$
$s=t_{2}^{\top}S^{\prime}\in\mathbb{R}^{r_{3}}\quad(\text{cost }\approx 2r_{2}r_{3}),\quad w=s^{\top}t_{3}\in\mathbb{R}\quad(\text{cost }\approx r_{3}).$
Thus, vector multiplication with Tucker factors reduce the computational cost from $O(p_{1}p_{2}p_{3})$ to $O(\sum_{k=1}^{3}p_{k}r_{k}+r_{1}r_{2}r_{3})$. The storage requirements of Tucker decomposition also drop significantly. Storing the full tensor $Y$ requires $O(p_{1}p_{2}p_{3})$ memory, while the Tucker representation requires only $\sum_{k=1}^{3}p_{k}r_{k}+r_{1}r_{2}r_{3}$ scalar storage.
Therefore, both computation and storage of Tucker decomposition scale with $O(\sum_{k=1}^{3}p_{k}r_{k}+r_{1}r_{2}r_{3})$, which is substantially smaller than the dense case when $r_{k}\ll p_{k}$.
3 Bias–variance tradeoff in matrix estimation
We consider the model
$Y=X^{*}+Z\in\mathbb{R}^{m\times n},$ (4)
where $X^{*}$ is an unknown matrix with arbitrary rank, and $Z$ is a perturbation (noise) matrix. Even when $\mathrm{rank}(X^{*})$ is not small, it is common to approximate $X^{*}$ by a rank-$r$ estimator, because (i) low-rank decomposition can dramatically increase the computational and memory efficiency, see Remark 3 for more details; and (ii) the approximation bias can be small, so balancing the bias and variance may lead to smaller estimation error. Let
$X^{*}=\sum_{i=1}^{\min\{m,n\}}\sigma_{i}(X^{*})\,u_{i}^{*}{v_{i}^{*}}^{\!\top},\qquad X_{(r)}^{*}=\sum_{i=1}^{r}\sigma_{i}(X^{*})\,u_{i}^{*}{v_{i}^{*}}^{\!\top}.$
By the Eckart–Young–Mirsky theorem (Eckart and Young,, 1936), $X_{(r)}^{*}$ is the best rank-$r$ approximation to $X^{*}$ in Frobenius norm in the sense that
$\|X^{*}-X^{*}_{(r)}\|_{\mathrm{F}}\leq\|X^{*}-W\|_{\mathrm{F}}\text{ for any }W\in\mathbb{R}^{m\times n}\text{ with }\operatorname*{rank}(W)\leq r.$
Suppose we observe $Y$ instead of $X^{*}$, and write the SVD of $Y$ as
$Y=\sum_{i=1}^{\min\{m,n\}}\sigma_{i}(Y)\,u_{i}v_{i}^{\top},\qquad Y_{(r)}=\sum_{i=1}^{r}\sigma_{i}(Y)\,u_{i}v_{i}^{\top}.$ (5)
That is, $Y_{(r)}$ is the rank-$r$ truncation of the SVD of $Y$. We now provide an upper bound between $Y_{(r)}$ and the unknown matrix $X^{*}$.
Theorem 2.
Under (4), let $Y_{(r)}$ be defined as in (5). Then
$\|Y_{(r)}-X^{*}\|_{\mathrm{F}}\ \leq\ (2+\sqrt{2})(\sqrt{r}\|Z\|+\xi_{(r)}),$ (6)
where $\xi_{(r)}=\|X_{(r)}^{*}-X^{*}\|_{\mathrm{F}}=\sqrt{\sum_{i=r+1}^{\min\{m,n\}}\sigma_{i}^{2}(X^{*})}$.
Proof of Theorem 2.
By the triangle inequality,
$\|Y_{(r)}-X^{*}\|_{\mathrm{F}}\leq\|X_{(r)}^{*}-X^{*}\|_{\mathrm{F}}+\|Y_{(r)}-X_{(r)}^{*}\|_{\mathrm{F}}=\xi_{(r)}+\|Y_{(r)}-X_{(r)}^{*}\|_{\mathrm{F}}.$ (7)
Since $\mathrm{rank}(Y_{(r)}-X_{(r)}^{*})\leq 2r$, we have
$\|Y_{(r)}-X_{(r)}^{*}\|_{\mathrm{F}}^{2}=\sum_{k=1}^{2r}\sigma_{k}^{2}(Y_{(r)}-X_{(r)}^{*}).$ (8)
Write
$Y_{(r)}-X_{(r)}^{*}=(X^{*}-X_{(r)}^{*})+Z+(Y_{(r)}-Y).$
It follows from Ky Fan’s Theorem (Lemma 25) that,
$\sqrt{\sum_{i=1}^{2r}\sigma_{i}^{2}(Y_{(r)}-X_{(r)}^{*})}\leq\sqrt{\sum_{i=1}^{2r}\sigma_{i}^{2}(X^{*}-X_{(r)}^{*})}+\sqrt{\sum_{i=1}^{2r}\sigma_{i}^{2}(Z)}+\sqrt{\sum_{i=1}^{2r}\sigma_{i}^{2}(Y_{(r)}-Y)}.$
Using the identity $\sigma_{i}(X^{*}-X^{*}_{(r)})=\sigma_{r+i}(X^{*})$ and likewise for $Y$, we obtain
$\|Y_{(r)}-X^{*}_{(r)}\|_{\mathrm{F}}$ $\leq\sqrt{\sum_{i=1}^{2r}\sigma_{r+i}^{2}(X^{*})}+\sqrt{\sum_{i=1}^{2r}\sigma_{i}^{2}(Z)}+\sqrt{\sum_{i=1}^{2r}\sigma_{r+i}^{2}(Y)}.$ (9)
To further simplify the third term, we apply Weyl’s inequality (Theorem 24) to obtain
$\sigma_{r+i}(Y)\leq\sigma_{r+i}(X^{*})+\sigma_{1}(Z),$
which gives
$\sigma_{r+i}^{2}(Y)\leq 2\left\{\sigma_{r+i}^{2}(X^{*})+\sigma_{1}^{2}(Z)\right\}.$
Substituting this into (9), we get
$\|Y_{(r)}-X^{*}_{(r)}\|_{\mathrm{F}}$ $\leq\sqrt{\sum_{i=1}^{2r}\sigma_{r+i}^{2}(X^{*})}+\sqrt{\sum_{i=1}^{2r}\sigma_{i}^{2}(Z)}+\sqrt{2\sum_{i=1}^{2r}\sigma_{r+i}^{2}(X^{*})}+2\sqrt{r}\sigma_{1}(Z)$
$\leq(1+\sqrt{2})\sqrt{\sum_{i=1}^{2r}\sigma_{r+i}^{2}(X^{*})}+({2+\sqrt{2}})\sqrt{r}\|Z\|$
$=(1+\sqrt{2})\sqrt{\sum_{i=r+1}^{3r}\sigma_{i}^{2}(X^{*})}+({2+\sqrt{2}})\sqrt{r}\|Z\|,$
where the second inequality follows from the fact that for any $k\geq 1$, we have $\sigma_{k}(Z)\leq\sigma_{1}(Z)=\|Z\|.$ Substituting the bound of $\|Y_{(r)}-X^{*}_{(r)}\|_{\mathrm{F}}$ into (7), we obtain
$\|Y_{(r)}-X^{*}\|_{\mathrm{F}}\leq(2+\sqrt{2})\sqrt{\sum_{i=r+1}^{\min\{m,n\}}\sigma^{2}_{i}(X^{*})}+(2+\sqrt{2})\sqrt{r}\|Z\|.$
Adaptivity in Theorem 2. For any prescribed rank $r$, the rank-$r$ truncated SVD estimator $Y_{(r)}$ satisfies the guarantee in Theorem 2. In other words, the theorem holds uniformly over all $r\geq 1$, so it applies to the vanilla SVD at any user-specified target rank.
Constant in (6). Tracing the proof of Theorem 2 shows that one may take the explicit constant $C=2+\sqrt{2}\approx 3.414$ in (6). We did not attempt to optimize this further.
Bias–variance tradeoff. Theorem 2 exhibits a clear bias–variance tradeoff for rank-$r$ estimation. The bias term $\xi_{(r)}=\|X^{*}-X_{(r)}^{*}\|_{\mathrm{F}}$ is the best possible rank-$r$ approximation error of $X^{*}$ and decreases as $r$ increases. The standard deviation term, being order $(\sqrt{r}\|Z\|)$, captures the cost of estimating additional singular components under noise and increases with $r$.
Remark 3.
Suppose $Y\in\mathbb{R}^{m\times n}$ and let $Y_{(r)}=U_{(r)}\,\Sigma_{(r)}\,V_{(r)}^{\top}$ be a rank-$r$ factorization (e.g. the truncated SVD), where $U_{(r)}\in\mathbb{R}^{m\times r}$, $V_{(r)}\in\mathbb{R}^{n\times r}$ satisfy $U_{(r)}^{\top}U_{(r)}=I_{r}$, $V_{(r)}^{\top}V_{(r)}=I_{r}$, and $\Sigma_{(r)}=\mathrm{diag}(\sigma_{1},\dots,\sigma_{r})\in\mathbb{R}^{r\times r}$. For a dense multiplication with $v\in\mathbb{R}^{n}$, computing $Yv$ costs $O(mn)$ operations. In contrast, using the factorization we compute successively
$t=V_{(r)}^{\top}v\in\mathbb{R}^{r}\ (\text{cost }\approx 2nr),\quad s=\Sigma_{(r)}t\in\mathbb{R}^{r}\ (\text{cost }\approx r),\quad w=U_{(r)}s\in\mathbb{R}^{m}\quad(\text{cost }\approx 2mr).$
Hence $Y_{(r)}v\;=\;U_{(r)}\,\Sigma_{(r)}\,V_{(r)}^{\top}v$ can be computed in $O\big(mr+nr\big)$ operations.
Beyond computational savings, the SVD factorization also reduces storage costs. Storing the full matrix $Y$ requires $O(mn)$ memory, whereas storing the factors $U_{(r)}\in\mathbb{R}^{m\times r}$, $\Sigma_{(r)}\in\mathbb{R}^{r\times r}$, and $V_{(r)}\in\mathbb{R}^{n\times r}$ requires only $O(mr+nr)$ scalars.
Therefore, both computation and storage using the truncated SVD are significantly smaller than in the full matrix case, especially when $r\ll\min\{m,n\}$.
In the following, we illustrate the application of Theorem 2 in three matrix estimation settings.
Corollary 3.
Suppose (4) holds and that the entries of $Z$ are independent sub-Gaussian random variables with
$\|Z_{ij}\|_{\psi_{2}}\leq\kappa.$
Then with probability at least $1-\exp(-(m+n))$, there exists a universal constant $C>0$ such that
$\|Y_{(r)}-X^{*}\|_{\mathrm{F}}\leq C\big(\|X_{(r)}^{*}-X^{*}\|_{\mathrm{F}}+\kappa\sqrt{r(m+n)}\big).$
Proof of Corollary 3.
From Vershynin, (2018) section 4.4.2, it follows that $\|Z\|\leq C_{1}\kappa\sqrt{m+n}$ with probability at least $1-\exp(-(m+n))$ . The desired result follows immediately from Theorem 2. ∎
Corollary 4.
Suppose $Z\in\mathbb{R}^{m\times n}$ is a sub-Gaussian random matrix in the sense that for any $v\in\mathbb{R}^{m}$ with $\|v\|_{2}=1$, and $w\in\mathbb{R}^{n}$ with $\|w\|_{2}=1$, it holds that
$\|v^{\top}Zw\|_{\psi_{2}}\leq\kappa.$
Then with probability at least $1-\exp(-(m+n))$, there exists a universal constant $C>0$ such that
$\|Y_{(r)}-X^{*}\|_{\mathrm{F}}\leq C\big(\|X_{(r)}^{*}-X^{*}\|_{\mathrm{F}}+\kappa\sqrt{r(m+n)}\big).$
Proof of Corollary 4.
By Lemma 12, $\|Z\|\leq C_{1}\sqrt{m+n}$ with probability at least $1-\exp(-(m+n))$. The desired result follows immediately from Theorem 2. ∎
Corollary 5.
Let $Z_{1},\dots,Z_{N}\in\mathbb{R}^{n}$ be i.i.d. mean-zero sub-Gaussian random vectors with covariance $\Sigma$. Define the sample covariance matrix
$Y=\frac{1}{N}\sum_{k=1}^{N}Z_{k}Z_{k}^{\top},$
and let $X^{*}=\mathbb{E}[Y]$ be the population covariance matrix. Suppose $Z_{k}$ is sub-Gaussian in the sense that
$\|u^{\top}Z_{k}\|_{\psi_{2}}\leq\kappa\quad\text{for all }u\in\mathbb{R}^{n},\ \|u\|_{2}=1.$
Then with probability at least $1-2\exp(-n)$, there exists a universal constant $C>0$ such that
$\|Y_{(r)}-X^{*}\|_{\mathrm{F}}\leq C\Bigl(\|X_{(r)}^{*}-X^{*}\|_{\mathrm{F}}+\kappa^{2}\sqrt{r}\Bigl[\sqrt{\frac{n}{N}}+\frac{n}{N}\Bigr]\Bigr).$
Proof of Corollary 5.
Let $\widetilde{Z}:=Y-X^{*}.$ By Remark 4.7.3 of Vershynin, (2018), with probability at least $1-\exp(-cn)$ one has $\|\widetilde{Z}\|\;\leq\;C_{1}\kappa^{2}\Bigl[\sqrt{\frac{n}{N}}+\frac{n}{N}\Bigr].$ Applying Theorem 2 with $Z$ replaced by $\widetilde{Z}$ yields the desired bound. ∎
4 Numerical experiments
In many applied settings, selecting the single best rank $r$ is inherently difficult in finite samples with unknown latent structure. Our analysis instead provides guarantees for singular-value decompositions in both the matrix (SVD) and tensor (HOSVD) regimes across a broad range of $r$. Consequently, rather than hinging on a brittle rank-selection heuristic, we empirically demonstrate that both truncated SVD and one-step HOSVD (Algorithm 1) are robust whenever ranks are chosen neither too small (underfitting/bias) nor too large (overfitting/variance). This is consistent with the bias–variance tradeoff formalized in Theorems 2 and 1. Moreover, in practice the choice of $r$, is frequently driven by diverse aims other than merely minimizing mean-squared error, including visualization, interpretability, or downstream decision tasks. This is reflected in widespread, long-standing use across biology/genomics (Price et al.,, 2006), economics (Filmer and Pritchett,, 2001), and computer vision (Turk and Pentland,, 1991).
We evaluate these predictions on (i) a real 3D brain-MRI dataset with controlled additive noise (IXI; T1-weighted volumes (IXI,, 2002)), and (ii) simulated data for both matrices and third–order tensors.
4.1 Real data: 3D brain MRI with controlled noise
We evaluate on a T1-weighted brain MRI volume drawn from the IXI dataset (IXI,, 2002). The selected volume is a 3D tensor of shape $256\times 256\times 150$. To probe low-rank behavior in a realistic preprocessing pipeline, the volume is prepared in two wavelet-smoothed variants using standard orthonormal families—(a) Daubechies-6 (db6) and (b) Symlet-8 (sym8) (Daubechies,, 1988, 1992).
We implement separable 3D multilevel decompositions (wavedecn) with periodic boundary handling, and apply subband-wise soft-thresholding via BayesShrink (Chang et al.,, 2000; Donoho and Johnstone,, 1994). The literature supports wavelet shrinkage as an effective denoising/smoothing step for MR magnitude images (Nowak,, 1999; Wood and Johnson,, 1999).
Noise model and ranks. For our prepared volume $X$ and each noise level $\lambda\in\{0.01,0.05,0.1,0.2\}$ we generate
$Y\;=\;X^{*}+Z,\qquad Z_{ijk}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathcal{N}(0,\kappa^{2}),\quad\kappa\;=\;\lambda\,\frac{\|X^{*}\|_{\mathrm{F}}}{\sqrt{p_{1}p_{2}p_{3}}}\,$,
which keeps the per-entry SNR comparable across images and variants. We compute one-step HOSVD reconstructions (Algorithm 1) $\widetilde{X}$ at Tucker rank $(r,r,r)$ for $r\in\{50,65,80\}.$
Findings. Figure 1 shows representative mid-sagittal slices from the reconstructions of the wavelet-smoothed variants of single volume. Qualitatively, the slices exhibit the expected progression: very small ranks $r$ lead to underfitting with over-smoothed, low-rank artifacts, while an intermediate range of $r$ preserves cortical detail and simultaneously suppresses background fluctuations.
Figure 1: Mid-sagittal MRI slice of the reconstructed volume at selected ranks and noise levels $\lambda$ for the db6 variant (left), and for the sym8 variant (right).
4.2 Synthetic data: Tensors
We generate third-order tensors $X^{*}\in\mathbb{R}^{p\times p\times p}$ in the Tucker form
$X^{*}\;=\;\mathcal{G}\times_{1}U_{1}\times_{2}U_{2}\times_{3}U_{3},\qquad U_{1},U_{2},U_{3}\in\mathbb{R}^{p\times s},\ U_{1}^{\top}U_{1}=U_{2}^{\top}U_{2}=U_{3}^{\top}U_{3}=I_{s},$
where $U_{1},U_{2},U_{3}$ are drawn independently with orthonormal columns. The core $\mathcal{G}\in\mathbb{R}^{s\times s\times s}$ is diagonal along its main fiber with entries $\{\gamma_{i}\}_{i=1}^{s}$ decaying exponentially, i.e. $\gamma_{i}=\beta^{i},\,\beta=0.8$. We observe $Y=X^{*}+Z$ with i.i.d. Gaussian noise $Z_{ijk}\sim\mathcal{N}(0,1)$ entry-wise. The SNR is controlled by scaling the signal such that
$\|X^{*}\|_{\mathrm{F}}\;=\;\lambda\,\sqrt{p^{3}},\qquad\lambda\in\{10,50\}.$
The estimator $\widetilde{X}$ is obtained via the one-step HOSVD (Algorithm 1) at the Tucker rank $(r,r,r)$.
Dimensions and evaluated ranks. We vary $p\in\{20,50,75,100\}$ and for each $(p,s)$, we evaluate two ranks $r$ as listed below:
$p$ $s$ $r$ (target)
20 15 10, 12
50 25 10, 15
75 20 10, 15
100 80 30, 40
Table 1: Synthetic tensor grid: dimensions and evaluated Tucker ranks.
Each configuration $(p,s,\lambda,r)$ is repeated $R=50$ times with freshly drawn $(U_{1},U_{2},U_{3})$ and noise. Implementations use exact batched SVD on GPU via TensorLy (PyTorch backend) (Kossaifi et al.,, 2019). We report the sample mean and standard error of the relative Frobenius error
$\mathrm{RelErr}(\widetilde{X};X^{*})\;=\;\frac{\|\widetilde{X}-X^{*}\|_{\mathrm{F}}}{\|X^{*}\|_{\mathrm{F}}}\,$,
summarized on the right panel of Table 2. We observe that the error consistently decreases as the SNR parameter $\lambda$ increases. Overall, one-step HOSVD is robust across the tested sizes and ranks, yielding accurate estimates on the synthetic tensors.
4.3 Synthetic data: Matrices
We examine the matrix analogue under a controlled SNR design that mirrors the tensor experiments. For each configuration, we draw an $m\times n$ latent signal
$X^{*}\;=\;U\,\Sigma\,V^{\top},\qquad U\in\mathbb{R}^{m\times n},\ \ V\in\mathbb{R}^{n\times n},\ \ U^{\top}U=V^{\top}V=I_{n},$
where $\Sigma=\mathrm{diag}(\sigma_{1},\ldots,\sigma_{n})$ has exponentially decaying singular values, i.e. $\sigma_{i}=\beta^{\,i}$ with $\beta=0.8$. To mirror the tensor ambient sizes while probing rectangularity, we set $m=5p$ and $n=s$ according to the tensor grid in Table 1. We observe $Y=X^{\star}+Z$, where $Z_{ij}\sim\mathcal{N}(0,1)$ is i.i.d. Gaussian noise entry-wise. The SNR is controlled by scaling the signal such that
$\|X^{*}\|_{\mathrm{F}}\;=\;\lambda\,\sqrt{mn},\qquad\lambda\in\{10,50\}.$
This design varies SNR via $\lambda$ while keeping the per–entry noise scale identical across sizes. Each configuration $(m,n,\lambda,r)$ is repeated $R=50$ times with independent draws of $(U,V)$ and noise, and we report the mean and standard error of $\mathrm{RelErr}$. We use the truncated SVD estimator $Y_{(r)}$ as in Theorem 2. Results are shown in the left panel of Table 2. The truncated SVD is robust over a wide range of selected ranks.
Table 2: Results for matrix (left) and tensor (right).
Matrix
$\lambda$ $m$ $n$ $r$ Mean (SE)
10 100 15 10 0.1178 (0.00125)
12 0.0906 (0.00164)
250 25 10 0.1254 (0.00044)
15 0.0868 (0.00077)
375 20 10 0.1279 (0.00045)
15 0.0927 (0.00076)
500 80 30 0.0698 (0.00032)
40 0.0795 (0.00035)
50 100 15 10 0.0844 (0.00007)
12 0.0181 (0.00037)
250 25 10 0.1079 (0.00002)
15 0.0378 (0.00009)
375 20 10 0.1068 (0.00002)
15 0.0349 (0.00008)
500 80 30 0.0134 (0.00008)
40 0.0156 (0.00006)
Tensor
$\lambda$ $p$ $s$ $r$ Mean (SE)
10 20 15 10 0.1092 (0.00031)
12 0.0776 (0.00058)
50 25 10 0.1081 (0.00002)
15 0.0402 (0.00012)
75 20 10 0.1081 (0.00002)
15 0.0353 (0.00004)
100 80 30 0.0204 (0.00007)
40 0.0294 (0.00007)
50 20 15 10 0.1018 (0.00001)
12 0.0599 (0.00002)
50 25 10 0.1073 (0.00000)
15 0.0352 (0.00000)
75 20 10 0.1067 (0.00000)
15 0.0333 (0.00000)
100 80 30 0.0039 (0.00001)
40 0.0058 (0.00001)
5 Discussion
This work provides a rank–adaptive analysis of HOSVD-based tensor denoising without assuming exact low-rank. Our main theorem (Theorem 1) yields an explicit bias–variance decomposition
$\|\widetilde{X}-X^{*}\|_{\mathrm{F}}\;\lesssim\;\kappa\,\sqrt{\textstyle\sum_{k=1}^{3}p_{k}r_{k}\;+\;r_{1}r_{2}r_{3}}\;+\;\xi_{(r_{1},r_{2},r_{3})},$
uniformly over all user–specified target Tucker ranks. The variance term scales with the effective degrees of freedom of the Tucker model, while the bias term is the best achievable approximation error at those ranks. Together with the matrix counterpart (Theorem 2), the results unify classical SVD intuition with the multilinear (tensor) setting and rigorously justify a practice common in applications: choose ranks large enough to suppress bias but not so large as to amplify noise. Our experiments on IXI brain MRI and controlled synthetic data corroborate this picture.
Future work.
(i) Beyond i.i.d. noise and full observations. Extending the theory to heteroskedastic or correlated perturbations (e.g. spatially correlated fields, Rician-like MRI noise) and to incomplete observations (tensor completion, masked entries) is natural. We expect the variance term to inherit problem-dependent effective dimensions (e.g. leverage scores or sampling densities), while the bias term remains $\xi_{(r_{1},r_{2},r_{3})}$. A key technical step is replacing isotropic concentration with mode-wise covariance-aware bounds for $\mathcal{M}_{k}(Z)$ and their projections. (ii) Higher orders and higher ranks. The proof strategy appears to scale to $d$th-order tensors with the clean generalization
$\|\widetilde{X}-X^{*}\|_{\mathrm{F}}\;\lesssim\;\kappa\,\sqrt{\textstyle\sum_{k=1}^{d}p_{k}r_{k}\;+\;\prod_{k=1}^{d}r_{k}}\;+\;\xi_{(r_{1},\ldots,r_{d})},$
under appropriately formulated spectral gaps for each unfolding. This would offer a unified bias–variance law across orders and ranks, and would help explain the strong empirical performance of HOSVD-style estimators in truly high-order settings.
Supplementary Materials for “Bias–variance Tradeoff in Tensor Estimation”
Without loss of generality, if $W\sim\mathrm{subGaussian}(0,\kappa^{2})$, we assume throughout, unless otherwise specified in the appendix, that $\mathbb{E}[W^{2}]=\kappa^{2}$.
Appendix A Proof of Theorem 1
Proof of Theorem 1.
For any orthogonal matrix $U\in\mathbb{O}^{p\times r}$, we denote $U_{\perp}\in\mathbb{O}^{p\times(p-r)}$ to be the orthogonal complement of $U$, and
$\mathcal{P}_{U}={U}{U}^{\top}\quad\text{and}\quad\mathcal{P}_{U_{\perp}}={U}_{\perp}{U}_{\perp}^{\top}=I_{p}-\mathcal{P}_{U}.$
For $k\in\{1,2,3\}$, let $U_{k}^{*}\in\mathbb{O}^{p_{k}\times r_{k}}$ be the matrix whose columns corresponds to the top $r_{k}$ singular vectors of $\mathcal{M}_{k}(X^{*})$.
Throughout this proof, we assume the following good events hold:
$\sup_{\begin{subarray}{c}A\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}},\\ \|A\|_{\mathrm{F}}\leq 1,A\in\mathcal{T}_{(r_{1},r_{2},r_{3})}\end{subarray}}\langle Z,A\rangle\leq C\kappa\sqrt{r_{1}r_{2}r_{3}+p_{1}r_{1}+p_{2}r_{2}+p_{3}r_{3}};$ (10)
$\left\|\mathcal{M}_{1}(Z)\cdot W_{2}\otimes W_{3}\right\|\leq C\kappa\left(\sqrt{p_{1}+s_{2}s_{3}}\right)\text{ for non-random }W_{2}\in\mathbb{O}^{p_{2}\times s_{2}},\,W_{3}\in\mathbb{O}^{p_{3}\times s_{3}};$ (11)
$\left\|\sin\Theta(U_{k}^{(0)},U_{k}^{*})\right\|=\|U_{k}^{*\top}U_{k}^{(0)}\|\leq\frac{1}{2\sqrt{r_{\max}}}\text{ for }k\in\{1,2,3\}.$ (12)
Indeed in Lemma 7, Lemma 8 and Corollary 14, we show that (10), (11), and (12) hold with probability at least $1-C\exp(-cp_{\min})$.
Note that
$\left\|\widetilde{X}-X^{*}\right\|_{\mathrm{F}}=\left\|Y\times_{1}\mathcal{P}_{U_{1}^{(1)}}\times_{2}\mathcal{P}_{U_{2}^{(1)}}\times_{3}\mathcal{P}_{U_{3}^{(1)}}-X^{*}\right\|_{\mathrm{F}}$
$\leq\left\|X^{*}\times_{1}\mathcal{P}_{U_{1}^{(1)}}\times_{2}\mathcal{P}_{U_{2}^{(1)}}\times_{3}\mathcal{P}_{U_{3}^{(1)}}-X^{*}\right\|_{\mathrm{F}}+\left\|Z\times_{1}\mathcal{P}_{U_{1}^{(1)}}\times_{2}\mathcal{P}_{U_{2}^{(1)}}\times_{3}\mathcal{P}_{U_{3}^{(1)}}\right\|_{\mathrm{F}}$
$=\mathrm{I}_{1}+\mathrm{I}_{2}.$
Step 1. For the term $\mathrm{I}_{2}$, observe that
$\mathrm{I}_{2}$ $=\left\|Z\times_{1}\mathcal{P}_{U_{1}^{(1)}}\times_{2}\mathcal{P}_{U_{2}^{(1)}}\times_{3}\mathcal{P}_{U_{3}^{(1)}}\right\|_{\mathrm{F}}$
$=\sup_{\begin{subarray}{c}W\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}},\|W\|_{\mathrm{F}}\leq 1\end{subarray}}\left\langle Z\times_{1}\mathcal{P}_{U_{1}^{(1)}}\times_{2}\mathcal{P}_{U_{2}^{(1)}}\times_{3}\mathcal{P}_{U_{3}^{(1)}},\,W\right\rangle$
$=\sup_{\begin{subarray}{c}W\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}},\|W\|_{\mathrm{F}}\leq 1\end{subarray}}\left\langle Z,\,\mathcal{W}\times_{1}\mathcal{P}_{U_{1}^{(1)}}\times_{2}\mathcal{P}_{U_{2}^{(1)}}\times_{3}\mathcal{P}_{U_{3}^{(1)}}\right\rangle$
$\leq\sup_{\begin{subarray}{c}A\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}},\\ \|A\|_{\mathrm{F}}\leq 1,A\in\mathcal{T}_{(r_{1},r_{2},r_{3})}\end{subarray}}\langle Z,A\rangle\leq C\kappa\left(r_{1}r_{2}r_{3}+\sum_{k=1}^{3}p_{k}r_{k}\right)^{1/2},$
where the second equality follows from the duality of the Frobenius norm, and the last inequality follows from (10).
Step 2. For the term $\mathrm{I}_{1}$, we have that
$\mathrm{I}_{1}$ $=\left\|X^{*}\times_{1}\mathcal{P}_{U_{1}^{(1)}}\times_{2}\mathcal{P}_{U_{2}^{(1)}}\times_{3}\mathcal{P}_{U_{3}^{(1)}}-X^{*}\right\|_{\mathrm{F}}$
$\leq\left\|X^{*}\times_{1}(I_{p_{1}}-\mathcal{P}_{U_{1}^{(1)}})\right\|_{\mathrm{F}}+\left\|X^{*}\times_{1}\mathcal{P}_{U_{1}^{(1)}}\times_{2}(I_{p_{2}}-\mathcal{P}_{U_{2}^{(1)}})\right\|_{\mathrm{F}}$
$\quad+\left\|X^{*}\times_{1}\mathcal{P}_{U_{1}^{(1)}}\times_{2}\mathcal{P}_{U_{2}^{(1)}}\times_{3}(I_{p_{3}}-\mathcal{P}_{U_{3}^{(1)}})\right\|_{\mathrm{F}}$
$\leq\sum_{k=1}^{3}\left\|X^{*}\times_{k}(I_{p_{k}}-\mathcal{P}_{U_{k}^{(1)}})\right\|_{\mathrm{F}},$
where the last inequality follows from the observation that
$\left\|X^{*}\times_{1}\mathcal{P}_{U_{1}^{(1)}}\times_{2}(I_{p_{2}}-\mathcal{P}_{U_{2}^{(1)}})\right\|_{\mathrm{F}}\leq\left\|X^{*}\times_{2}(I_{p_{2}}-\mathcal{P}_{U_{2}^{(1)}})\right\|_{\mathrm{F}}\|\mathcal{P}_{U_{1}^{(1)}}\|\leq\left\|X^{*}\times_{2}(I_{p_{2}}-\mathcal{P}_{U_{2}^{(1)}})\right\|_{\mathrm{F}}.$
We only consider the case when $k=1$, since the same arguments apply for $k=2,3$. We have that
$\left\|X^{*}\times_{1}(I_{p_{1}}-\mathcal{P}_{U_{1}^{(1)}})\right\|_{\mathrm{F}}$
$\leq\left\|X^{*}\times_{1}(I_{p_{1}}-\mathcal{P}_{U_{1}^{(1)}})\times_{2}\mathcal{P}_{U_{2}^{*}}\right\|_{\mathrm{F}}+\left\|X^{*}\times_{1}(I_{p_{1}}-\mathcal{P}_{U_{1}^{(1)}})\times_{2}(I_{p_{2}}-\mathcal{P}_{U_{2}^{*}})\right\|_{\mathrm{F}}$
$\leq\left\|X^{*}\times_{1}(I_{p_{1}}-\mathcal{P}_{U_{1}^{(1)}})\times_{2}\mathcal{P}_{U_{2}^{*}}\right\|_{\mathrm{F}}+\left\|X^{*}\times_{2}(I_{p_{2}}-\mathcal{P}_{U_{2}^{*}})\right\|_{\mathrm{F}}$
$\leq\left\|X^{*}\times_{1}(I_{p_{1}}-\mathcal{P}_{U_{1}^{(1)}})\times_{2}\mathcal{P}_{U_{2}^{*}}\times_{3}\mathcal{P}_{U_{3}^{*}}\right\|_{\mathrm{F}}+\left\|X^{*}\times_{1}(I_{p_{1}}-\mathcal{P}_{U_{1}^{(1)}})\times_{2}\mathcal{P}_{U_{2}^{*}}\times_{3}(I_{p_{3}}-\mathcal{P}_{U_{3}^{*}})\right\|_{\mathrm{F}}$
$+\left\|X^{*}\times_{2}(I_{p_{2}}-\mathcal{P}_{U_{2}^{*}})\right\|_{\mathrm{F}}$
$\leq\left\|X^{*}\times_{1}(I_{p_{1}}-\mathcal{P}_{U_{1}^{(1)}})\times_{2}\mathcal{P}_{U_{2}^{*}}\times_{3}\mathcal{P}_{U_{3}^{*}}\right\|_{\mathrm{F}}+\left\|X^{*}\times_{3}(I_{p_{3}}-\mathcal{P}_{U_{3}^{*}})\right\|_{\mathrm{F}}+\left\|X^{*}\times_{2}(I_{p_{2}}-\mathcal{P}_{U_{2}^{*}})\right\|_{\mathrm{F}}$
$=\left\|X^{*}\times_{1}(I_{p_{1}}-\mathcal{P}_{U_{1}^{(1)}})\times_{2}\mathcal{P}_{U_{2}^{*}}\times_{3}\mathcal{P}_{U_{3}^{*}}\right\|_{\mathrm{F}}+2\xi_{(r_{1},r_{2},r_{3})},$
where the second follows from that $\|I_{p_{1}}-\mathcal{P}_{U_{1}^{(1)}}\|\leq 1$, and the equality follows from Lemma 26.
Step 3. We bound $\|X^{*}\times_{1}(I_{p_{1}}-\mathcal{P}_{U_{1}^{(1)}})\times_{2}\mathcal{P}_{U_{2}^{*}}\times_{3}\mathcal{P}_{U_{3}^{*}}\|_{\mathrm{F}}$ in this step. Consider the different but relevant quantity
$\left\|X^{*}\times_{1}(I_{p_{1}}-\mathcal{P}_{U_{1}^{(1)}})\times_{2}\mathcal{P}_{U_{2}^{(0)}}\times_{3}\mathcal{P}_{U_{3}^{(0)}}\right\|_{\mathrm{F}}$
$=\left\|(I_{p_{1}}-\mathcal{P}_{U_{1}^{(1)}})\cdot\mathcal{M}_{1}(X^{*})\cdot(\mathcal{P}_{U_{2}^{(0)}}\otimes\mathcal{P}_{U_{3}^{(0)}})\right\|_{\mathrm{F}}$
$=\left\|(I_{p_{1}}-\mathcal{P}_{U_{1}^{(1)}})\cdot\mathcal{M}_{1}(X^{*})\cdot(U_{2}^{(0)}\otimes U_{3}^{(0)})(U_{2}^{(0)}\otimes U_{3}^{(0)})^{\top}\right\|_{\mathrm{F}}$
$=\left\|(I_{p_{1}}-\mathcal{P}_{U_{1}^{(1)}})\cdot\mathcal{M}_{1}(X^{*})\cdot(U_{2}^{(0)}\otimes U_{3}^{(0)})\right\|_{\mathrm{F}}$
$\geq\left\|(I_{p_{1}}-\mathcal{P}_{U_{1}^{(1)}})\cdot\mathcal{M}_{1}(X^{*})\cdot\mathcal{P}_{U_{2}^{*}\otimes U_{3}^{*}}\cdot(U_{2}^{(0)}\otimes U_{3}^{(0)})\right\|_{\mathrm{F}}$
$\quad-\left\|(I_{p_{1}}-\mathcal{P}_{U_{1}^{(1)}})\cdot\mathcal{M}_{1}(X^{*})\cdot(I_{p_{2}p_{3}}-\mathcal{P}_{U_{2}^{*}\otimes U_{3}^{*}})\cdot(U_{2}^{(0)}\otimes U_{3}^{(0)})\right\|_{\mathrm{F}}$
$=\mathrm{II}_{1}-\mathrm{II}_{2},$
where the first inequality follows from that $(U_{2}^{(0)}\otimes U_{3}^{(0)})\in\mathbb{O}^{(p_{2}p_{3})\times(r_{2}r_{3})}$, and Lemma 19.
Before analyzing the terms $\mathrm{II}_{1}$ and $\mathrm{II}_{2}$, we firstly note that
$(U_{2}^{*}\otimes U_{3}^{*})^{\top}\cdot(U_{2}^{(0)}\otimes U_{3}^{(0)})$ $=(U_{2}^{*\,\top}U_{2}^{(0)})\otimes(U_{3}^{*\,\top}U_{3}^{(0)}),$ (13)
$\sigma_{\min}\left((U_{2}^{*\,\top}U_{2}^{(0)})\otimes(U_{3}^{*\,\top}U_{3}^{(0)})\right)$ $=\sigma_{\min}\left(U_{2}^{*\,\top}U_{2}^{(0)}\right)\sigma_{\min}\left(U_{3}^{*\,\top}U_{3}^{(0)}\right),$
$\sigma_{\min}^{2}\left(U_{k}^{*\,\top}U_{k}^{(0)}\right)$ $=1-\left\|U_{k\,\perp}^{*\,\top}U_{k}^{(0)}\right\|^{2}=1-\left\|\sin\Theta(U_{k}^{*},U_{k}^{(0)})\right\|^{2},$
which hold due to the properties of the Kronecker product and Lemma 22.
For the term $\mathrm{II}_{1}$, we have that
$\mathrm{II}_{1}$ $=\left\|(I_{p_{1}}-\mathcal{P}_{U_{1}^{(1)}})\cdot\mathcal{M}_{1}(X^{*})\cdot(U_{2}^{*}\otimes U_{3}^{*})\cdot(U_{2}^{*}\otimes U_{3}^{*})^{\top}\cdot(U_{2}^{(0)}\otimes U_{3}^{(0)})\right\|_{\mathrm{F}}$
$\geq\left\|(I_{p_{1}}-\mathcal{P}_{U_{1}^{(1)}})\cdot\mathcal{M}_{1}(X^{*})\cdot(U_{2}^{*}\otimes U_{3}^{*})\right\|_{\mathrm{F}}\sigma_{\min}(U_{2}^{*\,\top}U_{2}^{(0)})\sigma_{\min}(U_{3}^{*\,\top}U_{3}^{(0)})$
$=\left\|X^{*}\times_{1}(I_{p_{1}}-\mathcal{P}_{U_{1}^{(1)}})\times_{2}\mathcal{P}_{U_{2}^{*}}\times_{3}\mathcal{P}_{U_{3}^{*}}\right\|_{\mathrm{F}}\sqrt{\left(1-\left\|\sin\Theta(U_{2}^{*},U_{2}^{(0)})\right\|^{2}\right)\left(1-\left\|\sin\Theta(U_{3}^{*},U_{3}^{(0)})\right\|^{2}\right)}$
$\geq\frac{3}{4}\left\|X^{*}\times_{1}(I_{p_{1}}-\mathcal{P}_{U_{1}^{(1)}})\times_{2}\mathcal{P}_{U_{2}^{*}}\times_{3}\mathcal{P}_{U_{3}^{*}}\right\|_{\mathrm{F}}$
where the first inequality follows from Lemma 18, and the second equality follows from (13), and the last inequality follows from (12).
For the term $\mathrm{II}_{2}$, we have that
$\mathrm{II}_{2}$ $=\left\|(I_{p_{1}}-\mathcal{P}_{U_{1}^{(1)}})\cdot\mathcal{M}_{1}(X^{*})\cdot(I_{p_{2}p_{3}}-\mathcal{P}_{U_{2}^{*}\otimes U_{3}^{*}})\cdot(U_{2}^{(0)}\otimes U_{3}^{(0)})\right\|_{\mathrm{F}}$
$=\left\|(I_{p_{1}}-\mathcal{P}_{U_{1}^{(1)}})\cdot\mathcal{M}_{1}(X^{*})\cdot(U_{2}^{*}\otimes U_{3}^{*})_{\perp}\cdot(U_{2}^{*}\otimes U_{3}^{*})_{\perp}^{\top}\cdot(U_{2}^{(0)}\otimes U_{3}^{(0)})\right\|_{\mathrm{F}}$
$\leq\left\|I_{p_{1}}-\mathcal{P}_{U_{1}^{(1)}}\right\|\left\|\mathcal{M}_{1}(X^{*})\cdot(U_{2}^{*}\otimes U_{3}^{*})_{\perp}\right\|_{\mathrm{F}}\left\|(U_{2}^{*}\otimes U_{3}^{*})_{\perp}^{\top}\cdot(U_{2}^{(0)}\otimes U_{3}^{(0)})\right\|$
$=\left\|\mathcal{M}_{1}(X^{*})\cdot(U_{2}^{*}\otimes U_{3}^{*})_{\perp}\right\|_{\mathrm{F}}.$
Note that
$\left\|\mathcal{M}_{1}(X^{*})\cdot(U_{2}^{*}\otimes U_{3}^{*})_{\perp}\right\|_{\mathrm{F}}$
$=\left\|\mathcal{M}_{1}(X^{*})\cdot[U_{2\,\perp}^{*}\otimes U_{3}^{*}\quad U_{2}^{*}\otimes U_{3\,\perp}^{*}\quad U_{2\,\perp}^{*}\otimes U_{3\,\perp}^{*}]\right\|_{\mathrm{F}}$
$=\sqrt{\left\|\mathcal{M}_{1}(X^{*})\cdot(U_{2\,\perp}^{*}\otimes U_{3}^{*})\right\|_{\mathrm{F}}^{2}+\left\|\mathcal{M}_{1}(X^{*})\cdot(U_{2}^{*}\otimes U_{3\,\perp}^{*})\right\|_{\mathrm{F}}^{2}+\left\|\mathcal{M}_{1}(X^{*})\cdot(U_{2\,\perp}^{*}\otimes U_{3\,\perp}^{*})\right\|_{\mathrm{F}}^{2}}$
$\leq\left\|\mathcal{M}_{1}(X^{*})\cdot(U_{2\,\perp}^{*}\otimes U_{3}^{*})\right\|_{\mathrm{F}}+\left\|\mathcal{M}_{1}(X^{*})\cdot(U_{2}^{*}\otimes U_{3\,\perp}^{*})\right\|_{\mathrm{F}}+\left\|\mathcal{M}_{1}(X^{*})\cdot(U_{2\,\perp}^{*}\otimes U_{3\,\perp}^{*})\right\|_{\mathrm{F}}$
$=\left\|X^{*}\times_{2}U_{2\,\perp}^{*}\times_{3}U_{3}^{*}\right\|_{\mathrm{F}}+\left\|X^{*}\times_{2}U_{2}^{*}\times_{3}U_{3\,\perp}^{*}\right\|_{\mathrm{F}}+\left\|X^{*}\times_{2}U_{2\,\perp}^{*}\times_{3}U_{3\,\perp}^{*}\right\|_{\mathrm{F}}$
$\leq\left\|X^{*}\times_{2}U_{2\,\perp}^{*}\right\|_{\mathrm{F}}+\left\|X^{*}\times_{3}U_{3\,\perp}^{*}\right\|_{\mathrm{F}}+\left\|X^{*}\times_{2}U_{2\,\perp}^{*}\right\|_{\mathrm{F}}$
$\leq 3\xi_{(r_{1},r_{2},r_{3})}$
where the last inequality follows from Lemma 26. Therefore,
$\mathrm{II}_{2}\leq 3\xi_{(r_{1},r_{2},r_{3})}.$
Combining $\mathrm{II}_{1}$ and $\mathrm{II}_{2}$, we have
$\left\|X^{*}\times_{1}(I_{p_{1}}-\mathcal{P}_{U_{1}^{(1)}})\times_{2}\mathcal{P}_{U_{2}^{*}}\times_{3}\mathcal{P}_{U_{3}^{*}}\right\|_{\mathrm{F}}\leq\frac{4}{3}\left\|X^{*}\times_{1}(I_{p_{1}}-\mathcal{P}_{U_{1}^{(1)}})\times_{2}\mathcal{P}_{U_{2}^{(0)}}\times_{3}\mathcal{P}_{U_{3}^{(0)}}\right\|_{\mathrm{F}}+4\xi_{(r_{1},r_{2},r_{3})}.$
Step 4. We bound $\left\|X^{*}\times_{1}(I_{p_{1}}-\mathcal{P}_{U_{1}^{(1)}})\times_{2}\mathcal{P}_{U_{2}^{(0)}}\times_{3}\mathcal{P}_{U_{3}^{(0)}}\right\|_{\mathrm{F}}$ in this step. Note that
$\|X^{*}\times_{1}(I_{p_{1}}-\mathcal{P}_{U_{1}^{(1)}})\times_{2}\mathcal{P}_{U_{2}^{(0)}}\times_{3}\mathcal{P}_{U_{3}^{(0)}}\|_{\mathrm{F}}$
$=$ $\|\mathcal{P}_{U_{1\perp}^{(1)}}\cdot\mathcal{M}_{1}(X^{*})\cdot(\mathcal{P}_{U_{2}^{(0)}}\otimes\mathcal{P}_{U_{3}^{(0)}})\|_{\mathrm{F}}=\|\mathcal{P}_{U_{1\perp}^{(1)}}\cdot\mathcal{M}_{1}(X^{*})\cdot(U_{2}^{(0)}\otimes U_{3}^{(0)})\|_{\mathrm{F}}$
It suffices to apply Lemma 20 to bound
$\|\mathcal{P}_{U_{1\perp}^{(1)}}\cdot\mathcal{M}_{1}(X^{*})\cdot(U_{2}^{(0)}\otimes U_{3}^{(0)})\|_{\mathrm{F}}.$
Since $U_{1\perp}^{(1)}$ corresponds to the SVD of $\mathcal{M}_{1}(Y)\cdot(U_{2}^{(0)}\otimes U_{3}^{(0)})$, let
$A=\mathcal{M}_{1}(Y)\cdot(U_{2}^{(0)}\otimes U_{3}^{(0)})\quad\text{and}\quad B=\mathcal{M}_{1}(X^{*})\cdot(U_{2}^{(0)}\otimes U_{3}^{(0)}).$
It follows from Lemma 20 that
$\left\|X^{*}\times_{1}(I_{p_{1}}-\mathcal{P}_{U_{1}^{(1)}})\times_{2}\mathcal{P}_{U_{2}^{(0)}}\times_{3}\mathcal{P}_{U_{3}^{(0)}}\right\|_{\mathrm{F}}=$ $\|\mathcal{P}_{U_{1\perp}^{(1)}}\cdot\mathcal{M}_{1}(X^{*})\cdot(U_{2}^{(0)}\otimes U_{3}^{(0)})\|_{\mathrm{F}}$
$\leq$ $C_{1}\|B-B_{r_{1}}\|_{F}+C_{2}\sqrt{r_{1}}\|A-B\|.$ (14)
Here
$\left\|A-B\right\|=$ $\left\|\mathcal{M}_{1}(Z)\cdot U_{2}^{(0)}\otimes U_{3}^{(0)}\right\|$
$=$ $\left\|\mathcal{M}_{1}(Z)\cdot(U^{*}_{2}U^{*\top}_{2}+U^{*}_{2\perp}U^{\top*}_{2\perp})U_{2}^{(0)}\otimes U_{3}^{(0)}\right\|$
$\leq$ $\left\|\mathcal{M}_{1}(Z)\cdot(U^{*}_{2}U^{*\top}_{2}U_{2}^{(0)})\otimes U_{3}^{(0)}\right\|+\left\|\mathcal{M}_{1}(Z)\cdot(U^{*}_{2\perp}U^{\top*}_{2\perp}U_{2}^{(0)})\otimes U_{3}^{(0)}\right\|$
$\leq$ $\left\|\mathcal{M}_{1}(Z)\cdot U^{*}_{2}\otimes U_{3}^{(0)}\right\|\|U^{*\top}_{2}U_{2}^{(0)}\|+\left\|\mathcal{M}_{1}(Z)\cdot U^{*}_{2\perp}\otimes U_{3}^{(0)}\right\|\|U^{\top*}_{2\perp}U_{2}^{(0)}\|.$ (15)
Note that
$\left\|\mathcal{M}_{1}(Z)\cdot U^{*}_{2}\otimes U_{3}^{(0)}\right\|\|U^{*\top}_{2}U_{2}^{(0)}\|=\left\|\mathcal{M}_{1}(Z)\cdot U^{*}_{2}\otimes(U^{*}_{3}U^{*\top}_{3}+U^{*}_{3\perp}U^{\top*}_{3\perp})U_{3}^{(0)}\right\|$
$\leq$ $\left\|\mathcal{M}_{1}(Z)\cdot U^{*}_{2}\otimes(U^{*}_{3}U^{*\top}_{3}U_{3}^{(0)})\right\|+\left\|\mathcal{M}_{1}(Z)\cdot U^{*}_{2}\otimes(U^{*}_{3\perp}U^{\top*}_{3\perp})U_{3}^{(0)}\right\|$
$\leq$ $\left\|\mathcal{M}_{1}(Z)\cdot U^{*}_{2}\otimes U^{*}_{3}\right\|\|U^{*\top}_{3}U_{3}^{(0)}\|+\left\|\mathcal{M}_{1}(Z)\cdot U^{*}_{2}\otimes U^{*}_{3\perp}\right\|\|U^{\top*}_{3\perp}U_{3}^{(0)}\|$
$\leq$ $C_{4}\kappa(\sqrt{p_{1}+r_{2}r_{3}})+C_{5}\kappa(\sqrt{p_{1}+p_{3}r_{2}})r^{-1/2}_{\max},$ (16)
where the last inequality follows from (11) and the fact that $\|U^{*\top}_{2}U_{2}^{(0)}\|\leq 1$. In addition
$\left\|\mathcal{M}_{1}(Z)\cdot U^{*}_{2\perp}\otimes U_{3}^{(0)}\right\|\|U^{\top*}_{2\perp}U_{2}^{(0)}\|\leq\left\|\mathcal{M}_{1}(Z)\cdot U^{*}_{2\perp}\otimes(U^{*}_{3}U^{\top*}_{3}+U^{*}_{3\perp}U^{\top*}_{3\perp})U_{3}^{(0)}\right\|\|U^{\top*}_{2\perp}U_{2}^{(0)}\|$
$\leq$ $\left\|\mathcal{M}_{1}(Z)\cdot U^{*}_{2\perp}\otimes(U^{*}_{3}U^{*\top}_{3}U_{3}^{(0)})\right\|\|U^{\top*}_{2\perp}U_{2}^{(0)}\|+\left\|\mathcal{M}_{1}(Z)\cdot U^{*}_{2}\otimes(U^{*}_{3\perp}U^{\top*}_{3\perp}U_{3}^{(0)})\right\|\|U^{\top*}_{2\perp}U_{2}^{(0)}\|$
$\leq$ $\left\|\mathcal{M}_{1}(Z)\cdot U^{*}_{2\perp}\otimes U^{*}_{3}\right\|\|U^{*\top}_{3}U_{3}^{(0)}\|\|U^{\top*}_{2\perp}U_{2}^{(0)}\|+\left\|\mathcal{M}_{1}(Z)\cdot U^{*}_{2}\otimes U^{*}_{3\perp}\right\|\|U^{\top*}_{3\perp}U_{3}^{(0)}\|\|U^{\top*}_{2\perp}U_{2}^{(0)}\|$
$\leq$ $C\kappa(\sqrt{p_{1}+p_{2}r_{3}})r_{\max}^{-1/2}+C\kappa(\sqrt{p_{1}+p_{3}r_{2}})r_{\max}^{-1}.$ (17)
Therefore (15), (16) and (17) leads to
$\|A-B\|\leq C_{4}\kappa(\sqrt{p_{1}}+\sqrt{r_{2}r_{3}}+\sqrt{p_{2}r_{3}}r_{\max}^{-1/2}+\sqrt{p_{3}r_{2}}r_{\max}^{-1/2}).$ (18)
In addition,
$\|B-B_{r_{1}}\|_{F}=$ $\|\mathcal{M}_{1}(X^{*})\cdot U_{2}^{(0)}\otimes U_{3}^{(0)}-\{\mathcal{M}_{1}(X^{*})\cdot U_{2}^{(0)}\otimes U_{3}^{(0)}\}_{r_{1}}\|_{\mathrm{F}}$
$\leq$ $\|\mathcal{M}_{1}(X^{*})\cdot U_{2}^{(0)}\otimes U_{3}^{(0)}-\{\mathcal{M}_{1}(X^{*})\}_{r_{1}}\cdot U_{2}^{(0)}\otimes U_{3}^{(0)}\|_{\mathrm{F}}$
$\leq$ $\|(\mathcal{M}_{1}(X^{*})-\{\mathcal{M}_{1}(X^{*})\}_{r_{1}})U_{2}^{(0)}\otimes U_{3}^{(0)}\|_{\mathrm{F}}$
$\leq$ $\|\mathcal{M}_{1}(X^{*})-\{\mathcal{M}_{1}(X^{*})\}_{r_{1}}\|_{\mathrm{F}}=\sqrt{\sum_{j=r_{1}+1}^{\operatorname*{rank}(\mathcal{M}_{1}(X^{*}))}\sigma_{j}^{2}(\mathcal{M}_{1}(X^{*}))}\leq\xi_{(r_{1},r_{2},r_{3})}.$ (19)
Here $\{\mathcal{M}_{1}(X^{*})\cdot U_{2}^{(0)}\otimes U_{3}^{(0)}\}_{r_{1}}$ indicate the best rank $r_{1}$ estimate of the matrix $\mathcal{M}_{1}(X^{*})\cdot U_{2}^{(0)}\otimes U_{3}^{(0)}$ in the first equality, and so for any rank $r_{1}$ matrix $\Phi$,
$\|\mathcal{M}_{1}(X^{*})\cdot U_{2}^{(0)}\otimes U_{3}^{(0)}-\{\mathcal{M}_{1}(X^{*})\cdot U_{2}^{(0)}\otimes U_{3}^{(0)}\}_{r_{1}}\|_{\mathrm{F}}\leq\|\mathcal{M}_{1}(X^{*})\cdot U_{2}^{(0)}\otimes U_{3}^{(0)}-\Phi\|_{\mathrm{F}};$
the second inequality holds because $\{\mathcal{M}_{1}(X^{*})\}_{r_{1}}\cdot U_{2}^{(0)}\otimes U_{3}^{(0)}$ is at most rank $r_{1}$, and the last inequality follows from Lemma 26.
It follows from (14), (18) and (19) that
$\left\|X^{*}\times_{1}(I_{p_{1}}-\mathcal{P}_{U_{1}^{(1)}})\times_{2}\mathcal{P}_{U_{2}^{(0)}}\times_{3}\mathcal{P}_{U_{3}^{(0)}}\right\|_{\mathrm{F}}\leq C_{5}\kappa(\sqrt{p_{1}r_{1}}+\sqrt{r_{1}r_{2}r_{3}}+\sqrt{p_{2}r_{2}}+\sqrt{p_{3}r_{3}})+C_{5}\xi_{(r_{1},r_{2},r_{3})}.$
Step 5. The conclusions of Step 3 and Step 4 lead to
$\left\|X^{*}\times_{1}(I_{p_{1}}-\mathcal{P}_{U_{1}^{(1)}})\times_{2}\mathcal{P}_{U_{2}^{*}}\times_{3}\mathcal{P}_{U_{3}^{*}}\right\|_{\mathrm{F}}\leq C_{6}\kappa(\sqrt{p_{1}r_{1}}+\sqrt{r_{1}r_{2}r_{3}}+\sqrt{p_{2}r_{2}}+\sqrt{p_{3}r_{3}})+C_{6}\xi_{(r_{1},r_{2},r_{3})}.$
This bound together with Step 1 and Step 2 leads to
$\left\|\widetilde{X}-X^{*}\right\|_{\mathrm{F}}\leq C_{7}\kappa(\sqrt{p_{1}r_{1}}+\sqrt{p_{2}r_{2}}+\sqrt{p_{3}r_{3}}+\sqrt{r_{1}r_{2}r_{3}})+C_{7}\xi_{(r_{1},r_{2},r_{3})}.$
Appendix B Deviation Bounds
Throughout this appendix we work with centered sub-Gaussian random variables with parameter $\kappa^{2}$, and we assume (without loss of generality) that each such variable $X$ satisfies $\mathbb{E}X^{2}=\kappa^{2}$. Any other case can be handled with additional constants in the bounds.
Lemma 6 (Theorem 4.4.3 in Vershynin, (2018)).
Assume all the entries of $Z\in\mathbb{R}^{m\times n}$ are independent mean-zero sub-Gaussian random variables, i.e.
$\|Z_{ij}\|_{\psi_{2}}=\sup_{q\geq 1}\mathbb{E}(|Z_{ij}|^{q})^{1/q}/q^{1/2}\leq\kappa.$
Then there exist some universal constant $C>0$, such that
$\|Z\|\leq C\kappa\left(\sqrt{m}+\sqrt{n}\right)$
with probability at least $1-\exp(-(m+n)))$.
Lemma 7.
Suppose all the entries of $Z\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}}$ are independent mean-zero sub-Gaussian random variables, i.e.
$\|Z_{ijk}\|_{\psi_{2}}=\sup_{q\geq 1}\mathbb{E}(|Z_{ijk}|^{q})^{1/q}/q^{1/2}\leq\kappa.$
Then there exist some universal constants $C,c>0$, such that
$\sup_{\begin{subarray}{c}A\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}},\\ \|A\|_{\mathrm{F}}\leq 1,A\in\mathcal{T}_{(r_{1},r_{2},r_{3})}\end{subarray}}\langle Z,A\rangle\leq C\kappa\left(r_{1}r_{2}r_{3}+\sum_{k=1}^{3}p_{k}r_{k}\right)^{1/2}$
with probability at least $1-\exp(-c\sum_{k=1}^{3}p_{k}r_{k})$.
Proof.
This directly follows from Lemma E.5 in Han et al., (2022). ∎
Lemma 8.
Suppose all the entries of $Z\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}}$ are independent mean-zero sub-Gaussian random variables, i.e.
$\|Z_{ijk}\|_{\psi_{2}}=\sup_{q\geq 1}\mathbb{E}(|Z_{ijk}|^{q})^{1/q}/q^{1/2}\leq\kappa.$
Let $W_{2}\in\mathbb{O}^{p_{2}\times s_{2}}$ and $W_{3}\in\mathbb{O}^{p_{3}\times s_{3}}$ be non-random. Then there exists absolute positive constants $C_{1},C_{2}$ and $c$ such that
$\mathbb{P}\bigg(\big\|\mathcal{M}_{1}(Z)\,(W_{2}\otimes W_{3})\big\|\geq C_{1}\kappa(\sqrt{p_{1}+s_{2}s_{3}})\bigg)\leq C_{2}\exp\left(-cp_{1}\right),$
Proof.
It suffices to observe that $W_{2}\otimes W_{3}\in\mathbb{O}^{p_{2}p_{3}\times r_{2}r_{3}}$. The desired result is a direct consequence of Lemma 10. ∎
Lemma 9.
Suppose all the entries of $Z\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}}$ are independent mean-zero sub-Gaussian random variables, i.e.
$\|Z_{ijk}\|_{\psi_{2}}=\sup_{q\geq 1}\mathbb{E}(|Z_{ijk}|^{q})^{1/q}/q^{1/2}\leq\kappa.$
Then there exists absolute positive constants $C_{1},C_{2}$ and $c$ such that
$\mathbb{P}\bigg(\sup_{\begin{subarray}{c}V_{2}\in\mathbb{R}^{p_{2}\times r_{2}},\,\|V_{2}\|\leq 1\\ V_{3}\in\mathbb{R}^{p_{3}\times r_{3}},\,\|V_{3}\|\leq 1\end{subarray}}\big\|\mathcal{M}_{1}(Z)\,(V_{2}\otimes V_{3})\big\|\geq C_{1}\kappa\Big(\sqrt{p_{1}+r_{2}r_{3}+p_{2}r_{2}+p_{3}r_{3}}\Bigg)$
$\leq C_{2}\exp\left(-c(p_{1}+p_{2}+p_{3})\right),$
Proof.
It follows from the assumption that $\left(\mathcal{M}_{1}(Z)\right)_{i,j}\overset{i.i.d}{\sim}\mathrm{subGaussian}(0,\kappa^{2})$.
Step 1. For fixed $V_{2}\in\mathbb{R}^{p_{2}\times r_{2}}$ with $\|V_{2}\|\leq 1$ and $V_{3}\in\mathbb{R}^{p_{3}\times r_{3}}$ with $\|V_{3}\|\leq 1$, it follows that
$\left\|V_{2}\otimes V_{3}\right\|=\left\|V_{2}\right\|\left\|V_{3}\right\|\leq 1,$
We upper bound $\|\mathcal{M}_{1}(Z)\cdot(V_{2}\otimes V_{3})\|$ for any fixed $V_{2}$ and $V_{3}$. Since,
$\mathcal{M}(Z)\in\mathbb{R}^{p_{1}\times p_{2}p_{3}},\left(\mathcal{M}(Z)\right)_{i,j}\overset{i.i.d}{\sim}\mathrm{subGaussian}(0,\kappa^{2}),\quad\text{and}\quad\left\|V_{2}\otimes V_{3}\right\|=\left\|V_{2}\right\|\left\|V_{3}\right\|=1.$
It follows from Lemma 10 that
$\mathbb{P}\Big(\left\|\mathcal{M}_{1}(Z)\cdot(V_{2}\otimes V_{3})\right\|>x\Big)\leq 2\exp\left(C(p_{1}+r_{2}r_{3})-cx^{2}\kappa^{-2}\right).$
Step 2. Let $\mathcal{N}_{p_{2},r_{2}}(\epsilon)$ denote an $\epsilon$ net of the set
$\{A\in\mathbb{R}^{p_{2}\times r_{2}}:\|A\|\leq 1\}$
with respect to the operator norm $\|\cdot\|$. Similarly, let Let $\mathcal{N}_{p_{3},r_{3}}(\epsilon)$ denote an $\epsilon$ net of the set
$\{B\in\mathbb{R}^{p_{3}\times r_{3}}:\|B\|\leq 1\}$
with respect to the operator norm $\|\cdot\|$. Denote the random quantity
$\psi=\sup_{\begin{subarray}{c}V_{2}\in\mathbb{R}^{p_{2}\times r_{2}},\,\|V_{2}\|\leq 1\\ V_{3}\in\mathbb{R}^{p_{3}\times r_{3}},\,\|V_{3}\|\leq 1\end{subarray}}\big\|\mathcal{M}_{1}(Z)\,(V_{2}\otimes V_{3})\big\|.$
For any given $V_{2}\in\mathbb{R}^{p_{2}\times r_{2}}$ and $V_{3}\in\mathbb{R}^{q}$ with $\|V_{2}\|\leq 1$ and $\|V_{3}\|\leq 1$, let $\widetilde{V}_{2}\in\mathcal{N}_{p_{2},r_{2}}(1/4)$ and $\widetilde{V}_{3}\in\mathcal{N}_{p_{3},r_{3}}(1/4)$ be such that
$\|V_{2}-\widetilde{V}_{2}\|\leq 1/4\quad\text{and}\quad\|V_{3}-\widetilde{V}_{3}\|\leq 1/4.$
Then
$\|\mathcal{M}_{1}(Z)V_{2}\otimes V_{3}\|\leq\|\mathcal{M}_{1}(Z)(V_{2}-\widetilde{V}_{2})\otimes V_{3}\|+\|\mathcal{M}_{1}(Z)\widetilde{V}_{2}\otimes(V_{3}-\widetilde{V}_{3})\|+\|\mathcal{M}_{1}(Z)\widetilde{V}_{2}\otimes\widetilde{V}_{3}\|.$
Note that $\|V_{2}-\widetilde{V}_{2}\|\leq 1/4$. So
$\|\mathcal{M}_{1}(Z)(V_{2}-\widetilde{V}_{2})\otimes V_{3}\|=\frac{1}{4}\|\mathcal{M}_{1}(Z)\{4(V_{2}-\widetilde{V}_{2})\}\otimes V_{3}\|\leq\frac{\psi}{4}.$
Similarly
$\|\mathcal{M}_{1}(Z)V_{2}\otimes(V_{3}-\widetilde{V}_{3})\|\leq\frac{\psi}{4}.$
In addition,
$\|\mathcal{M}_{1}(Z)\widetilde{V}_{2}\otimes\widetilde{V}_{3}\|\leq\sup_{V_{2}\in\mathcal{N}_{p_{2},r_{2}}(1/4),b\in\mathcal{N}_{p_{3},r_{3}}(1/4)}\|\mathcal{M}_{1}(Z)V_{2}\otimes V_{3}\|.$
So for any $V_{2}$ and $V_{3}$,
$\|\mathcal{M}_{1}(Z)V_{2}\otimes V_{3}\|\leq\frac{1}{2}\psi+\sup_{V_{2}\in\mathcal{N}_{p_{2},r_{2}}(1/4),b\in\mathcal{N}_{p_{3},r_{3}}(1/4)}\|\mathcal{M}_{1}(Z)V_{2}\otimes V_{3}\|.$
Taking sup over all $V_{2}\in\{A\in\mathbb{R}^{p_{2}\times r_{2}}:\|A\|\leq 1\}$ and $V_{3}\in\{B\in\mathbb{R}^{p_{3}\times r_{3}}:\|B\|\leq 1\}$, it follows that
$\psi\leq\frac{1}{2}\psi+\sup_{V_{2}\in\mathcal{N}_{p_{2},r_{2}}(1/4),b\in\mathcal{N}_{p_{3},r_{3}}(1/4)}\|\mathcal{M}_{1}(Z)V_{2}\otimes V_{3}\|,$
or simply
$\psi\leq 2\sup_{V_{2}\in\mathcal{N}_{p_{2},r_{2}}(1/4),b\in\mathcal{N}_{p_{3},r_{3}}(1/4)}\|\mathcal{M}_{1}(Z)V_{2}\otimes V_{3}\|.$
Step 3. By Proposition 8 in Pajor, (1998), the cardinality of $\mathcal{N}_{p_{2},r_{2}}(\epsilon)$ is bounded $(\frac{C}{\epsilon})^{p_{2}r_{2}}$, and $\mathcal{N}_{p_{3},r_{3}}(\epsilon)$ is bounded $(\frac{C}{\epsilon})^{p_{3}r_{3}}$. Therefore
$\mathbb{P}(\psi\geq 2t)\leq$ $\mathbb{P}\bigg(\sup_{V_{2}\in\mathcal{N}_{p_{2},r_{2}}(1/4),b\in\mathcal{N}_{p_{3},r_{3}}(1/4)}\|\mathcal{M}_{1}(Z)V_{2}\otimes V_{3}\|\geq t\bigg)$
$\leq$ $C_{2}^{p_{2}r_{2}}C_{3}^{p_{3}r_{3}}\sup_{V_{2}\in\mathcal{N}_{p_{2},r_{2}}(1/4),b\in\mathcal{N}_{p_{3},r_{3}}(1/4)}\mathbb{P}\bigg(\|\mathcal{M}_{1}(Z)V_{2}\otimes V_{3}\|\geq t\bigg)$
$\leq$ $2\exp\left(C(p_{1}+r_{2}r_{3}+p_{2}r_{2}+p_{3}r_{3})-ct^{2}\kappa^{-2}\right).$
Here $C$ and $C_{3}$ are positive constants. The desired result follows by noting
$\psi=\sup_{\begin{subarray}{c}V_{2}\in\mathbb{R}^{p_{2}\times r_{2}},\,\|V_{2}\|\leq 1\\ V_{3}\in\mathbb{R}^{p_{3}\times r_{3}},\,\|V_{3}\|\leq 1\end{subarray}}\big\|\mathcal{M}_{1}(Z)\,(V_{2}\otimes V_{3})\big\|.$
Lemma 10.
Suppose $Z\in\mathbb{R}^{n\times m}$, with $Z_{ij}\overset{i.i.d}{\sim}\mathrm{subGaussian}(0,\kappa^{2})$. Let $A\in\mathbb{R}^{p\times n}$ and $B\in\mathbb{R}^{m\times q}$ be non-random matrices. Then for any $t>0$
$\mathbb{P}\left(\left\|A\,Z\,B\right\|>t\right)\leq C_{1}\,\exp\left(C(p+q)-\frac{c\,t^{2}}{\kappa^{2}\,\left\|A\right\|^{2}\,\left\|B\right\|^{2}}\right).$ (20)
Proof.
Step 1. Let $u\in\mathbb{R}^{n}$ and $v\in\mathbb{R}^{m}$ be non-random. Then that $u^{\top}\,Z\,v=\sum_{i=1}^{n}\sum_{j=1}^{n}u_{i}Z_{ij}v_{j}$. Since $Z_{ij}$ are i.i.d. sub-Gaussian with parameter $\kappa^{2}$, it follows that $u^{\top}\,Z\,v$ is sub-Gaussian with parameter $\kappa^{2}\|u\|_{2}^{2}\|v\|_{2}^{2}$. Consequently by Hoeffding’s inequality,
$\mathbb{P}\left(\left|u^{\top}\,Z\,v\right|\,>t\,\right)\leq 2\,\exp\left(-\frac{c\,t^{2}}{\kappa^{2}\,\left\|u\right\|_{2}^{2}\,\left\|v\right\|_{2}^{2}}\right).$
Step 2. Let $a\in\mathbb{R}^{p}$ and $b\in\mathbb{R}^{q}$ be non-random vectors such that $\|a\|_{2},\|b\|_{2}\leq 1$. By Step 1, it follows that
$\mathbb{P}\left(\left|a^{\top}AZ\,Bb\right|\,>t\,\right)\leq 2\,\exp\left(-\frac{ct^{2}}{\kappa^{2}\,\left\|Aa\right\|_{2}^{2}\,\left\|Bb\right\|_{2}^{2}}\right)\leq 2\exp\left(-\frac{ct^{2}}{\kappa^{2}\,\left\|A\right\|^{2}\,\left\|B\right\|^{2}}\right).$
Step 3. Let $\mathcal{N}_{p}(\epsilon)$ be an $\epsilon$-net of the unit ball in $\mathbb{R}^{p}$. It follows that for any $a\in\mathbb{R}^{p}$ with $\|a\|_{2}=1$, there exists $\widetilde{a}\in\mathcal{N}_{p}(\epsilon)$ such that
$\|a-\widetilde{a}\|_{2}\leq\epsilon.$
Similarly let $\mathcal{N}_{q}(\epsilon)$ be an $\epsilon$-net of the unit ball in $\mathbb{R}^{q}$.
Denote the random quantity
$\psi=\|AZB\|=\sup_{a\in\mathbb{R}^{p},b\in\mathbb{R}^{q}\|a\|_{2}=\|b\|_{2}=1}|a^{\top}AZBb|.$
For any given $a\in\mathbb{R}^{p}$ and $b\in\mathbb{R}^{q}$, let $\widetilde{a}\in\mathcal{N}_{p}(1/4)$ and $\widetilde{b}\in\mathcal{N}_{q}(1/4)$ be such that
$\|a-\widetilde{a}\|_{2}\leq 1/4\quad\text{and}\quad\|b-\widetilde{b}\|_{2}\leq 1/4.$
Then
[…]
Note that $\|a-\widetilde{a}\|\leq\frac{1}{4}$. So
$|(a-\widetilde{a})^{\top}AZBb|=\frac{1}{4}|\{4(a-\widetilde{a})^{\top}\}AZBb|\leq\frac{\psi}{4}.$
Similarly
$|\widetilde{a}^{\top}AZB(\widetilde{b}-b)|\leq\frac{\psi}{4}.$
In addition,
$|\widetilde{a}^{\top}AZB\widetilde{b}|\leq\sup_{a\in\mathcal{N}_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}|a^{\top}AZBb|.$
So for any $a$ and $b$,
$|a^{\top}AZBb|\leq\frac{1}{2}\psi+\sup_{a\in N_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}|a^{\top}AZBb|.$
Taking sup over all unit vectors $a\in\mathbb{R}^{p}$ and $b\in\mathbb{R}^{q}$, it follows that
$\psi\leq\frac{1}{2}\psi+\sup_{a\in\mathcal{N}_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}|a^{\top}AZBb|,$
or simply
$\psi\leq 2\sup_{a\in\mathcal{N}_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}|a^{\top}AZBb|.$
Step 4. By Vershynin, (2018), the cardinality of $\mathcal{N}_{p}(\epsilon)$ is bounded by $(\frac{C}{\epsilon})^{p}$, and the cardinality of $\mathcal{N}_{q}(\epsilon)$ is bounded by $(\frac{C}{\epsilon})^{q}$. Therefore
$\mathbb{P}(\psi\geq 2t)=\mathbb{P}\bigg(\sup_{a\in\mathcal{N}_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}|a^{\top}AZBb|\geq t\bigg)\leq$ ${C_{2}}^{p}{C_{2}}^{q}\sup_{a\in\mathcal{N}_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}\mathbb{P}(|a^{\top}AZBb|\geq t)$
$\leq$ $2{C_{2}}^{p}{C_{2}}^{q}\exp\left(-\frac{ct^{2}}{\kappa^{2}\,\left\|A\right\|^{2}\,\left\|B\right\|^{2}}\right),$
where $c,C$ are positive constants. The desired result follows by noting $\psi=\|AZB\|$. ∎
Lemma 11.
Suppose $Z\in\mathbb{R}^{n\times m}$, with $Z_{ij}\overset{i.i.d}{\sim}\mathrm{subGaussian}(0,\kappa^{2})$. Let $A\in\mathbb{R}^{m\times p}$ and $B\in\mathbb{R}^{m\times q}$ be non-random matrices. Then for any $t>0$
$\mathbb{P}\left(\left\|A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B\right\|>t\right)\leq C_{1}\,\exp\left(C(p+q)-\min\left(\frac{t^{2}}{n\,\kappa^{4}\,\left\|B\right\|^{2}\,\left\|A\right\|^{2}},\,\frac{t}{\kappa^{2}\left\|B\right\|\,\left\|A\right\|}\right)\right),$ (21)
where $C$ and $C_{1}$ are positive constants.
Proof.
For any non-random $u,v\in\mathbb{R}^{m}$ , it follows that
$u^{\top}Z^{\top}Zv-n\kappa^{2}u^{\top}v=\sum_{j=1}^{n}(u^{\top}Z_{j})(v^{\top}Z_{j})-\mathbb{E}\{(u^{\top}Z_{j})(v^{\top}Z_{j})\},$
where $Z_{j}$ is the $j$-th row of $Z$. Note that $(u^{\top}Z_{j})$ is sub-Gaussian with parameter $\kappa^{2}\|u\|_{2}^{2}$, and $(v^{\top}Z_{j})$ is sub-Gaussian with parameter $\kappa^{2}\|v\|_{2}^{2}$. Since $Z$ have i.i.d. entries, it follows that $\{(u^{\top}Z_{j})(v^{\top}Z_{j})\}_{j=1}^{n}$ are i.i.d. sub-exponential with parameter $\kappa^{4}\|u\|_{2}^{2}\|v\|_{2}^{2}$. So
$\mathbb{P}\bigg(|u^{\top}Z^{\top}Zv-n\kappa^{2}u^{\top}v|\geq t\bigg)\leq 2\exp\bigg(-c\min\bigg\{\frac{t^{2}}{n\kappa^{4}\|u\|_{2}^{2}\|v\|_{2}^{2}},\frac{t}{\kappa^{2}\|u\|_{2}\|v\|_{2}}\bigg\}\bigg).$
Step 1. Let $\mathcal{N}_{p}(\epsilon)$ be the $\epsilon$-net of the unit ball in $\mathbb{R}^{p}$. It follows that for any $a\in\mathbb{R}^{p}$ with $\|a\|_{2}=1$, there exists $\widetilde{a}\in\mathcal{N}_{p}(\epsilon)$ such that
$\|a-\widetilde{a}\|_{2}\leq\epsilon.$
Similarly let $\mathcal{N}_{q}(\epsilon)$ be the $\epsilon$-net of the unit ball in $\mathbb{R}^{q}$.
Denote the random quantity
$\psi=\|A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B\|=\sup_{a\in\mathbb{R}^{p},b\in\mathbb{R}^{q}\|a\|_{2}=\|b\|_{2}=1}\bigl|a^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b\bigr|.$
For any given $a\in\mathbb{R}^{p}$ and $b\in\mathbb{R}^{q}$, let $\widetilde{a}\in\mathcal{N}_{p}(1/4)$ and $\widetilde{b}\in\mathcal{N}_{q}(1/4)$ be such that
$\|a-\widetilde{a}\|_{2}\leq 1/4\quad\text{and}\quad\|b-\widetilde{b}\|_{2}\leq 1/4.$
Then
$|a^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b|\leq|(a-\widetilde{a})^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b|$
[…]
Note that $\|a-\widetilde{a}\|_{2}\leq\frac{1}{4}$. So
$|(a-\widetilde{a})^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b|=\frac{1}{4}|\{4(a-\widetilde{a})^{\top}\}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b|\leq\frac{\psi}{4}.$
Similarly
$|\widetilde{a}^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)(\widetilde{b}-b)|\leq\frac{\psi}{4}.$
In addition,
$|\widetilde{a}^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)\widetilde{b}|\leq\sup_{a\in\mathcal{N}_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}|a^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b|.$
So for any $a$ and $b$,
$|a^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b|\leq\frac{1}{2}\psi+\sup_{a\in N_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}|a^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b|.$
Taking sup over all unit vectors $a\in\mathbb{R}^{p}$ and $b\in\mathbb{R}^{q}$, it follows that
$\psi\leq\frac{1}{2}\psi+\sup_{a\in\mathcal{N}_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}|a^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b|,$
or simply
$\psi\leq 2\sup_{a\in\mathcal{N}_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}|a^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b|.$
Step 2. By Vershynin, (2018), the cardinality of $\mathcal{N}_{p}(\epsilon)$ is bounded by $(\frac{C}{\epsilon})^{p}$, and the cardinality of $\mathcal{N}_{q}(\epsilon)$ is bounded by $(\frac{C}{\epsilon})^{q}$, for a positive constant $C$. Therefore
$\mathbb{P}(\psi\geq 2t)=$ $\mathbb{P}\bigg(\sup_{a\in\mathcal{N}_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}|a^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b|\geq t\bigg)$
$\leq$ ${C_{2}}^{p}{C_{2}}^{q}\sup_{a\in\mathcal{N}_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}\mathbb{P}(|a^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b|\geq t)$
$=$ ${C_{2}}^{p}{C_{2}}^{q}\sup_{a\in\mathcal{N}_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}\mathbb{P}(|(Aa)^{\top}Z^{\top}Z(Bb)-n\kappa^{2}(Aa)^{\top}(Bb)|\geq t)$
$\leq$ $2{C_{2}}^{p}{C_{2}}^{q}\exp\bigg(-c\min\bigg\{\frac{t^{2}}{n\kappa^{4}\|Aa\|_{2}^{2}\|Bb\|_{2}^{2}},\frac{t}{\kappa^{2}\|Aa\|_{2}\|Bb\|_{2}}\bigg\}\bigg),$
where $C_{2}$ is a positive constant. The desired result follows from the observation that $\|Aa\|_{2}\leq\|A\|\|a\|_{2}\leq\|A\|$, and $\|Bb\|_{2}\leq\|B\|\|b\|_{2}\leq\|B\|$. ∎
Lemma 12.
Suppose $Z\in\mathbb{R}^{m\times n}$ is a sub-Gaussian random matrix in the sense that for any $u\in\mathbb{R}^{m},v\in\mathbb{R}^{n}$, it holds that
$\|u^{\top}Zv\|_{\psi_{2}}\leq\kappa\|u\|_{2}\|v\|_{2}.$
Then with probability at least $1-\exp(-c(m+n))$, it holds that
$\mathbb{P}\left(\left\|Z\right\|>t\right)\leq C_{1}\,\exp\left(C(m+n)-\frac{c\,t^{2}}{\kappa^{2}\,}\right),$
where $c,C$ and $C_{1}$ are positive constants.
Proof.
By assumption,
$\mathbb{P}\left(\left|u^{\top}\,Z\,v\right|\,>t\,\right)\leq 2\,\exp\left(-\frac{c\,t^{2}}{\kappa^{2}\,\left\|u\right\|_{2}^{2}\,\left\|v\right\|_{2}^{2}}\right).$
The rest of the proof is similar and simpler than Lemma 10 and is omitted.
B.1 SVD for Unbalanced Matrices
Lemma 13.
Suppose
$Y=X+Z\in\mathbb{R}^{n\times m},$
where $X$ is a non-random matrix of arbitrary rank, and $Z$ is a random matrix whose entries are i.i.d. sub-Gaussian random variables with mean zero and the sub-Gaussian norm $\|Z_{ij}\|_{\psi_{2}}=\kappa<\infty$. For any $r\leq\min\{n,m\}$, write the full SVD of $X$ as
$X=U\Sigma V^{\top}=\begin{bmatrix}U_{r}\ U_{\perp}\end{bmatrix}\cdot\begin{bmatrix}\Sigma_{r}&\\ &\Sigma_{\perp}\end{bmatrix}\cdot\begin{bmatrix}V_{r}^{\top}\\ V_{\perp}^{\top}\end{bmatrix}=X_{r}+X_{\perp}.$
Here $U_{r}\in\mathbb{O}_{m,r}$, $V_{r}\in\mathbb{O}_{m,r}$ correspond to the leading $r$ left and right singular vectors of $X$. Suppose that
$\{\sigma_{r}(X)-\sigma_{r+1}(X)\}^{2}\geq C_{\mathrm{gap}}\kappa^{2}\{\sqrt{mn}+m\}$
where $C_{\mathrm{gap}}>0$ is a sufficient large constant. Then with probability at least $1-C_{1}\exp(-C_{2}n)$, it holds that
$\left\|\sin\Theta\left(\widehat{V}_{r},V_{r}\right)\right\|^{2}\leq C_{3}\bigg\{\frac{m\kappa^{2}}{(\sigma_{r}(X)-\sigma_{r+1}(X))^{2}}+\frac{\kappa^{4}nm}{(\sigma_{r}(X)-\sigma_{r+1}(X))^{4}}\bigg\},$
where $C_{1},C_{2},C_{3}>0$ are absolute constants only depending on $C_{\mathrm{gap}}$.
Proof.
Note that by assumption, we have
$\mathbb{E}[Z^{\top}Z]=n\kappa^{2}I_{m},\quad\mathbb{E}[Y^{\top}Y]=V_{r}\Sigma_{r}^{2}V_{r}^{\top}+V_{\perp}\Sigma_{\perp}^{2}V_{\perp}^{\top}+n\kappa^{2}I_{m},\quad\mathbb{E}[V_{r}^{\top}Y^{\top}YV_{r}]=\Sigma_{r}^{2}+n\kappa^{2}I_{r}.$
Define the diagonal weighting matrix
$M=\operatorname*{diag}\left((\sigma_{1}^{2}+n\kappa^{2})^{-1/2},\dots,(\sigma_{r}^{2}+n\kappa^{2})^{-1/2}\right)\in\mathbb{R}^{r\times r}.$
Then it holds that
$YV_{r}M$ $=\left(X_{r}+X_{\perp}+Z\right)V_{r}M=\left(X_{r}+Z\right)V_{r}M,$
$M^{\top}\,\mathbb{E}\left[V_{r}^{\top}Y^{\top}YV_{r}\right]\,M$ $=M^{\top}\,\mathbb{E}\left[V_{r}^{\top}\left(X_{r}+Z\right)^{\top}\left(X_{r}+Z\right)V_{r}\right]\,M=I_{r}.$
Step 1. For $\sigma_{r}(YV_{r})$, observe that
$\sigma_{r}^{2}(YV_{r})=$ $\sigma_{r}^{2}\left(\{X_{r}+Z\}V_{r}\right)=\sigma_{r}^{2}\left(\{X_{r}+Z\}V_{r}MM^{-1}\right)\geq\sigma_{r}^{2}(\{X_{r}+Z\}V_{r}M)\sigma_{\min}^{2}(M^{-1})$
$=$ $\sigma_{r}^{2}\left(\{X_{r}+Z\}V_{r}M\right)\{\sigma^{2}_{r}(X)+n\kappa^{2}\}$
$=$ $\sigma_{r}\big(M^{\top}V_{r}^{\top}\{X_{r}+Z\}^{\top}\{X_{r}+Z\}V_{r}M\big)\{\sigma^{2}_{r}(X)+n\kappa^{2}\}$ (22)
where the inequality follows from Lemma 17.
Consider the term $\sigma_{r}\big(M^{\top}V_{r}^{\top}\{X_{r}+Z\}^{\top}\{X_{r}+Z\}V_{r}M\big)$. Note that
$M^{\top}V_{r}^{\top}\left(X_{r}+Z\right)^{\top}\left(X_{r}+Z\right)V_{r}M-I_{r}$
$=\,$ $M^{\top}V_{r}^{\top}\left(X_{r}+Z\right)^{\top}\left(X_{r}+Z\right)V_{r}M-\mathbb{E}\left[M^{\top}V_{r}^{\top}\left(X_{r}+Z\right)^{\top}\left(X_{r}+Z\right)V_{r}M\right]$
$=\,$ $\underbrace{M^{\top}V_{r}^{\top}X_{r}^{\top}X_{r}V_{r}M-\mathbb{E}\left[M^{\top}V_{r}^{\top}X_{r}^{\top}X_{r}V_{r}M\right]}_{{=0}}\,+\,M^{\top}V_{r}^{\top}X_{r}^{\top}ZV_{r}M$
$\,+\,M^{\top}V_{r}^{\top}Z^{\top}X_{r}V_{r}M\,+\,M^{\top}V_{r}^{\top}Z^{\top}ZV_{r}M-\mathbb{E}\left[M^{\top}V_{r}^{\top}Z^{\top}ZV_{r}M\right]$
$\,+\,\underbrace{\mathbb{E}\left[M^{\top}V_{r}^{\top}X_{r}^{\top}ZV_{r}M\right]}_{=0}\,+\,\underbrace{\mathbb{E}\left[M^{\top}V_{r}^{\top}Z^{\top}X_{r}V_{r}M\right]}_{=0}$
$=\,$ $M^{\top}V_{r}^{\top}X_{r}^{\top}ZV_{r}M\,+\,M^{\top}V_{r}^{\top}Z^{\top}X_{r}V_{r}M\,+\,M^{\top}V_{r}^{\top}\left(Z^{\top}Z-n\kappa^{2}I_{m}\right)V_{r}M.$ (23)
Since,
$\left\|X_{r}V_{r}M\right\|^{2}=\max_{k=1,\ldots,r}\frac{\sigma_{k}^{2}(X)}{\sigma_{k}^{2}(X)+n\kappa^{2}}\leq 1,\qquad\text{and}\qquad\left\|V_{r}M\right\|^{2}=\left\|M\right\|^{2}=\frac{1}{\sigma_{r}^{2}(X)+n\kappa^{2}},$
it follows by Lemma 10 that
$\mathbb{P}\Big(\left\|M^{\top}V_{r}^{\top}X_{r}^{\top}ZV_{r}M\right\|\geq x\Big)\leq 2\,\exp\left(Cr-cx^{2}\,\frac{\sigma_{r}^{2}(X)+n\kappa^{2}}{\kappa^{2}}\right).$ (24)
Similarly, Lemma 11 implies that
$\begin{split}&\mathbb{P}\Big(\left\|M^{\top}V_{r}^{\top}\left(Z^{\top}Z-n\kappa^{2}I_{m}\right)V_{r}M\right\|\geq x\Big)\\ &\leq 2\,\exp\left(Cr-c\min\left\{x^{2}\,\frac{\left(\sigma_{r}^{2}(X)+n\kappa^{2}\right)^{2}}{n\kappa^{4}},\,x\,\frac{\sigma_{r}^{2}(X)+n\kappa^{2}}{\kappa^{2}}\right\}\right),\\ &\leq 2\,\exp\left(Cr-c\,\frac{\sigma_{r}^{2}(X)+n\kappa^{2}}{\kappa^{2}}\min\left\{x^{2},\,x\right\}\right),\end{split}$ (25)
where the last inequality follows from the fact that $\frac{\{\sigma_{r}^{2}(X)+n{\kappa}^{2}\}^{2}}{\kappa^{4}n}\geq\frac{\sigma_{r}^{2}(X)+n{\kappa}^{2}}{\kappa^{2}}$. Thus, combining (23), (24), and (25), we have
$\mathbb{P}\!\left(\big\|M^{\top}V_{r}^{\top}(X_{r}+Z)^{\top}(X_{r}+Z)V_{r}M-I_{r}\big\|\geq x\right)\leq 6\exp\!\left(Cr-c\frac{\sigma_{r}^{2}(X)+n\kappa^{2}}{\kappa^{2}}\,\min\{x^{2},x\}\right).$ (26)
Here we apply a union bound to the three terms in (23) with thresholds $x/3$ each. The same tail bound as in (24) holds for $M^{\top}V_{r}^{\top}Z^{\top}X_{r}V_{r}M$ by symmetry, and (25) controls the centered quadratic term $M^{\top}V_{r}^{\top}(Z^{\top}Z-n\kappa^{2}I_{m})V_{r}M$. All absolute constants are absorbed into $C,c$. This implies that
$\mathbb{P}\bigg(\sigma_{r}\big(M^{\top}V_{r}^{\top}\{X_{r}+Z\}^{\top}\{X_{r}+Z\}V_{r}M\big)\geq 1-x\bigg)$
$\geq 1-6\exp\left(Cr-c\,\frac{\sigma_{r}^{2}(X)+n\kappa^{2}}{\kappa^{2}}\min\left\{x^{2},\,x\right\}\right),$ (27)
(27) and (22) together imply that
$\mathbb{P}\Big(\sigma_{r}^{2}(YV_{r})\geq\{\sigma_{r}^{2}(X)+n\kappa^{2}\}(1-x)\Big)\geq 1-6\,\exp\left(Cr-c\,\frac{\sigma_{r}^{2}(X)+n\kappa^{2}}{\kappa^{2}}\min\left\{x^{2},\,x\right\}\right).$
Setting $x=\tfrac{1}{6}\,\tfrac{\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X)}{\sigma_{r}^{2}(X)+n\kappa^{2}}$, we have
$\mathbb{P}\left(\sigma_{r}^{2}(YV_{r})\geq\sigma_{r}^{2}(X)+n\kappa^{2}-\frac{\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X)}{6}\right)$
$\geq 1-6\exp\left(Cr-c\min\left\{\frac{1}{36\kappa^{2}}\,\frac{\left(\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X)\right)^{2}}{\sigma^{2}_{r}(X)+n\kappa^{2}},\,\frac{\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X)}{6\kappa^{2}}\right\}\right).$ (28)
Step 2. We upper bound the term $\sigma_{r+1}^{2}(Y)$. Note that
$\sigma_{r+1}(Y)=\min_{\operatorname*{rank}(B)\leq r}\|Y-B\|\leq\left\|Y-Y\cdot V_{r}V_{r}^{\top}\right\|=\sigma_{\max}(YV_{\perp}).$
Moreover,
$\sigma_{\max}^{2}(YV_{\perp})$ $=\left\|V_{\perp}^{\top}Y^{\top}YV_{\perp}\right\|=\left\|V_{\perp}^{\top}(X_{r}+X_{\perp}+Z)^{\top}(X_{r}+X_{\perp}+Z)V_{\perp}\right\|$
$=\left\|V_{\perp}^{\top}(X_{\perp}+Z)^{\top}(X_{\perp}+Z)V_{\perp}\right\|$
$\leq\left\|V_{\perp}^{\top}Z^{\top}ZV_{\perp}\right\|+\left\|V_{\perp}^{\top}Z^{\top}X_{\perp}V_{\perp}\right\|+\left\|V_{\perp}^{\top}X_{\perp}^{\top}ZV_{\perp}\right\|+\left\|V_{\perp}^{\top}X_{\perp}^{\top}X_{\perp}V_{\perp}\right\|$
$=\underbrace{\left\|V_{\perp}^{\top}Z^{\top}ZV_{\perp}\right\|}_{\mathrm{I}_{1}}+2\underbrace{\left\|V_{\perp}^{\top}Z^{\top}X_{\perp}V_{\perp}\right\|}_{\mathrm{I}_{2}}+\underbrace{\left\|V_{\perp}^{\top}X_{\perp}^{\top}X_{\perp}V_{\perp}\right\|}_{\mathrm{I}_{3}}.$ (29)
For the term $\mathrm{I}_{3}$, we have
$\mathrm{I}_{3}=\sigma^{2}_{1}(X_{\perp}V_{\perp})=\sigma^{2}_{1}(U_{\perp}\Sigma_{\perp}V_{\perp}^{\top}V_{\perp})=\sigma^{2}_{1}(\Sigma_{\perp})=\sigma^{2}_{r+1}(X).$ (30)
For $\mathrm{I}_{2}$, note that
$\left\|X_{\perp}V_{\perp}\right\|^{2}=\left\|U_{\perp}\Sigma_{\perp}\right\|^{2}=\sigma_{r+1}^{2}(X),\qquad\text{and}\qquad\left\|V_{\perp}\right\|^{2}=1.$
it follows from Lemma 10 that
$\mathbb{P}\Big(\mathrm{I}_{2}\geq x\Big)\leq 2\,\exp\left(C_{1}m-\frac{c_{1}x^{2}}{\kappa^{2}\,\sigma_{r+1}^{2}(X)}\right).$ (31)
For $\mathrm{I}_{1}$, note that by Lemma 11 and the fac that $V_{\perp}^{\top}V_{\perp}=I_{m-r}$, we have
$\begin{split}&\mathbb{P}\Big(\left\|V_{\perp}^{\top}Z^{\top}ZV_{\perp}-n\kappa^{2}I_{m-r}\right\|\geq t\Big)\leq 2\,\exp\left(C_{2}m-c_{2}\min\left\{\frac{t^{2}}{n\,\kappa^{4}},\,\frac{t}{\kappa^{2}}\right\}\right),\\ \implies&\mathbb{P}\Big(\mathrm{I}_{1}\geq n\kappa^{2}(1+t)\Big)\leq 2\,\exp\left(C_{2}m-c_{2}n\min\left\{t^{2},\,t\right\}\right).\end{split}$ (32)
Combining the calculations in this step, with $x=\kappa\sigma_{r+1}(X)\sqrt{\tfrac{2C_{1}}{c_{1}}m}$ in (31), and with $t=\tfrac{\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X)}{6n\kappa^{2}}$ in (32), it follows that
$\begin{split}&\mathbb{P}\left(\sigma^{2}_{r+1}(Y)\geq n\kappa^{2}+\frac{\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X)}{6}+\kappa\sigma_{r+1}(X)\sqrt{\frac{2C_{1}}{c_{1}}m\,}+\sigma_{r+1}^{2}(X)\right)\\ &\quad\leq 2\,\exp\left(C_{2}m-c_{2}\,\min\!\left\{\,\frac{\left(\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X)\right)^{2}}{36\,\kappa^{4}\,n},\frac{\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X)}{6\,\kappa^{2}}\right\}\right)+2\,\exp\bigl(-C_{3}\,m\bigr).\end{split}$ (33)
Step 3. Recall we define
$M=\operatorname*{diag}\left((\sigma_{1}^{2}+n\kappa^{2})^{-1/2},\dots,(\sigma_{r}^{2}+n\kappa^{2})^{-1/2}\right)\in\mathbb{R}^{r\times r}.$
We have
$\left\|\mathcal{P}_{YV_{r}}YV_{\perp}\right\|$ $=\left\|\mathcal{P}_{YV_{r}M}YV_{\perp}\right\|$
$=\left\|(YV_{r}M)\left((YV_{r}M)^{\top}(YV_{r}M)\right)^{-1}(YV_{r}M)^{\top}YV_{\perp}\right\|$
$\leq\left\|(YV_{r}M)\left((YV_{r}M)^{\top}(YV_{r}M)\right)^{-1}\right\|\left\|M^{\top}V_{r}^{\top}Y^{\top}YV_{\perp}\right\|$
$\leq\sigma_{\min}^{-1}(YV_{r}M)\left\|M^{\top}V_{r}^{\top}Y^{\top}YV_{\perp}\right\|=\sigma_{r}^{-1}(YV_{r}M)\left\|M^{\top}V_{r}^{\top}Y^{\top}YV_{\perp}\right\|,$ (34)
where the first equality follows from the fact that $YV_{r}$ and $YV_{r}M$ have the same column spaces (since $M$ is invertible), the last inequality follows from Lemma 15, and for the last equality we use that the singular values of $YV_{r}M$ are in nonincreasing order so that its smallest singular value equals $\sigma_{r}(YV_{r}M)$.
By (26), we have for every $x>0$
$\mathbb{P}\Bigl(\bigl\|M^{\top}V_{r}^{\top}(X_{r}+Z)^{\top}(X_{r}+Z)V_{r}M-I_{r}\bigr\|\geq x\Bigr)\leq 6\exp\!\left(Cr-c\,\frac{\sigma_{r}^{2}(X)+n\kappa^{2}}{\kappa^{2}}\min\{x^{2},x\}\right).$
Taking $x=\frac{1}{2}$ gives
$\mathbb{P}\Bigl(\bigl\|M^{\top}V_{r}^{\top}(X_{r}+Z)^{\top}(X_{r}+Z)V_{r}M-I_{r}\bigr\|<\tfrac{1}{2}\Bigr)\geq 1-6\exp\!\left(Cr-c\,\frac{\sigma_{r}^{2}(X)+n\kappa^{2}}{4\kappa^{2}}\right).$ (35)
[…]
$M^{\top}V_{r}^{\top}Y^{\top}YV_{\perp}$ $=M^{\top}V_{r}^{\top}\left(X_{r}+X_{\perp}+Z\right)^{\top}\left(X_{r}+X_{\perp}+Z\right)V_{\perp}$
$=M^{\top}V_{r}^{\top}X_{r}^{\top}ZV_{\perp}+M^{\top}V_{r}^{\top}Z^{\top}X_{\perp}V_{\perp}+M^{\top}V_{r}^{\top}Z^{\top}ZV_{\perp}$
$=M^{\top}V_{r}^{\top}X_{r}^{\top}ZV_{\perp}+M^{\top}V_{r}^{\top}Z^{\top}X_{\perp}V_{\perp}+M^{\top}V_{r}^{\top}Z^{\top}ZV_{\perp}-\underbrace{M^{\top}V_{r}^{\top}\left(n\kappa^{2}I_{m}\right)V_{\perp}}_{=0},$
Since,
$\left\|X_{r}V_{r}M\right\|^{2}\leq 1,\,\left\|V_{r}M\right\|^{2}=\frac{1}{\sigma_{r}^{2}(X)+n\kappa^{2}}\text{ and }\left\|X_{\perp}V_{\perp}\right\|^{2}=\left\|X_{\perp}\right\|^{2}=\sigma^{2}_{r+1}(X),$
it follows from Lemma 10 that
$\begin{split}\mathbb{P}\Big(\left\|M^{\top}V_{r}^{\top}X_{r}^{\top}ZV_{\perp}\right\|\geq x\Big)&\leq 2\,\exp\left(Cm-\frac{cx^{2}}{\kappa^{2}}\right),\\ \mathbb{P}\Big(\left\|M^{\top}V_{r}^{\top}Z^{\top}X_{\perp}V_{\perp}\right\|\geq x\Big)&\leq 2\,\exp\left(Cm-\frac{cx^{2}}{\kappa^{2}}\,\frac{\sigma_{r}^{2}(X)+n\kappa^{2}}{\sigma_{r+1}^{2}(X)}\right)\leq 2\,\exp\left(Cm-\frac{cx^{2}}{\kappa^{2}}\right).\end{split}$ (36)
Similarly, Lemma 11 implies that
$\begin{split}&\mathbb{P}\Big(\left\|M^{\top}V_{r}^{\top}\left(Z^{\top}Z-n\kappa^{2}I_{m}\right)V_{\perp}\right\|\geq x\Big)\\ &\leq 2\,\exp\left(Cm-c\min\left\{x^{2}\,\frac{\sigma_{r}^{2}(X)+n\kappa^{2}}{n\kappa^{4}},\,x\,\frac{\sqrt{\sigma_{r}^{2}(X)+n\kappa^{2}}}{\kappa^{2}}\right\}\right),\\ &\leq 2\,\exp\left(Cm-c\min\left\{\frac{x^{2}}{\kappa^{2}},\,x\,\frac{\sqrt{\sigma_{r}^{2}(X)+n\kappa^{2}}}{\kappa^{2}}\right\}\right),\end{split}$ (37)
where the last inequality uses $\frac{\sigma_{R}^{2}(X)+n{\kappa}^{2}}{n\,\kappa^{2}}\geq 1$. Thus, combining (B.1), (35), (36) and (37) with $x=\kappa\sqrt{\frac{2C}{c}m\,}$, we have
$\mathbb{P}\!\left(\left\|\mathcal{P}_{YV_{r}}YV_{\perp}\right\|^{2}\geq\frac{36C}{c}\,m\kappa^{2}\right)$ $\leq 6\exp\!\left(Cr-c\frac{\sigma_{r}^{2}(X)+n\kappa^{2}}{4\kappa^{2}}\right)+4\exp(-Cm)$ (38)
$\quad+2\exp\!\left(Cm-c\min\!\left\{\tfrac{2C}{c}\,m,\,\sqrt{\tfrac{2C}{c}\,m\;\tfrac{\sigma_{r}^{2}(X)+n\kappa^{2}}{\kappa^{2}}}\right\}\right).$
Step 4. Define the event
$\mathcal{E}=\Bigg\{$ $\sigma_{r}^{2}(YV_{r})\geq\sigma_{r}^{2}(X)+n\kappa^{2}-\frac{\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X)}{6}\,;$
$\sigma_{r+1}^{2}(Y)\leq n\kappa^{2}+\frac{\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X)}{6}+\kappa\sigma_{r+1}(X)\sqrt{\frac{2C}{c}m\,}+\sigma_{r+1}^{2}(X)\,;$
$\left\|\mathcal{P}_{YV_{r}}YV_{\perp}\right\|^{2}\leq\frac{8C}{c}\,m\kappa^{2}\Bigg\}.$
It follows from (B.1), (33) and (38) that by the union bound,
$\mathbb{P}(\mathcal{E}^{c})$ $\leq 6\exp\left(Cr-c\min\left\{\frac{1}{36\kappa^{2}}\,\frac{\left(\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X)\right)^{2}}{\sigma^{2}_{r}(X)+n\kappa^{2}},\,\frac{\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X)}{6\kappa^{2}}\right\}\right)$ (39)
$\quad+2\,\exp\left(Cm\;-\;c\,\min\!\left\{\,\frac{\left(\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X)\right)^{2}}{36\,\kappa^{4}\,n},\frac{\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X)}{6\,\kappa^{2}}\right\}\right)+2\,\exp\bigl(-C\,m\bigr)$ (40)
$\quad+6\exp\left(Cr-c\frac{\sigma_{r}^{2}(X)+n\kappa^{2}}{4\kappa^{2}}\right)+4\exp(-Cm)$ (41)
$\quad+2\exp\left(Cm-c\min\left\{\frac{2C}{c}\,m,\,\sqrt{\frac{2C}{c}\,m\frac{\sigma_{r}^{2}(X)+n\kappa^{2}}{\kappa^{2}}}\right\}\right).$ (42)
In what follows, we show that under the SNR assumption
$\Big(\sigma_{r}(X)-\sigma_{r+1}(X)\Big)^{2}\geq C_{\mathrm{gap}}\kappa^{2}\Big(\sqrt{nm}+m\Big)$
with sufficient large absolute constant $C_{\mathrm{gap}}>0$, we have
$\mathbb{P}(\mathcal{E}^{c})$ $\leq C\exp\left(-Cm\right),$
where $C>0$ is an absolute constant, appropriately scaled to absorb the other constants.
We illustrate how to bound (39), as the rest of the terms can be bounded in a similar and simpler way. Note that
$\frac{1}{36\kappa^{2}}\,\frac{\left(\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X)\right)^{2}}{\sigma^{2}_{r}(X)+n\kappa^{2}}$ $\geq\frac{\left(\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X)\right)^{2}}{72\kappa^{2}}\min\left\{\frac{1}{\sigma_{r}^{2}(X)},\,\frac{1}{n\kappa^{2}}\right\}$
$\geq\min\left\{\frac{\left(\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X)\right)^{2}}{72\,\kappa^{2}\sigma_{r}^{2}(X)},\,\frac{\left(\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X)\right)^{2}}{72n\kappa^{4}}\right\}.$
We have
$\frac{\big(\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X)\big)^{2}}{72\,\kappa^{2}\sigma_{r}^{2}(X)}=\frac{\big(\sigma_{r}(X)+\sigma_{r+1}(X)\big)^{2}\,\big(\sigma_{r}(X)-\sigma_{r+1}(X)\big)^{2}}{72\,\kappa^{2}\sigma_{r}^{2}(X)}\geq\frac{\big(\sigma_{r}(X)-\sigma_{r+1}(X)\big)^{2}}{72\,\kappa^{2}}\geq\frac{2C}{c}m,$
$\frac{\left(\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X)\right)^{2}}{72\,\kappa^{4}\,n}\geq\frac{\left(\sigma_{r}(X)-\sigma_{r+1}(X)\right)^{4}}{72\,\kappa^{4}\,n}\geq\frac{2C}{c}m.$
So
$\frac{1}{36\kappa^{2}}\,\frac{\left(\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X)\right)^{2}}{\sigma^{2}_{r}(X)+n\kappa^{2}}\geq\frac{2C}{c}m.$
In addition,
$\frac{\sigma_{R}^{2}(X)-\sigma_{R+1}^{2}(X)}{6\kappa^{2}}\geq\frac{\big(\sigma_{R}^{2}(X)-\sigma_{R+1}^{2}(X)\big)^{2}}{72\,\kappa^{2}\sigma_{R}^{2}(X)}\geq\frac{2C}{c}m.$
[…]
Step 5. Under the event $\mathcal{E}$, by Lemma 21, we have that
$\left\|\sin\Theta\left(\widehat{V}_{r},V_{r}\right)\right\|^{2}$ $\leq\frac{\sigma_{r}^{2}(YV_{r})\left\|\mathcal{P}_{YV_{r}}YV_{\perp}\right\|^{2}}{\left(\sigma_{r}^{2}(YV_{r})-\sigma_{r+1}^{2}(Y)\right)^{2}}\leq C_{7}\frac{\sigma_{r}^{2}(YV_{r})m\kappa^{2}}{\left(\sigma_{r}^{2}(YV_{r})-\sigma_{r+1}^{2}(Y)\right)^{2}}$
$\leq C_{8}\frac{\left(\sigma_{r}^{2}(X)+n\kappa^{2}-\frac{\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X)}{6}\right)m\kappa^{2}}{\left(\sigma_{r}^{2}(X)+n\kappa^{2}-\frac{\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X)}{6}-\sigma_{r+1}^{2}(Y)\right)^{2}}$
$\leq C_{8}\frac{\left(\sigma_{r}^{2}(X)+n\kappa^{2}-\frac{\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X)}{6}\right)m\kappa^{2}}{\left((1-\frac{1}{3})\left(\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X)\right)-\kappa\sigma_{r+1}(X)\sqrt{\frac{2C}{c}m\,}\right)^{2}}$
$\leq C_{9}\frac{\left(\sigma_{r}^{2}(X)+n\kappa^{2}-\frac{\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X)}{6}\right)m\kappa^{2}}{(1-\frac{1}{2})^{2}\big(\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X)\big)^{2}}$
$\leq C_{10}\frac{\left(\sigma_{r}^{2}(X)+n\kappa^{2}-\frac{\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X)}{6}\right)m\kappa^{2}}{\big(\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X)\big)^{2}},$
Here, the third inequality follows from the fact that $x^{2}/(x^{2}-y^{2})^{2}$ is a decreasing function of $x$ and an increasing function of $y$ when $x>y\geq 0$, together with the fact that the event $\mathcal{E}$ holds. The fifth inequality follows from the fact that, under the assumption $\Big(\sigma_{r}(X)-\sigma_{r+1}(X)\Big)^{2}\geq C_{\mathrm{gap}}\kappa^{2}(\sqrt{nm}+m)$ with $C_{\mathrm{gap}}>0$ being large enough,
$\frac{\Big(\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X)\Big)^{2}}{36}=$ $\Big(\sigma_{r}(X)+\sigma_{r+1}(X)\Big)^{2}\frac{\Big(\sigma_{r}(X)-\sigma_{r+1}(X)\Big)^{2}}{36}$
$\geq$ $\sigma_{r}^{2}(X)\frac{C_{\mathrm{gap}}m\kappa^{2}}{36}\geq Cm\kappa^{2}\sigma_{r}^{2}(X)\geq Cm\kappa^{2}\sigma_{r+1}^{2}(X).$
Therefore, with probability at least $1-C\exp(-Cm)$,
$\left\|\sin\Theta\left(\widehat{V}_{r},V_{r}\right)\right\|^{2}$ $\leq C_{3}\frac{\left(\sigma_{r}^{2}(X)+n\kappa^{2}-\frac{\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X)}{6}\right)m\kappa^{2}}{(\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X))^{2}}$
$\leq C_{3}\frac{\left(\sigma_{r}^{2}(X)+n\kappa^{2}\right)m\kappa^{2}}{(\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X))^{2}}=C_{3}\frac{\sigma_{r}^{2}(X)m\kappa^{2}}{(\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X))^{2}}+C_{3}\frac{nm\kappa^{4}}{(\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X))^{2}}$
$\leq C_{4}\bigg\{\frac{m\kappa^{2}}{(\sigma_{r}(X)-\sigma_{r+1}(X))^{2}}+\frac{\kappa^{4}nm}{(\sigma_{r}(X)-\sigma_{r+1}(X))^{4}}\bigg\},$
where the third inequality follows from the observation that
$\frac{\sigma_{r}^{2}(X)}{(\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X))^{2}}=\frac{\sigma_{r}^{2}(X)}{(\sigma_{r}(X)+\sigma_{r+1}(X))^{2}(\sigma_{r}(X)-\sigma_{r+1}(X))^{2}}\leq\frac{1}{(\sigma_{r}(X)-\sigma_{r+1}(X))^{2}},$
and last display follows from
$\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X)=(\sigma_{r}(X)+\sigma_{r+1}(X))(\sigma_{r}(X)-\sigma_{r+1}(X))\geq\Big(\sigma_{r}(X)-\sigma_{r+1}(X)\Big)^{2}.$
Corollary 14.
Suppose the conditions of Theorem 1 hold, in particular the condition in Equation (2) for each mode-$k$ unfolding $\mathcal{M}_{k}(X^{*})$ with target rank $r_{k}$. Let $U_{k}^{(0)}$ be the matrix of top $r_{k}$ left singular vectors of the mode-$k$ unfolding $Y^{(k)}:=\mathcal{M}_{k}(Y)$. Then for each $k\in\{1,2,3\}$, with probability at least $1-C_{1}\exp(-C_{2}p_{k})$,
$\bigl\|\sin\Theta\!\left(U_{k}^{(0)},\,U_{k}^{*}\right)\bigr\|\;\leq\;\frac{1}{2\sqrt{r_{\max}}}\,.$
Proof.
Apply Lemma 13 to $Y^{(k)}=\mathcal{M}_{k}(Y)\in\mathbb{R}^{p_{k}\times p_{\neg k}}$ with signal $X^{(k)}=\mathcal{M}_{k}(X^{*})$ and noise $Z^{(k)}=\mathcal{M}_{k}(Z)$, where $p_{\neg k}:=\prod_{j\neq k}p_{j}$. Use rank $r=r_{k}$ and identify $n:=p_{k}$ (rows) and $m:=p_{\neg k}$ (columns). The lemma (applied to left singular vectors, or equivalently to the transpose) gives
$\bigl\|\sin\Theta(U_{k}^{(0)},U_{k}^{*})\bigr\|^{2}\;\leq\;C\left\{\frac{p_{k}\,\kappa^{2}}{\Delta_{k}^{2}}+\frac{\kappa^{4}\,p_{k}\,p_{\neg k}}{\Delta_{k}^{4}}\right\},$
where $\Delta_{k}:=\sigma_{r_{k}}\!\bigl(\mathcal{M}_{k}(X^{*})\bigr)-\sigma_{r_{k}+1}\!\bigl(\mathcal{M}_{k}(X^{*})\bigr)$ is the mode-$k$ spectral gap.
By the assumption (2) in Theorem 1 (applied to mode $k$),
$\Delta_{k}^{2}\;\geq\;C_{\mathrm{gap}}\kappa^{2}\!\left(\sqrt{p_{k}\,p_{\neg k}\,r_{\max}}+r_{\max}\sum_{j=1}^{3}p_{j}\right).$
Choosing $C_{\mathrm{gap}}$ sufficiently large makes the right–hand side above at most $1/(4r_{\max})$, hence
$\bigl\|\sin\Theta(U_{k}^{(0)},U_{k}^{*})\bigr\|\;\leq\;\frac{1}{2\sqrt{r_{\max}}}\,.$
The probability bound $1-C_{1}\exp(-C_{2}p_{k})$ matches the row dimension in the matrix lemma. ∎
Appendix C Matrix Perturbation Bounds
Lemma 15.
Suppose that $A\in\mathbb{R}^{n\times r}$. Then
$\|A(A^{\top}A)^{-1}\|\leq\sigma_{r}^{-1}(A).$
Proof.
If $\sigma_{r}(A)=0$, then the desired result trivially follows. So suppose $\operatorname*{rank}(A)=r$. Therefore $A^{\top}A$ is invertible. Let the SVD of $A$ satisfies $A=U_{A}\Sigma_{A}V_{A}^{\top}$, then
$\left\|A(A^{\top}A)^{-1}\right\|=\left\|U_{A}\Sigma_{A}V_{A}^{\top}(V_{A}\Sigma_{A}^{2}V_{A}^{\top})^{-1}\right\|=\left\|U_{A}\Sigma_{A}^{-1}V_{A}^{\top}\right\|=\sigma_{\min}^{-1}(A)=\sigma_{r}^{-1}(A).$
Lemma 16.
Suppose $A\in\mathbb{R}^{m\times n}$ and $B\in\mathbb{R}^{n\times k}$ are any two matrices. Then
$\sigma_{j}(AB)\leq\sigma_{j}(A)\sigma_{\max}(B).$
Proof.
Let $\lambda_{j}(M)$ denote the jthj-th eigenvalues of $M$ in the absolute value order. Then
$\sigma_{j}(AB)=\sqrt{\lambda_{j}(ABB^{\top}A^{\top})}\quad\text{and}\quad\sigma_{j}(A)=\sqrt{\lambda_{j}(AA^{\top})}.$ (43)
Since $BB^{\top}\preceq\sigma^{2}_{\max}(B)I_{n}$, it follows that
$ABB^{\top}A^{\top}\preceq A(\sigma^{2}_{\max}(B)I_{n})A^{\top}=\sigma^{2}_{\max}(B)AA^{\top}.$
By the monotonicity of eigenvalues under the positive definite matrices, it follows that
$\lambda_{j}(ABB^{\top}A^{\top})\leq\sigma_{\max}^{2}(B)\lambda_{j}(AA^{\top}).$ (44)
The desired result follows from (43). ∎
Lemma 17.
Suppose $A\in\mathbb{R}^{m\times n}$ and $B\in\mathbb{R}^{n\times n}$ are any two matrices. Then
$\sigma_{j}(AB)\geq\sigma_{j}(A)\sigma_{\min}(B).$
Proof.
Suppose $\sigma_{\min}(B)=0$. Then the desired result immediately follows. Therefore it suffices to assume $\sigma_{\min}(B)>0$ and $B$ is invertible. It suffices to observe that
$\sigma_{j}(A)=\sigma_{j}(ABB^{-1})\leq\sigma_{j}(AB)\sigma_{\max}(B^{-1})=\sigma_{j}(AB)\sigma^{-1}_{\min}(B),$
where the inequality follows from Lemma 16.
Lemma 18.
For any real matrices $A\in\mathbb{R}^{n\times m}$ and $B\in\mathbb{R}^{m\times m}$, it holds that
$\|AB\|_{\mathrm{F}}\geq\sigma_{\min}(B)\|A\|_{\mathrm{F}}.$
Proof of Lemma 18.
Since $B$ is a square matrix, it follows that
$\lambda_{\min}(BB^{\top})=\sigma_{\min}^{2}(B),$
where $\lambda_{\min}(\cdot)$ denotes the minimum eigenvalue. Note that
$BB^{\top}\succeq\lambda_{\min}(BB^{\top})I_{m}.$
Therefore
$ABB^{\top}A^{\top}\succeq A\{\lambda_{\min}(BB^{\top})I_{m}\}A^{\top},$
and so
$\operatorname*{tr}(ABB^{\top}A^{\top})\geq\operatorname*{tr}(A\{\lambda_{\min}(BB^{\top})I_{m}\}A^{\top}).$
Then
$\|AB\|_{\mathrm{F}}^{2}=\operatorname*{tr}(ABB^{\top}A^{\top})\geq\operatorname*{tr}(A\{\lambda_{\min}(BB^{\top})I_{m}\}A^{\top})=\lambda_{\min}(BB^{\top})\operatorname*{tr}(AA^{\top})=\sigma_{\min}^{2}(B)\|A\|^{2}_{\mathrm{F}}.$
Lemma 19.
Let $A\in\mathbb{R}^{p\times q}$ and $U\in\mathbb{O}^{q\times r}$. Then
$\|AUU^{\top}\|_{\mathrm{F}}=\|AU\|_{\mathrm{F}}.$
Proof.
Observe that
$\|AUU^{\top}\|_{\mathrm{F}}^{2}=\operatorname*{tr}(AUU^{\top}UU^{\top}A^{\top})=\operatorname*{tr}(AUU^{\top}A^{\top})=\|AU\|_{\mathrm{F}}^{2}.$
Lemma 20.
Suppose $B,Z\in\mathbb{R}^{n\times m}$. For all $1\leq R\leq\min\{n,m\}$, write the full SVD of $A$ as
$A=B+Z=\widehat{U}\widehat{\Sigma}\widehat{V}^{\top}=\begin{bmatrix}\widehat{U}_{(R)}\ \widehat{U}_{\perp}\end{bmatrix}\cdot\begin{bmatrix}\widehat{\Sigma}_{(R)}&\\ &\widehat{\Sigma}_{\perp}\end{bmatrix}\cdot\begin{bmatrix}\widehat{V}_{(R)}^{\top}\\ \widehat{V}_{\perp}^{\top}\end{bmatrix}$,
where $\widehat{U}_{(R)}\in\mathbb{O}_{n,R}$, $\widehat{V}_{(R)}\in\mathbb{O}_{m,R}$ correspond to the leading $R$ left and right singular vectors; and $\widehat{U}_{\perp}\in\mathbb{O}_{n,n-R}$, $\widehat{V}_{\perp}\in\mathbb{O}_{m,m-R}$ correspond to their orthonormal complement. We have
$\left\|\mathcal{P}_{\widehat{U}_{\perp}}B\right\|_{\mathrm{F}}\leq$ $3\sqrt{\sum_{j=R+1}^{\min\{n,m\}}\sigma_{j}^{2}(B)}+2\min\left\{\sqrt{R}\|Z\|,\|Z\|_{\mathrm{F}}\right\}$
$=$ $3\left\|B_{(R)}-B\right\|_{\mathrm{F}}+2\min\left\{\sqrt{R}\|Z\|,\|Z\|_{\mathrm{F}}\right\},$
where $B_{(R)}$ denote the rank-$R$ truncated SVD of $B$, this is, if $B=U\Sigma V^{\top}$ then $B_{(R)}:=U_{(R)}\Sigma_{(R)}V_{(R)}^{\top}$.
Proof.
Without loss of generality, assume $n\leq m$. For $A\in\mathbb{R}^{n\times m}$, let $\Sigma(A)\in\mathbb{R}^{n\times m}$ denote the non-negative diagonal matrices whose diagonal entries are the non-increasingly ordered singular values of $A$. For all $1\leq R\leq n$, let $B_{(R)}$ denote the truncated SVD of $B$ with rank $R$, and we have by the Eckart–Young–Mirsky theorem
$\left\|B_{(R)}-B\right\|_{\mathrm{F}}=\sqrt{\sum_{j=R+1}^{n}\sigma_{j}^{2}(B)}.$
For a matrix $A\in\mathbb{R}^{m\times n}$, let $\Sigma(A)\in\mathbb{R}^{m\times n}$ be a non-negative (rectangular) diagonal matrix whose diagonal entries are the non-increasingly ordered singular values of $A$.
We have that
$\left\|\mathcal{P}_{\widehat{U}_{\perp}}B\right\|_{\mathrm{F}}\leq\left\|\mathcal{P}_{\widehat{U}_{\perp}}B_{(R)}\right\|_{\mathrm{F}}+\left\|\mathcal{P}_{\widehat{U}_{\perp}}(B-B_{(R)})\right\|_{\mathrm{F}}=\sqrt{\sum_{j=1}^{R}\sigma_{j}^{2}(\mathcal{P}_{\widehat{U}_{\perp}}B_{(R)})}+\left\|\mathcal{P}_{\widehat{U}_{\perp}}(B-B_{(R)})\right\|_{\mathrm{F}}$
$\leq$ $\sqrt{\sum_{j=1}^{R}\sigma_{j}^{2}(\mathcal{P}_{\widehat{U}_{\perp}}B_{(R)})}+\left\|B-B_{(R)}\right\|_{\mathrm{F}}=\sqrt{\sum_{j=1}^{R}\sigma_{j}^{2}(\mathcal{P}_{\widehat{U}_{\perp}}B_{(R)})}+\sqrt{\sum_{j=R+1}^{n}\sigma_{j}^{2}(B)}$
$\leq$ $\left\|(\sigma_{1}(\mathcal{P}_{\widehat{U}_{\perp}}B_{(R)})-\sigma_{1}(\mathcal{P}_{\widehat{U}_{\perp}}B),\dots,\sigma_{R}(\mathcal{P}_{\widehat{U}_{\perp}}B_{(R)})-\sigma_{R}(\mathcal{P}_{\widehat{U}_{\perp}}B))^{\top}\right\|_{2}+\left\|(\sigma_{1}(\mathcal{P}_{\widehat{U}_{\perp}}B),\dots,\sigma_{R}(\mathcal{P}_{\widehat{U}_{\perp}}B))^{\top}\right\|_{2}$
$+\sqrt{\sum_{j=R+1}^{n}\sigma_{j}^{2}(B)}$
$\leq$ $\left\|\Sigma(\mathcal{P}_{\widehat{U}_{\perp}}B_{(R)})-\Sigma(\mathcal{P}_{\widehat{U}_{\perp}}B)\right\|_{\mathrm{F}}+\left\|(\sigma_{1}(\mathcal{P}_{\widehat{U}_{\perp}}B),\dots,\sigma_{R}(\mathcal{P}_{\widehat{U}_{\perp}}B))^{\top}\right\|_{2}+\sqrt{\sum_{j=R+1}^{n}\sigma_{j}^{2}(B)}$
$\leq$ $\left\|\mathcal{P}_{\widehat{U}_{\perp}}(B_{(R)}-B)\right\|_{\mathrm{F}}+\sqrt{\sum_{j=1}^{R}\sigma_{j}^{2}(\mathcal{P}_{\widehat{U}_{\perp}}B)}+\sqrt{\sum_{j=R+1}^{n}\sigma_{j}^{2}(B)}$
$\leq$ $\sqrt{\sum_{j=1}^{R}\sigma_{j}^{2}(\mathcal{P}_{\widehat{U}_{\perp}}B)}+2\sqrt{\sum_{j=R+1}^{n}\sigma_{j}^{2}(B)},$
where the first equality follows from $\operatorname*{rank}(B_{(R)})=R$, and the fifth inequality follows from Theorem 23. To upper bound $\sqrt{\sum_{j=1}^{R}\sigma_{j}^{2}(\mathcal{P}_{\widehat{U}_{\perp}}B)}$, we first consider $\sqrt{\sum_{j=1}^{R}\sigma_{j}^{2}(\mathcal{P}_{\widehat{U}_{\perp}}A)}$. Note that
$\mathcal{P}_{\widehat{U}_{\perp}}A=\sum_{j=R+1}^{n}\sigma_{j}(A)\widehat{u}_{j}\widehat{v}_{j}^{\top},$
where $\widehat{u}_{j}$ and $\widehat{v}_{j}$ are the left and right singular vectors associated with the $j$th largest singular value $\sigma_{j}(A)$. Note that $\sigma_{j}(A)=\sigma_{j}(B)=0$ for $j>n$. It follows that
$\sqrt{\sum_{j=1}^{R}\sigma_{j}^{2}(\mathcal{P}_{\widehat{U}_{\perp}}A)}=\sqrt{\sum_{j=R+1}^{2R}\sigma_{j}^{2}(A)}=\left\|(\sigma_{R+1}(A),\dots,\sigma_{2R}(A))^{\top}\right\|$
$\leq$ $\left\|(\sigma_{R+1}(A)-\sigma_{R+1}(B),\dots,\sigma_{2R}(A)-\sigma_{2R}(B))^{\top}\right\|+\left\|(\sigma_{R+1}(B),\dots,\sigma_{2R}(B))^{\top}\right\|$
$\leq$ $\min\left\{\sqrt{R}\|Z\|,\|Z\|_{\mathrm{F}}\right\}+\sqrt{\sum_{j=R+1}^{n}\sigma_{j}^{2}(B)},$ (45)
where the first inequality follows from the triangle inequality, and second inequality follows from Weyl’s inequality (Weyl,, 1912), i.e. $|\sigma_{j}(A)-\sigma_{j}(B)|\leq\|A-B\|$ for all $1\leq j\leq n$, as well as the fact that
$\left\|(\sigma_{R+1}(A)-\sigma_{R+1}(B),\dots,\sigma_{2R}(A)-\sigma_{2R}(B))^{\top}\right\|\leq\left\|\Sigma(A)-\Sigma(B)\right\|_{\mathrm{F}}\leq\left\|Z\right\|_{\mathrm{F}},$
where the last inequality follows from Theorem 23. It then follows from (C),
$\sqrt{\sum_{j=1}^{R}\sigma_{j}^{2}(\mathcal{P}_{\widehat{U}_{\perp}}B)}=\left\|(\sigma_{1}(\mathcal{P}_{\widehat{U}_{\perp}}(A-Z)),\dots,\sigma_{R}(\mathcal{P}_{\widehat{U}_{\perp}}(A-Z))^{\top}\right\|$
$\leq$ $\left\|(\sigma_{1}(\mathcal{P}_{\widehat{U}_{\perp}}(A-Z))-\sigma_{1}(\mathcal{P}_{\widehat{U}_{\perp}}A),\dots,\sigma_{R}(\mathcal{P}_{\widehat{U}_{\perp}}(A-Z))-\sigma_{R}(\mathcal{P}_{\widehat{U}_{\perp}}A))^{\top}\right\|$
$+\left\|(\sigma_{1}(\mathcal{P}_{\widehat{U}_{\perp}}A),\dots,\sigma_{R}(\mathcal{P}_{\widehat{U}_{\perp}}A))^{\top}\right\|$
$\leq$ $\min\left\{\sqrt{R}\|\mathcal{P}_{\widehat{U}_{\perp}}Z\|,\|\mathcal{P}_{\widehat{U}_{\perp}}Z\|_{\mathrm{F}}\right\}+\sqrt{\sum_{j=1}^{R}\sigma_{j}^{2}(\mathcal{P}_{\widehat{U}_{\perp}}A)}$
$\leq$ $\min\left\{\sqrt{R}\|Z\|,\|Z\|_{\mathrm{F}}\right\}+\sqrt{\sum_{j=1}^{R}\sigma_{j}^{2}(P_{\widehat{U}_{\perp}}A)}$
$\leq$ $2\min\left\{\sqrt{R}\|Z\|,\|Z\|_{\mathrm{F}}\right\}+\sqrt{\sum_{j=R+1}^{n}\sigma_{j}^{2}(B)},$
where the first two inequalities follow from the same arguments as in (C). Consequently,
$\left\|\mathcal{P}_{\widehat{U}_{\perp}}B\right\|_{\mathrm{F}}\leq 3\sqrt{\sum_{j=R+1}^{n}\sigma_{j}^{2}(B)}+2\min\left\{\sqrt{R}\|Z\|,\|Z\|_{\mathrm{F}}\right\}.$
Lemma 21 (Proposition 1 of Cai and Zhang, (2018)).
Suppose $Y\in\mathbb{R}^{m\times n}$, $\widehat{V}=[\widehat{V}_{r}\ \widehat{V}_{\perp}]\in\mathbb{O}_{n}$ where $\widehat{V}_{r}\in\mathbb{O}_{n,r}$, $\widehat{V}_{\perp}\in\mathbb{O}_{n,n-r}$ correspond to the first $r$ and last (nr)(n-r) right singular vectors of $Y$ respectively. Let $V=[V_{r}\ V_{\perp}]\in\mathbb{O}_{n,n}$ be any orthogonal matrix with $V_{r}\in\mathbb{O}_{n,r}$, $V_{\perp}\in\mathbb{O}_{n,n-r}$. Given that $\sigma_{R}(YV_{r})>\sigma_{r+1}(Y)$, we have
$\|\sin\Theta(V_{r},\widehat{V}_{r})\|\leq\frac{\sigma_{r}(YV_{r})\|\mathcal{P}_{YV_{r}}YV_{\perp}\|}{\sigma_{r}^{2}(YV_{r})-\sigma_{r+1}^{2}(Y)}\wedge 1,$ (46)
where $\mathcal{P}_{A}$ is the projection operator onto the column space of $A$.
[…]
The following properties hold for the $\sin\Theta$ distances.
  1. 1.
    (Equivalent Expressions) Suppose $V,\widehat{V}\in\mathbb{O}_{p,R}$. If $V_{\perp}$ is an orthogonal extension of $V$, namely $\begin{bmatrix}V&V_{\perp}\end{bmatrix}\in\mathbb{O}_{p}$, we have the following equivalent forms for $\|\sin\Theta(\widehat{V},V)\|$ and $\|\sin\Theta(\widehat{V},V)\|_{\mathrm{F}}$,
    $\|\sin\Theta(\widehat{V},V)\|=\sqrt{1-\sigma_{\min}^{2}(\widehat{V}^{T}V)}=\|\widehat{V}^{T}V_{\perp}\|,$
    $\|\sin\Theta(\widehat{V},V)\|_{\mathrm{F}}=\sqrt{r-\|V^{T}\widehat{V}\|_{\mathrm{F}}^{2}}=\|\widehat{V}^{T}V_{\perp}\|_{\mathrm{F}}.$
  2. 2.
    (Triangle Inequality) For all $V_{1},V_{2},V_{3}\in\mathbb{O}_{p,R}$,
    $\|\sin\Theta(V_{2},V_{3})\|\leq\|\sin\Theta(V_{1},V_{2})\|+\|\sin\Theta(V_{1},V_{3})\|,$
    $\|\sin\Theta(V_{2},V_{3})\|_{\mathrm{F}}\leq\|\sin\Theta(V_{1},V_{2})\|_{\mathrm{F}}+\|\sin\Theta(V_{1},V_{3})\|_{\mathrm{F}}.$
  3. 3.
    (Equivalence with Other Metrics)
    $\|\sin\Theta(\widehat{V},V)\|\leq\sqrt{2}\|\sin\Theta(\widehat{V},V)\|,$
    $\|\sin\Theta(\widehat{V},V)\|_{\mathrm{F}}\leq\sqrt{2}\|\sin\Theta(\widehat{V},V)\|_{\mathrm{F}},$
    $\|\sin\Theta(\widehat{V},V)\|\leq\|\widehat{V}\widehat{V}^{\top}-VV^{\top}\|\leq 2\|\sin\Theta(\widehat{V},V)\|,$
    $\|\widehat{V}\widehat{V}^{\top}-VV^{\top}\|_{\mathrm{F}}=\sqrt{2}\|\sin\Theta(\widehat{V},V)\|_{\mathrm{F}}.$
Theorem 23 (Mirsky’s singular value inequality in Mirsky, (1960)).
For any matrices $A,B\in\mathbb{R}^{m\times n}$, let $A=V_{1}\Sigma(A)W_{1}^{\top}$ and $B=V_{2}\Sigma(B)W_{2}^{\top}$ be the full SVDs of $A$ and $B$, respectively. Note that $\Sigma(A),\Sigma(B)\in\mathbb{R}^{m\times n}$ are non-negative (rectangular) diagonal matrices whose diagonal entries are the non-increasingly ordered singular values of $A$ and $B$, respectively. Then
$\|\Sigma(A)-\Sigma(B)\|\leq\|A-B\|$ (47)
for any unitarily invariant norm $\|\cdot\|$ on $\mathbb{R}^{m\times n}$.
Theorem 24 (Weyl’s Inequality for Singular Values ).
Let $A,B\in\mathbb{R}^{m\times n}$ and denote their singular values (in nonincreasing order) by $\{\sigma_{i}(A)\}$ and $\{\sigma_{i}(B)\}$ respectively. In addition denote the singular values of $A+B$ as $\{\sigma_{i}(A+B)\}$. Then for all indices $i,j$ satisfying $i+j-1\leq\min\{m,n\}$,
$\sigma_{i+j-1}(A+B)\leq\sigma_{i}(A)+\sigma_{j}(B).$
Lemma 25 (Ky Fan-type Inequality for Sums of Matrices).
Let $A,B\in\mathbb{R}^{m\times n}$ and denote their singular values (in nonincreasing order) by $\{\sigma_{i}(A)\}$ and $\{\sigma_{i}(B)\}$ respectively. In addition denote the singular values of $A+B$ as $\{\sigma_{i}(A+B)\}$. Then for any $1\leq k\leq\min\{m,n\}$, it holds that
$\sqrt{\sum_{i=1}^{k}\sigma_{i}^{2}(A+B)}\leq\sqrt{\sum_{i=1}^{k}\sigma_{i}^{2}(A)}+\sqrt{\sum_{i=1}^{k}\sigma_{i}^{2}(B)}.$
Proof.
For a symmetric matrix $M\in\mathbb{R}^{n\times n}$, by Ky Fan’s maximum principle (see e.g. II.1.13 in Bhatia, (2013)), for any $1\leq k\leq n$,
$\sum_{i=1}^{k}\lambda_{i}(M)=\sup_{P\in\mathbb{O}^{n\times k}}\operatorname*{tr}(P^{\top}MP).$
Therefore
$\sum_{i=1}^{k}\sigma_{i}^{2}(A)=\sum_{i=1}^{k}\lambda_{i}(A^{\top}A)=\sup_{P\in\mathbb{O}^{n\times k}}\operatorname*{tr}(P^{\top}A^{\top}AP)=\sup_{P\in\mathbb{O}^{n\times k}}\|AP\|_{\mathrm{F}}^{2},$
and so
$\sqrt{\sum_{i=1}^{k}\sigma_{i}^{2}(A)}=\sup_{P\in\mathbb{O}^{n\times k}}\|AP\|_{\mathrm{F}}.$
Then
$\sqrt{\sum_{i=1}^{k}\sigma_{i}^{2}\left(A+B\right)}=$ $\max_{U\in\mathbb{O}_{n,k}}\|(A+B)U\|_{\mathrm{F}}\leq\max_{U\in\mathbb{O}_{n,k}}\|AU\|_{\mathrm{F}}+\max_{U\in\mathbb{O}_{n,k}}\|BU\|_{\mathrm{F}}$
$=\sqrt{\sum_{i=1}^{k}\sigma_{i}^{2}\left(A\right)}+\sqrt{\sum_{i=1}^{k}\sigma_{i}^{2}\left(B\right)}.$
Let $\mathcal{T}_{(r_{1},r_{2},r_{3})}$ denote the class of tensor in $\mathbb{R}^{p_{1}\times p_{2}\times p_{3}}$ with tucker ranks at most $(r_{1},r_{2},r_{3})$. More precisely
$\mathcal{T}_{(r_{1},r_{2},r_{3})}=\{A\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}}:\operatorname*{rank}(\mathcal{M}_{k}(A))\leq r_{k},k=1,2,3\}.$
Lemma 26.
Let $X^{*}\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}}$. For $k\in\{1,2,3\}$, suppose the $k$-th matricization of $X^{*}$ satisfies
$\mathcal{M}_{k}(X^{*})=\begin{bmatrix}U_{k}^{*}&U_{k\,\perp}^{*}\end{bmatrix}\begin{bmatrix}\Sigma_{k}^{*}&0\\ 0&\Sigma_{k\,\perp}^{*}\end{bmatrix}\begin{bmatrix}V_{k}^{*}&V_{k\,\perp}^{*}\end{bmatrix}^{\top}$
where $U_{k}^{*}\in\mathbb{O}_{p_{k},r_{k}}$ corresponds to the the top $r_{k}$ singular vectors of $\mathcal{M}_{k}(X^{*})$. Then for $k\in\{1,2,3\}$, it holds that
$\left\|X^{*}\times_{k}U_{k\perp}^{*}\right\|_{\mathrm{F}}=\left\|X^{*}\times_{k}(I_{p_{k}}-\mathcal{P}_{U_{k}^{*}})\right\|_{\mathrm{F}}=\sqrt{\sum_{j=r_{k}+1}^{\operatorname*{rank}(\mathcal{M}_{k}(X^{*}))}\sigma^{2}_{j}(\mathcal{M}_{k}(X^{*}))}\leq\xi_{(r_{1},r_{2},r_{3})},$
where
$\xi_{(r_{1},r_{2},r_{3})}=\inf_{A\in\mathcal{T}_{(r_{1},r_{2},r_{3})}}\|A-X^{*}\|_{\mathrm{F}}.$
Proof.
By symmetry, it suffices to consider $k=1$. Note that
$\left\|X^{*}\times_{k}(I_{p_{k}}-\mathcal{P}_{U_{k}^{*}})\right\|_{\mathrm{F}}=\left\|X^{*}\times_{k}\mathcal{P}_{U_{k\perp^{*}}}\right\|_{\mathrm{F}}=\left\|X^{*}\times_{k}U_{k\perp}^{*}\right\|_{\mathrm{F}}.$
In addition
$\left\|X^{*}\times_{1}(I_{p_{1}}-\mathcal{P}_{U_{1}^{*}})\right\|_{\mathrm{F}}=$ $\left\|(I_{p_{1}}-\mathcal{P}_{U_{1}^{*}})\cdot\mathcal{M}_{1}(X^{*})\right\|_{\mathrm{F}}=\left\|U_{1\,\perp}^{*}U_{1\,\perp}^{*\,\top}\cdot(U_{1}^{*}\Sigma^{*}_{1}V_{1}^{*\,\top}+U_{1\,\perp}^{*}\Sigma^{*}_{1\,\perp}V_{1\,\perp}^{*\,\top})\right\|_{\mathrm{F}}$
$=$ $\left\|U_{1\,\perp}^{*}U_{1\,\perp}^{*\,\top}\cdot(U_{1\,\perp}^{*}\Sigma^{*}_{1\,\perp}V_{1\,\perp}^{*\,\top})\right\|_{\mathrm{F}}=\sqrt{\sum_{j=r_{1}+1}^{\operatorname*{rank}(\mathcal{M}_{1}(X^{*}))}\sigma^{2}_{j}(\mathcal{M}_{1}(X^{*}))}.$
Note that by the properties of SVD, for any $W\in\mathbb{R}^{p_{1}\times p_{2}p_{3}}$ such that $\operatorname*{rank}(W)\leq r_{1}$, it holds that
$\sqrt{\sum_{j=r_{1}+1}^{\operatorname*{rank}(\mathcal{M}_{1}(X^{*}))}\sigma^{2}_{j}(\mathcal{M}_{1}(X^{*}))}\leq\|\mathcal{M}_{1}(X^{*})-W\|_{\mathrm{F}}.$
For any $A\in\mathcal{T}_{(r_{1},r_{2},r_{3})}$, it holds that $\operatorname*{rank}(\mathcal{M}_{1}(A))\leq r_{1}$. Therefore for any $A\in\mathcal{T}_{(r_{1},r_{2},r_{3})}$,
$\sqrt{\sum_{j=r_{1}+1}^{\operatorname*{rank}(\mathcal{M}_{1}(X^{*}))}\sigma^{2}_{j}(\mathcal{M}_{1}(X^{*}))}\leq\|\mathcal{M}_{1}(X^{*})-A\|_{\mathrm{F}}.$
Taking the inf over all $A\in\mathcal{T}_{(r_{1},r_{2},r_{3})}$, it follows that
$\sqrt{\sum_{j=r_{1}+1}^{\operatorname*{rank}(\mathcal{M}_{1}(X^{*}))}\sigma^{2}_{j}(\mathcal{M}_{1}(X^{*}))}\leq\xi_{(r_{1},r_{2},r_{3})}.$

### v3

@@
-Bias–variance Tradeoff in Tensor Estimation
+Optimal Bias-variance Tradeoff in Matrix and Tensor Estimation
 Abstract
-We study denoising of a third-order tensor when the ground-truth tensor is not necessarily Tucker low-rank. Specifically, we observe
+We study matrix and tensor denoising when the underlying signal is not necessarily low-rank. In the tensor setting, we observe
 $Y=X^{\ast}+Z\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}},$
-where $X^{\ast}$ is the ground-truth tensor, and $Z$ is the noise tensor.
-We propose a simple variant of the higher-order tensor SVD estimator $\widetilde{X}$.
-We show that uniformly over all user-specified Tucker ranks $(r_{1},r_{2},r_{3})$,
-$\|\widetilde{X}-X^{*}\|_{\mathrm{F}}^{2}=O\Big(\kappa^{2}\Big\{r_{1}r_{2}r_{3}+\sum_{k=1}^{3}p_{k}r_{k}\Big\}\;+\;\xi_{(r_{1},r_{2},r_{3})}^{2}\Big)\quad\text{ with high probability.}$
-Here, the bias term $\xi_{(r_{1},r_{2},r_{3})}$ corresponds to the best achievable approximation error of $X^{\ast}$ over the class of tensors with Tucker ranks $(r_{1},r_{2},r_{3})$; $\kappa^{2}$ quantifies the noise level; and the variance term $\kappa^{2}\{r_{1}r_{2}r_{3}+\sum_{k=1}^{3}p_{k}r_{k}\}$ scales with the effective number of free parameters in the estimator $\widetilde{X}$.
-Our analysis achieves a clean rank-adaptive bias–variance tradeoff: as we increase the ranks of estimator $\widetilde{X}$, the bias $\xi(r_{1},r_{2},r_{3})$ decreases and the variance increases.
-As a byproduct we also obtain a convenient bias-variance decomposition for the vanilla low-rank SVD matrix estimators.
+where $X^{\ast}$ is an unknown signal tensor and $Z$ is a noise tensor.
+We propose a one-step variant of the higher-order SVD (HOSVD) estimator, denoted $\widetilde{X}$, and show that, uniformly over any user-specified Tucker ranks $(r_{1},r_{2},r_{3})$, with high probability,
+$\|\widetilde{X}-X^{\ast}\|_{\mathrm{F}}^{2}=O\Big(\kappa^{2}\Big\{r_{1}r_{2}r_{3}+\sum_{k=1}^{3}p_{k}r_{k}\Big\}+\xi_{(r_{1},r_{2},r_{3})}^{2}\Big).$
+Here, $\xi_{(r_{1},r_{2},r_{3})}$ is the best achievable Tucker rank-$(r_{1},r_{2},r_{3})$ approximation error of $X^{\ast}$ (bias), $\kappa^{2}$ quantifies the noise level, and $\kappa^{2}\{r_{1}r_{2}r_{3}+\sum_{k=1}^{3}p_{k}r_{k}\}$ is the variance term scaling with the effective degrees of freedom of $\widetilde{X}$.
+This yields a rank-adaptive bias–variance tradeoff: increasing $(r_{1},r_{2},r_{3})$ decreases the bias $\xi_{(r_{1},r_{2},r_{3})}$ while increasing variance.
+In the matrix setting, we show that truncated SVD achieves an analogous bias–variance tradeoff for arbitrary signal matrices.
+Notably, our matrix result requires no assumptions on the signal matrix, such as finite rank or spectral gaps.
+Finally, we complement our upper bounds with matching information-theoretic lower bounds, showing that the resulting bias–variance tradeoff is minimax optimal up to universal constants in both the matrix and tensor settings.
 1 Introduction
-Low-rank tensor models are a cornerstone of modern machine learning.
-They capture essential structure in high-dimensional signals while offering substantial gains in computation and storage.
-Applications include recommender systems (Koren et al.,, 2009), topic modeling (Blei et al.,, 2003), community detection (Anandkumar et al.,, 2014; Abbe,, 2018), and, more recently, latent variable learning (Diakonikolas and Kane,, 2024) and generative modeling (Hur et al.,, 2023; Peng et al.,, 2023).
-The effectiveness of these methods demonstrate that many real-world signals embedded in high-dimensional ambient spaces can be well-approximated by low-rank structures.
-The study of exactly low-rank estimation is well-developed, particularly in the matrix setting.
-A large body of work has introduced efficient algorithms for tasks such as matrix completion and denoising, including nuclear norm relaxation (Koltchinskii et al.,, 2011), convex optimization (Candes and Recht,, 2012), nonconvex methods (Chi et al.,, 2019), and SVD guarantees with unbalanced matrices (Cai and Zhang,, 2018).
-Parallel efforts in low-rank tensor estimation are well-developed, including the study of higher-order tensor singular value decompositions in the non-random setting (De Lathauwer et al., 2000a, ; De Lathauwer et al., 2000b, ) and Riemannian optimization methods for low-rank Tucker tensor completion (Kressner et al.,, 2014).
-Nevertheless, the tensor setting poses additional challenges: many fundamental problems are computationally intractable, multiple inequivalent notions of rank (e.g.
-CP and Tucker) exist, and best low-rank approximations need not be unique (e.g.
-Hillar and Lim,, 2013).
-Most theoretical guarantees in tensor estimation (e.g.
-Zhang and Xia,, 2018; Han et al.,, 2022), like their matrix counterparts, rely on the assumption that the underlying signal is exactly low-rank.
-In practice, however, signals are not exactly low-rank: their singular spectra typically decay gradually rather than vanishing at a finite rank.
-This creates an interesting gap between theory and practice, since existing analyses provide guarantees under idealized exact low-rankness and leave open the question of how to analyze estimation error in more realistic settings.
-We address this gap by designing and analyzing low-rank tensor estimation when the signal is not low-rank.
-In such settings, estimation error is governed by a bias–variance tradeoff: truncating to a smaller rank incurs approximation bias, while larger ranks inflate variance by amplifying noise.
-In particular, we propose a simple variant of the higher-order singular value decomposition (HOSVD) algorithm and provide the analysis of tensor estimation without assuming low-rankness.
-Our main result (Theorem 1) establishes an explicit bias–variance decomposition of the estimation error:
-  • The bias term quantifies the approximation error incurred by truncating a spectrum.
-  • The variance term matches the optimal rate for exactly low-rank models.
-The success of SVD and HOSVD has been well established by their decades of widespread use.
-The objective of our manuscript is to provide mathematical justification for the information-efficient bias-variance tradeoff achieved by these two classical algorithms, and to establish uniform performance guarantees that hold for any user-specified ranks.
-To achieve this clean bias-variance tradeoff, we utilize classical linear algebra results such as Mirsky’s and Ky Fan’s theorems to develop non-trivial extensions of classical perturbation tools to general tensors.
-As a convenient side result, we also recover a clean bias–variance characterization for matrices, thereby unifying the matrix and tensor perspectives.
-Finally, we validate our theory with simulations that confirm the existence of tradeoffs across different regimes of spectral decay and noise.
+Low-rank matrix and tensor estimation is a cornerstone of modern machine learning.
+Such models capture essential structure in high-dimensional signals while offering substantial gains in computation and storage.
+Applications include recommender systems (Koren et al., 2009), topic modeling (Blei et al., 2003), community detection (Abbe, 2018; Diakonikolas and Kane, 2024), and, more recently, latent variable learning (Anandkumar et al., 2014; Sherman and Kolda, 2020; Zhang and Kileel, 2023) and generative modeling (Hur et al., 2023; Peng et al., 2023).
+The success of these methods suggests that many real-world signals in high-dimensional ambient spaces can be well approximated by low-rank structure.
+The theory of low-rank estimation is well developed in the matrix setting.
+Many efficient methods have been proposed for matrix completion and denoising, including nuclear-norm relaxation (Koltchinskii et al., 2011), convex optimization (Candes and Recht, 2012), nonconvex approaches (Chi et al., 2019), and refined SVD guarantees for highly rectangular matrices (Cai and Zhang, 2018).
+Nevertheless, most existing guarantees are tailored to exactly low-rank signals.
+When the signal is not exactly low-rank, accurate estimation is typically established under additional structural conditions, such as spectral gap assumptions.
+Substantial efforts have advanced low-rank tensor estimation, including higher-order tensor singular value decompositions in deterministic settings (De Lathauwer et al., 2000a, b), symmetric tensor decomposition in deterministic settings (Jin et al., 2024), and Riemannian optimization methods for low-rank Tucker tensor completion (Kressner et al., 2014).
+At the same time, the tensor setting poses additional challenges: many fundamental problems are computationally intractable, multiple inequivalent notions of rank (e.g., CP and Tucker) coexist, and best low-rank approximations need not be unique (e.g.
+Hillar and Lim, 2013).
+Accordingly, most theoretical guarantees in tensor estimation (e.g.
+Anandkumar et al., 2017; Zhang and Xia, 2018; Han et al., 2022), like their matrix counterparts, assume that the underlying signal is exactly low-rank.
+In practice, however, signals are rarely exactly low-rank: their singular values typically decay gradually rather than dropping to zero at a finite rank.
+This creates a gap between theory and practice, since many existing guarantees are tailored to idealized exactly low-rank models and do not directly characterize performance in realistic regimes.
+We close this gap with a rank-adaptive analysis of classical spectral estimators in the approximately low-rank regime. Our contributions are as follows:
+  • Rank-adaptive bias–variance theory. We establish an explicit bias–variance decomposition for the estimation error that holds uniformly over user-specified target ranks.
+  • Matrix setting. We show that truncated SVD achieves the optimal bias–variance tradeoff without any structural assumptions on the signal matrix.
+  • Tensor setting. We analyze a simple one-step variant of HOSVD and obtain analogous guarantees without assuming that the ground truth is exactly Tucker low-rank.
+  • Minimax optimality. We complement our upper bounds with matching information-theoretic lower bounds, establishing minimax optimality up to universal constants.
+  • Methodology.
+Our analysis combines classical linear algebra tools, including Mirsky’s and Ky Fan’s theorems, with modern concentration arguments.
+As a byproduct, we develop new matrix and tensor perturbation bounds that may be of independent interest.
+  • Numerical validation. We validate the predicted tradeoffs through simulations and a real 3D brain MRI example.
 1.1 Notation
@@
 The mode-$1$ product of $B$ with a matrix $M\in\mathbb{R}^{q\times p_{1}}$ is defined as
-$(B\times_{1}M)(j,i_{2},i_{3})=\sum_{i_{1}=1}^{p_{1}}M(i_{1},j)B(i_{1},i_{2},i_{3}).$
+$(B\times_{1}M)_{j,i_{2}i_{3}}=\sum_{i_{1}=1}^{p_{1}}M_{i_{1},j}B_{i_{1},i_{2},i_{3}}.$
 We write the Tucker rank of $B$ as $(r_{1},r_{2},r_{3})$ if
@@
 Universal constants: We use $C_{1},C_{2},\dots$ and $c_{1},c_{2},\dots$ to denote positive constants whose values may differ from place to place.
-2 Bias–variance tradeoff in tensor estimation
+2 Optimal bias–variance tradeoff in matrix estimation
+We consider the model
+$Y=X^{*}+Z\in\mathbb{R}^{m\times n},$ (1)
+where $X^{*}$ is an unknown population matrix with arbitrary rank, and $Z$ is a noise matrix.
+In particular, we do not assume that $X^{*}$ is low-rank, nor do we assume any spectral gap conditions on $X^{*}$.
+Even when $\mathrm{rank}(X^{*})$ is not small, it is common to approximate $X^{*}$ by a rank-$r$ estimator, because (i) low-rank decomposition can dramatically increase the computational and memory efficiency, see Remark 1 for more details; and (ii) the approximation bias can be small, so balancing the bias and variance may lead to smaller estimation error.
+Let
+$X^{*}=\sum_{i=1}^{\min\{m,n\}}\sigma_{i}(X^{*})\,u_{i}^{*}{v_{i}^{*}}^{\!\top},\qquad X_{(r)}^{*}=\sum_{i=1}^{r}\sigma_{i}(X^{*})\,u_{i}^{*}{v_{i}^{*}}^{\!\top}.$
+By the Eckart–Young–Mirsky theorem (Eckart and Young, 1936), $X_{(r)}^{*}$ is the best rank-$r$ approximation to $X^{*}$ in Frobenius norm in the sense that
+$\|X^{*}-X^{*}_{(r)}\|_{\mathrm{F}}\leq\|X^{*}-W\|_{\mathrm{F}}\text{ for any }W\in\mathbb{R}^{m\times n}\text{ with }\operatorname*{rank}(W)\leq r.$
+Suppose we observe $Y$ instead of $X^{*}$, and write the SVD of $Y$ as
+$Y=\sum_{i=1}^{\min\{m,n\}}\sigma_{i}(Y)\,u_{i}v_{i}^{\top},\qquad Y_{(r)}=\sum_{i=1}^{r}\sigma_{i}(Y)\,u_{i}v_{i}^{\top}.$ (2)
+That is, $Y_{(r)}$ is the rank-$r$ truncation of the SVD of $Y$. We now provide an upper bound between $Y_{(r)}$ and the unknown matrix $X^{*}$.
+Theorem 1.
+Under (1), let $Y_{(r)}$ be defined as in (2). Then
+$\|Y_{(r)}-X^{*}\|_{\mathrm{F}}\ \leq\ (2+\sqrt{2})(\sqrt{r}\|Z\|+\xi_{(r)}),$ (3)
+where $\xi_{(r)}=\|X_{(r)}^{*}-X^{*}\|_{\mathrm{F}}=\sqrt{\sum_{i=r+1}^{\min\{m,n\}}\sigma_{i}^{2}(X^{*})}$.
+Assumption-free with respect to $X^{*}$.
+Theorem 1 requires no structural conditions on $X^{*}$.
+In particular, $\operatorname*{rank}(X^{*})$ may be arbitrary, and no spectral gap assumptions are needed.
+Adaptivity in Theorem 1.
+For any prescribed rank $r$, the rank-$r$ truncated SVD estimator $Y_{(r)}$ satisfies the guarantee in Theorem 1.
+In other words, the theorem holds uniformly over all $r\geq 1$, so it applies to the vanilla SVD at any user-specified target rank.
+Constant in (3). Tracing the proof of Theorem 1 shows that one may take the explicit constant $C=2+\sqrt{2}\approx 3.414$ in (3). We did not attempt to optimize this further.
+Bias–variance tradeoff.
+Theorem 1 exhibits a clear bias–variance tradeoff for rank-$r$ estimation.
+The bias term $\xi_{(r)}=\|X^{*}-X_{(r)}^{*}\|_{\mathrm{F}}$ is the best possible rank-$r$ approximation error of $X^{*}$ and decreases as $r$ increases.
+The standard deviation term, being order $(\sqrt{r}\|Z\|)$, captures the cost of estimating additional singular components under noise and increases with $r$.
+In what follows, we illustrate the applicability and optimality of Theorem 1 in two popular matrix estimation settings.
+Corollary 1.
+Suppose $Z\in\mathbb{R}^{m\times n}$ is a sub-Gaussian random matrix in the sense that for any $v\in\mathbb{R}^{m}$ with $\|v\|_{2}=1$, and $w\in\mathbb{R}^{n}$ with $\|w\|_{2}=1$, it holds that
+$\|v^{\top}Zw\|_{\psi_{2}}\leq\kappa.$
+Then with probability at least $1-\exp(-(m+n))$, there exists a universal constant $C>0$ such that
+[…]
+where $\xi_{(r)}=\|X_{(r)}^{*}-X^{*}\|_{\mathrm{F}}=\sqrt{\sum_{i=r+1}^{\min\{m,n\}}\sigma_{i}^{2}(X^{*})}$.
+Next, we show that the error bound in Corollary 1 is minimax optimal.
+Theorem 2.
+Let $r,m,n$ be arbitrary positive integers such that $r\leq\min\{m,n\}$. For $\xi>0$, let $\mathcal{M}_{r}^{\xi}$ be such that
+[…]
+Consider the model $Y=X^{\ast}+Z\in\mathbb{R}^{m\times n}$ where the entries of $Z$ are i.i.d.
+$\mathrm{subGaussian}(0,\kappa^{2})$.
+Suppose in addition that $\xi^{2}\leq(m-r)n\kappa^{2}$.
+Then there exist a universal constant $c>0$ such that
+$\inf_{\widehat{X}}\sup_{X^{\ast}\in\mathcal{M}_{r}^{\xi}}\ \mathbb{E}\bigl\|\widehat{X}-X^{\ast}\bigr\|_{\mathrm{F}}^{2}\$$\geq\ c\Bigl\{\kappa^{2}r(m+n)+\xi^{2}\Bigr\}.$
+Corollary 1 and Theorem 2 together show that (3) is minimax optimal when $Z$ is sub-Gaussian. Next, we illustrate the applicability of Theorem 1 in covariance matrix estimation.
+Corollary 2.
+Let $Z_{1},\dots,Z_{N}\in\mathbb{R}^{n}$ be i.i.d. mean-zero sub-Gaussian random vectors with covariance $\Sigma$. Define the sample covariance matrix
+$Y=\frac{1}{N}\sum_{k=1}^{N}Z_{k}Z_{k}^{\top},$
+and let $X^{*}=\mathbb{E}[Y]$ be the population covariance matrix. Suppose $Z_{k}$ is sub-Gaussian in the sense that
+$\|u^{\top}Z_{k}\|_{\psi_{2}}\leq\kappa\quad\text{for all }u\in\mathbb{R}^{n},\ \|u\|_{2}=1.$
+Then with probability at least $1-2\exp(-n)$, there exists a universal constant $C>0$ such that
+$\|Y_{(r)}-X^{*}\|_{\mathrm{F}}\leq C\Bigl(\kappa\sqrt{r}\Bigl[\sqrt{\frac{n}{N}}+\frac{n}{N}\Bigr]+\xi_{(r)}\Bigr),$
+where $\xi_{(r)}=\|X_{(r)}^{*}-X^{*}\|_{\mathrm{F}}=\sqrt{\sum_{i=r+1}^{\min\{m,n\}}\sigma_{i}^{2}(X^{*})}$.
+Remark 1 (Computational and storage savings from matrix low-rank factorization).
+Suppose $Y\in\mathbb{R}^{m\times n}$ and let $Y_{(r)}=U_{(r)}\,\Sigma_{(r)}\,V_{(r)}^{\top}$ be a rank-$r$ factorization (e.g.
+the truncated SVD), where $U_{(r)}\in\mathbb{R}^{m\times r}$, $V_{(r)}\in\mathbb{R}^{n\times r}$ satisfy $U_{(r)}^{\top}U_{(r)}=I_{r}$, $V_{(r)}^{\top}V_{(r)}=I_{r}$, and $\Sigma_{(r)}=\mathrm{diag}(\sigma_{1},\dots,\sigma_{r})\in\mathbb{R}^{r\times r}$.
+For a dense multiplication with $v\in\mathbb{R}^{n}$, computing $Yv$ costs $O(mn)$ operations.
+In contrast, using the factorization we compute successively
+$t=V_{(r)}^{\top}v\in\mathbb{R}^{r}\ (\text{cost}\approx 2nr),\quad s=\Sigma_{(r)}t\in\mathbb{R}^{r}\ (\text{cost}\approx r),\quad w=U_{(r)}s\in\mathbb{R}^{m}\ (\text{cost}\approx 2mr).$
+Hence $Y_{(r)}v\;=\;U_{(r)}\,\Sigma_{(r)}\,V_{(r)}^{\top}v$ can be computed in $O\big(mr+nr\big)$ operations.
+The SVD factorization also reduces storage costs.
+Storing the full matrix $Y$ requires $O(mn)$ memory, whereas storing the factors $U_{(r)}\in\mathbb{R}^{m\times r}$, $\Sigma_{(r)}\in\mathbb{R}^{r\times r}$, and $V_{(r)}\in\mathbb{R}^{n\times r}$ requires only $O(mr+nr)$ scalars.
+Therefore, both computation and storage using the truncated SVD are significantly smaller than in the full matrix case, especially when $r\ll\min\{m,n\}$.
+3 Optimal bias–variance tradeoff in tensor estimation
 We consider the noisy tensor model
 $Y=X^{*}+Z\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}},$ (4)
-where $X^{*}$ is an unknown signal and $Z$ is a random perturbation.
-Even when $X^{*}$ has a full Tucker rank $(p_{1},p_{2},p_{3})$, it is common to approximate $X^{*}$ by a Tucker rank-$(r_{1},r_{2},r_{3})$ estimator.
-This is because (i) low-rank structure yields dramatic gains in computation and memory, see Remark 2; and (ii) the approximation bias can decay quickly when approximating $X^{*}$ by a Tucker low-rank tensor, so balancing bias and variance may lead to a smaller estimation error.
-Given a target Tucker rank $(r_{1},r_{2},r_{3})$, we consider the following simple variant of the HOSVD algorithm (De Lathauwer et al., 2000b, ).
+where $X^{*}$ is an unknown signal tensor and $Z$ is a noise tensor.
+Importantly, we note that the results in this section do not assume that $X^{*}$ is exactly Tucker low-rank.
+Given a user-specified Tucker rank $(r_{1},r_{2},r_{3})$, we consider the following simple variant of the HOSVD algorithm (De Lathauwer et al., 2000b).
 Algorithm 1 One-step HOSVD
@@
 Let $\mathcal{T}_{(r_{1},r_{2},r_{3})}$ denote the class of tensors in $\mathbb{R}^{p_{1}\times p_{2}\times p_{3}}$ with Tucker rank at most $(r_{1},r_{2},r_{3})$, i.e.
-$\mathcal{T}_{(r_{1},r_{2},r_{3})}=\big\{A\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}}:\ \operatorname*{rank}(\mathcal{M}_{k}(A))\leq r_{k},\ k=1,2,3\big\}.$
+$\mathcal{T}_{(r_{1},r_{2},r_{3})}=\big\{A\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}}:\ \operatorname*{rank}(\mathcal{M}_{k}(A))\leq r_{k},\ k=1,2,3\big\}.$ (5)
 For any tensor $X^{*}\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}}$, let $\xi_{(r_{1},r_{2},r_{3})}$ denote the best Tucker rank-$(r_{1},r_{2},r_{3})$ approximation error:
@@
 Balancing these two leads to improved accuracy.
-Tightness.
-Our upper bound in (3) attains the minimax optimal rate (up to constants) over the class of tensors with Tucker rank at most $(r_{1},r_{2},r_{3})$.
-Indeed, by definition, $\xi_{(r_{1},r_{2},r_{3})}$ is the smallest possible approximation error of $X^{*}$ over $\mathcal{T}_{(r_{1},r_{2},r_{3})}$.
-In addition, Zhang and Xia, (2018) prove that, when $X^{*}$ is exactly low-rank with Tucker rank $(r_{1},r_{2},r_{3})$, the term $\kappa^{2}\Bigl(\sum_{k=1}^{3}p_{k}r_{k}+r_{1}r_{2}r_{3}\Bigr)$ is minimax optimal up to a universal constant.
-For arbitrary (e.g.
-full-rank) tensors, our bound should be interpreted as an upper bound; a matching minimax lower bound for that broader class is not known.
 Proof techniques.
 For the bias, we leverage classical linear algebra tools, such as Mirsky’s and Ky Fan’s theorems, to characterize the optimal approximation error appearing in (7).
-To handle the variance of the estimator $\widetilde{X}$, we adopt existing techniques from Zhang and Xia, (2018).
+To handle the variance of the estimator $\widetilde{X}$, we adopt existing techniques from Zhang and Xia (2018).
 Remark 2 (Assumptions in Theorem 3).
-In Theorem 1, we assume that the entries of the noise tensor $Z$ is i.i.d. sub-Gaussian. This is a commonly seen condition in the tensor literature (e.g. Zhang and Xia,, 2018; Han et al.,, 2022).
+In Theorem 3, we assume that the entries of the noise tensor $Z$ are i.i.d. sub-Gaussian. This is a commonly seen condition in the tensor literature (e.g. Zhang and Xia, 2018; Han et al., 2022).
 The signal-to-noise ratio (SNR) condition (6) in Theorem 3 is also commonly seen in literature.
-In fact, if the ground truth $X^{*}$ is exactly Tucker low-rank with Tucker rank $(r_{1},r_{2},r_{3})$, then $\sigma_{r_{k}+1}(\mathcal{M}_{k}(X^{*}))=0$ and (2) reduces to the SNR condition used in Zhang and Xia, (2018), when $r_{\max}$ is a bounded constant.
-Remark 2.
+In fact, suppose the ground truth $X^{*}$ is exactly Tucker low-rank with Tucker rank $(r_{1},r_{2},r_{3})$.
+Then $\sigma_{r_{k}+1}(\mathcal{M}_{k}(X^{*}))=0$, and (6) reduces to the SNR condition used in Zhang and Xia (2018) when $r_{\max}$ is a bounded constant.
+We present a relative Frobenius error bound as a corollary of Theorem 3.
+Corollary 3 (Relative estimation error).
+Suppose the conditions of Theorem 3 hold and $r_{k}\leq\sqrt{p_{\min}}$ for $k=1,2,3$. Then, with probability at least $1-C_{1}p_{1}p_{2}p_{3}\exp(-C_{2}p_{\min})$,
+$\frac{\|\widetilde{X}-X^{*}\|_{\mathrm{F}}}{\|X^{*}\|_{\mathrm{F}}}\;\leq\;C_{3}\Bigg\{\sqrt{\kappa^{2}\sum_{k=1}^{3}\frac{p_{k}}{\sigma^{2}_{r_{k}}\!\big(\mathcal{M}_{k}(X^{*})\big)}}\;+\;\frac{\xi_{(r_{1},r_{2},r_{3})}}{\|X^{*}\|_{\mathrm{F}}}\Bigg\}.$ (8)
+We next show that the rate in (7) is minimax optimal by proving a matching lower bound.
+Theorem 4 (Minimax lower bound).
+Let $r_{1},r_{2},r_{3},p_{1},p_{2},p_{3}$ be arbitrary integers such that $r_{i}\leq\min\{p_{i},\prod_{j\not=i}p_{j}\}$ for all $i\in\{1,2,3\}$.
+Let $\mathcal{T}_{(r_{1},r_{2},r_{3})}$ be the class of Tucker low-rank tensor as in (5) and for $\xi>0$, let
+$\mathcal{T}_{(r_{1},r_{2},r_{3})}^{\xi}=\big\{B\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}}:\inf_{A\in\mathcal{T}_{(r_{1},r_{2},r_{3})}}\|B-A\|_{\mathrm{F}}\leq\xi\big\}.$
+Consider the model $Y=X^{\ast}+Z\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}}$ where the entries of $Z$ are i.i.d.
+$\mathrm{subGaussian}(0,\kappa^{2})$.
+For any ranks $(r_{1},r_{2},r_{3})$ and any $0\leq\xi\leq\kappa^{2}p_{1}p_{2}p_{3}$, there exist a universal constant $c>0$ such that
+$\inf_{\widehat{X}}\ \sup_{X^{\ast}\in\mathcal{T}_{(r_{1},r_{2},r_{3})}^{\xi}}\ \mathbb{E}\bigl\|\widehat{X}-X^{\ast}\bigr\|_{\mathrm{F}}^{2}\$$\geq\ c\Bigl\{\kappa^{2}\big(\textstyle\sum_{k=1}^{3}p_{k}r_{k}+r_{1}r_{2}r_{3}\big)+\xi^{2}\Bigr\}.$ (9)
+Tightness.
+The upper bound in Theorem 3 and the matching lower bound in Theorem 4 together show that the error bound in (7) is minimax optimal up to universal constants over the approximately Tucker low-rank class $\mathcal{T}^{\xi}_{(r_{1},r_{2},r_{3})}$.
+Advantages of low-rank representations.
+Even when $X^{*}$ has full Tucker rank $(p_{1},p_{2},p_{3})$, it is common to approximate $X^{*}$ by a Tucker rank-$(r_{1},r_{2},r_{3})$ estimator.
+This is because (i) the approximation bias can be small when $X^{*}$ is well approximated by a low-rank tensor, so balancing bias and variance may reduce estimation error; and (ii) low-rank structure can yield substantial gains in computation and storage, as demonstrated in the following remark.
+Remark 3 (Computational and storage savings from tensor low-rank estimation).
 Suppose $Y\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}}$ and let
-$Y_{(r_{1},r_{2},r_{3})}=S\times_{1}U_{1}\times_{2}U_{2}\times_{3}U_{3},$
+$Y_{(r_{1},r_{2},r_{3})}=S\times_{1}U_{1}\times_{2}U_{2}\times_{3}U_{3}$
 be a Tucker rank-$(r_{1},r_{2},r_{3})$ approximation of $Y$, where $U_{k}\in\mathbb{R}^{p_{k}\times r_{k}}$ satisfying $U_{k}^{\top}U_{k}=I_{r_{k}}$ for $k=1,2,3$, and $S\in\mathbb{R}^{r_{1}\times r_{2}\times r_{3}}$ is the core tensor.
@@
 $t_{k}=U_{k}^{\top}v_{k}\in\mathbb{R}^{r_{k}}\ (\text{cost }\approx 2\sum_{k=1}^{3}p_{k}r_{k}),\quad S^{\prime}=S\times_{1}t_{1}\in\mathbb{R}^{r_{2}\times r_{3}}\ (\text{cost }\approx 2r_{1}r_{2}r_{3}),$
-$s=t_{2}^{\top}S^{\prime}\in\mathbb{R}^{r_{3}}\quad(\text{cost }\approx 2r_{2}r_{3}),\quad w=s^{\top}t_{3}\in\mathbb{R}\quad(\text{cost }\approx r_{3}).$
+$s=t_{2}^{\top}S^{\prime}\in\mathbb{R}^{r_{3}}\ (\text{cost }\approx 2r_{2}r_{3}),\quad w=s^{\top}t_{3}\in\mathbb{R}\ (\text{cost }\approx r_{3}).$
 Thus, vector multiplication with Tucker factors reduce the computational cost from $O(p_{1}p_{2}p_{3})$ to $O(\sum_{k=1}^{3}p_{k}r_{k}+r_{1}r_{2}r_{3})$.
@@
 Therefore, both computation and storage of Tucker decomposition scale with $O(\sum_{k=1}^{3}p_{k}r_{k}+r_{1}r_{2}r_{3})$, which is substantially smaller than the dense case when $r_{k}\ll p_{k}$.
-3 Bias–variance tradeoff in matrix estimation
-We consider the model
-$Y=X^{*}+Z\in\mathbb{R}^{m\times n},$ (4)
-where $X^{*}$ is an unknown matrix with arbitrary rank, and $Z$ is a perturbation (noise) matrix.
-Even when $\mathrm{rank}(X^{*})$ is not small, it is common to approximate $X^{*}$ by a rank-$r$ estimator, because (i) low-rank decomposition can dramatically increase the computational and memory efficiency, see Remark 3 for more details; and (ii) the approximation bias can be small, so balancing the bias and variance may lead to smaller estimation error.
-Let
-$X^{*}=\sum_{i=1}^{\min\{m,n\}}\sigma_{i}(X^{*})\,u_{i}^{*}{v_{i}^{*}}^{\!\top},\qquad X_{(r)}^{*}=\sum_{i=1}^{r}\sigma_{i}(X^{*})\,u_{i}^{*}{v_{i}^{*}}^{\!\top}.$
-By the Eckart–Young–Mirsky theorem (Eckart and Young,, 1936), $X_{(r)}^{*}$ is the best rank-$r$ approximation to $X^{*}$ in Frobenius norm in the sense that
-$\|X^{*}-X^{*}_{(r)}\|_{\mathrm{F}}\leq\|X^{*}-W\|_{\mathrm{F}}\text{ for any }W\in\mathbb{R}^{m\times n}\text{ with }\operatorname*{rank}(W)\leq r.$
-Suppose we observe $Y$ instead of $X^{*}$, and write the SVD of $Y$ as
-$Y=\sum_{i=1}^{\min\{m,n\}}\sigma_{i}(Y)\,u_{i}v_{i}^{\top},\qquad Y_{(r)}=\sum_{i=1}^{r}\sigma_{i}(Y)\,u_{i}v_{i}^{\top}.$ (5)
-That is, $Y_{(r)}$ is the rank-$r$ truncation of the SVD of $Y$. We now provide an upper bound between $Y_{(r)}$ and the unknown matrix $X^{*}$.
-Theorem 2.
-Under (4), let $Y_{(r)}$ be defined as in (5). Then
-$\|Y_{(r)}-X^{*}\|_{\mathrm{F}}\ \leq\ (2+\sqrt{2})(\sqrt{r}\|Z\|+\xi_{(r)}),$ (6)
-where $\xi_{(r)}=\|X_{(r)}^{*}-X^{*}\|_{\mathrm{F}}=\sqrt{\sum_{i=r+1}^{\min\{m,n\}}\sigma_{i}^{2}(X^{*})}$.
-Proof of Theorem 2.
-By the triangle inequality,
-$\|Y_{(r)}-X^{*}\|_{\mathrm{F}}\leq\|X_{(r)}^{*}-X^{*}\|_{\mathrm{F}}+\|Y_{(r)}-X_{(r)}^{*}\|_{\mathrm{F}}=\xi_{(r)}+\|Y_{(r)}-X_{(r)}^{*}\|_{\mathrm{F}}.$ (7)
-Since $\mathrm{rank}(Y_{(r)}-X_{(r)}^{*})\leq 2r$, we have
-$\|Y_{(r)}-X_{(r)}^{*}\|_{\mathrm{F}}^{2}=\sum_{k=1}^{2r}\sigma_{k}^{2}(Y_{(r)}-X_{(r)}^{*}).$ (8)
-Write
-$Y_{(r)}-X_{(r)}^{*}=(X^{*}-X_{(r)}^{*})+Z+(Y_{(r)}-Y).$
-It follows from Ky Fan’s Theorem (Lemma 25) that,
-$\sqrt{\sum_{i=1}^{2r}\sigma_{i}^{2}(Y_{(r)}-X_{(r)}^{*})}\leq\sqrt{\sum_{i=1}^{2r}\sigma_{i}^{2}(X^{*}-X_{(r)}^{*})}+\sqrt{\sum_{i=1}^{2r}\sigma_{i}^{2}(Z)}+\sqrt{\sum_{i=1}^{2r}\sigma_{i}^{2}(Y_{(r)}-Y)}.$
-Using the identity $\sigma_{i}(X^{*}-X^{*}_{(r)})=\sigma_{r+i}(X^{*})$ and likewise for $Y$, we obtain
-$\|Y_{(r)}-X^{*}_{(r)}\|_{\mathrm{F}}$ $\leq\sqrt{\sum_{i=1}^{2r}\sigma_{r+i}^{2}(X^{*})}+\sqrt{\sum_{i=1}^{2r}\sigma_{i}^{2}(Z)}+\sqrt{\sum_{i=1}^{2r}\sigma_{r+i}^{2}(Y)}.$ (9)
-To further simplify the third term, we apply Weyl’s inequality (Theorem 24) to obtain
-$\sigma_{r+i}(Y)\leq\sigma_{r+i}(X^{*})+\sigma_{1}(Z),$
-which gives
-$\sigma_{r+i}^{2}(Y)\leq 2\left\{\sigma_{r+i}^{2}(X^{*})+\sigma_{1}^{2}(Z)\right\}.$
-Substituting this into (9), we get
-$\|Y_{(r)}-X^{*}_{(r)}\|_{\mathrm{F}}$ $\leq\sqrt{\sum_{i=1}^{2r}\sigma_{r+i}^{2}(X^{*})}+\sqrt{\sum_{i=1}^{2r}\sigma_{i}^{2}(Z)}+\sqrt{2\sum_{i=1}^{2r}\sigma_{r+i}^{2}(X^{*})}+2\sqrt{r}\sigma_{1}(Z)$
-$\leq(1+\sqrt{2})\sqrt{\sum_{i=1}^{2r}\sigma_{r+i}^{2}(X^{*})}+({2+\sqrt{2}})\sqrt{r}\|Z\|$
-$=(1+\sqrt{2})\sqrt{\sum_{i=r+1}^{3r}\sigma_{i}^{2}(X^{*})}+({2+\sqrt{2}})\sqrt{r}\|Z\|,$
-where the second inequality follows from the fact that for any $k\geq 1$, we have $\sigma_{k}(Z)\leq\sigma_{1}(Z)=\|Z\|.$ Substituting the bound of $\|Y_{(r)}-X^{*}_{(r)}\|_{\mathrm{F}}$ into (7), we obtain
-$\|Y_{(r)}-X^{*}\|_{\mathrm{F}}\leq(2+\sqrt{2})\sqrt{\sum_{i=r+1}^{\min\{m,n\}}\sigma^{2}_{i}(X^{*})}+(2+\sqrt{2})\sqrt{r}\|Z\|.$
-Adaptivity in Theorem 2.
-For any prescribed rank $r$, the rank-$r$ truncated SVD estimator $Y_{(r)}$ satisfies the guarantee in Theorem 2.
-In other words, the theorem holds uniformly over all $r\geq 1$, so it applies to the vanilla SVD at any user-specified target rank.
-Constant in (6). Tracing the proof of Theorem 2 shows that one may take the explicit constant $C=2+\sqrt{2}\approx 3.414$ in (6). We did not attempt to optimize this further.
-Bias–variance tradeoff.
-Theorem 2 exhibits a clear bias–variance tradeoff for rank-$r$ estimation.
-The bias term $\xi_{(r)}=\|X^{*}-X_{(r)}^{*}\|_{\mathrm{F}}$ is the best possible rank-$r$ approximation error of $X^{*}$ and decreases as $r$ increases.
-The standard deviation term, being order $(\sqrt{r}\|Z\|)$, captures the cost of estimating additional singular components under noise and increases with $r$.
-Remark 3.
-Suppose $Y\in\mathbb{R}^{m\times n}$ and let $Y_{(r)}=U_{(r)}\,\Sigma_{(r)}\,V_{(r)}^{\top}$ be a rank-$r$ factorization (e.g.
-the truncated SVD), where $U_{(r)}\in\mathbb{R}^{m\times r}$, $V_{(r)}\in\mathbb{R}^{n\times r}$ satisfy $U_{(r)}^{\top}U_{(r)}=I_{r}$, $V_{(r)}^{\top}V_{(r)}=I_{r}$, and $\Sigma_{(r)}=\mathrm{diag}(\sigma_{1},\dots,\sigma_{r})\in\mathbb{R}^{r\times r}$.
-For a dense multiplication with $v\in\mathbb{R}^{n}$, computing $Yv$ costs $O(mn)$ operations.
-In contrast, using the factorization we compute successively
-$t=V_{(r)}^{\top}v\in\mathbb{R}^{r}\ (\text{cost }\approx 2nr),\quad s=\Sigma_{(r)}t\in\mathbb{R}^{r}\ (\text{cost }\approx r),\quad w=U_{(r)}s\in\mathbb{R}^{m}\quad(\text{cost }\approx 2mr).$
-Hence $Y_{(r)}v\;=\;U_{(r)}\,\Sigma_{(r)}\,V_{(r)}^{\top}v$ can be computed in $O\big(mr+nr\big)$ operations.
-Beyond computational savings, the SVD factorization also reduces storage costs.
-Storing the full matrix $Y$ requires $O(mn)$ memory, whereas storing the factors $U_{(r)}\in\mathbb{R}^{m\times r}$, $\Sigma_{(r)}\in\mathbb{R}^{r\times r}$, and $V_{(r)}\in\mathbb{R}^{n\times r}$ requires only $O(mr+nr)$ scalars.
-Therefore, both computation and storage using the truncated SVD are significantly smaller than in the full matrix case, especially when $r\ll\min\{m,n\}$.
-In the following, we illustrate the application of Theorem 2 in three matrix estimation settings.
-Corollary 3.
-Suppose (4) holds and that the entries of $Z$ are independent sub-Gaussian random variables with
-$\|Z_{ij}\|_{\psi_{2}}\leq\kappa.$
-Then with probability at least $1-\exp(-(m+n))$, there exists a universal constant $C>0$ such that
-$\|Y_{(r)}-X^{*}\|_{\mathrm{F}}\leq C\big(\|X_{(r)}^{*}-X^{*}\|_{\mathrm{F}}+\kappa\sqrt{r(m+n)}\big).$
-Proof of Corollary 3.
-From Vershynin, (2018) section 4.4.2, it follows that $\|Z\|\leq C_{1}\kappa\sqrt{m+n}$ with probability at least $1-\exp(-(m+n))$ . The desired result follows immediately from Theorem 2. ∎
-Corollary 4.
-Suppose $Z\in\mathbb{R}^{m\times n}$ is a sub-Gaussian random matrix in the sense that for any $v\in\mathbb{R}^{m}$ with $\|v\|_{2}=1$, and $w\in\mathbb{R}^{n}$ with $\|w\|_{2}=1$, it holds that
-$\|v^{\top}Zw\|_{\psi_{2}}\leq\kappa.$
-Then with probability at least $1-\exp(-(m+n))$, there exists a universal constant $C>0$ such that
-$\|Y_{(r)}-X^{*}\|_{\mathrm{F}}\leq C\big(\|X_{(r)}^{*}-X^{*}\|_{\mathrm{F}}+\kappa\sqrt{r(m+n)}\big).$
-Proof of Corollary 4.
-By Lemma 12, $\|Z\|\leq C_{1}\sqrt{m+n}$ with probability at least $1-\exp(-(m+n))$. The desired result follows immediately from Theorem 2. ∎
-Corollary 5.
-Let $Z_{1},\dots,Z_{N}\in\mathbb{R}^{n}$ be i.i.d. mean-zero sub-Gaussian random vectors with covariance $\Sigma$. Define the sample covariance matrix
-$Y=\frac{1}{N}\sum_{k=1}^{N}Z_{k}Z_{k}^{\top},$
-and let $X^{*}=\mathbb{E}[Y]$ be the population covariance matrix. Suppose $Z_{k}$ is sub-Gaussian in the sense that
-$\|u^{\top}Z_{k}\|_{\psi_{2}}\leq\kappa\quad\text{for all }u\in\mathbb{R}^{n},\ \|u\|_{2}=1.$
-Then with probability at least $1-2\exp(-n)$, there exists a universal constant $C>0$ such that
-$\|Y_{(r)}-X^{*}\|_{\mathrm{F}}\leq C\Bigl(\|X_{(r)}^{*}-X^{*}\|_{\mathrm{F}}+\kappa^{2}\sqrt{r}\Bigl[\sqrt{\frac{n}{N}}+\frac{n}{N}\Bigr]\Bigr).$
-Proof of Corollary 5.
-Let $\widetilde{Z}:=Y-X^{*}.$ By Remark 4.7.3 of Vershynin, (2018), with probability at least $1-\exp(-cn)$ one has $\|\widetilde{Z}\|\;\leq\;C_{1}\kappa^{2}\Bigl[\sqrt{\frac{n}{N}}+\frac{n}{N}\Bigr].$ Applying Theorem 2 with $Z$ replaced by $\widetilde{Z}$ yields the desired bound.
-∎
 4 Numerical experiments
-In many applied settings, selecting the single best rank $r$ is inherently difficult in finite samples with unknown latent structure.
-Our analysis instead provides guarantees for singular-value decompositions in both the matrix (SVD) and tensor (HOSVD) regimes across a broad range of $r$.
-Consequently, rather than hinging on a brittle rank-selection heuristic, we empirically demonstrate that both truncated SVD and one-step HOSVD (Algorithm 1) are robust whenever ranks are chosen neither too small (underfitting/bias) nor too large (overfitting/variance).
-This is consistent with the bias–variance tradeoff formalized in Theorems 2 and 1.
-Moreover, in practice the choice of $r$, is frequently driven by diverse aims other than merely minimizing mean-squared error, including visualization, interpretability, or downstream decision tasks.
-This is reflected in widespread, long-standing use across biology/genomics (Price et al.,, 2006), economics (Filmer and Pritchett,, 2001), and computer vision (Turk and Pentland,, 1991).
-We evaluate these predictions on (i) a real 3D brain-MRI dataset with controlled additive noise (IXI; T1-weighted volumes (IXI,, 2002)), and (ii) simulated data for both matrices and third–order tensors.
+We numerically demonstrate the bias-variance tradeoff developed in our theory.
+We begin by describing a practical approach of selecting the rank parameters used in both SVD and one-step HOSVD (Algorithm 1).
+Remark 4 (Practical choice of target ranks).
+There are several well-known approaches for selecting the target Tucker ranks.
+To select the rank parameters in a data-driven manner, we recommend following the discussion in Ozdemir et al.
+(2017) and adopting an adaptive thresholding scheme.
+More precisely, one selects $r_{k}$ such that $\sigma_{r_{k}}^{2}(\mathcal{M}_{k}(Y))\geq\tau$, where $\tau>0$ is prescribed threshold.
+A simple and effective choice is $\tau=\alpha\|Y\|_{\mathrm{F}}^{2}$ with $\alpha=0.01$.
+Under this rule, all singular values of $\mathcal{M}_{k}(Y)$ up to the largest index $j$ satisfying $\sigma_{j}^{2}(\mathcal{M}_{k}(Y))\geq\tau$ are retained, and the remainder are discarded.
+In many applied settings, however, selecting the single best rank $r$ is inherently difficult in finite samples, particularly with unknown latent structure.
+Our analysis instead provides guarantees for singular-value decompositions in both the matrix and tensor regimes across a broad range of $r$.
+Consequently, rather than relying on a specific rank-selection heuristic, we empirically demonstrate that both truncated SVD and one-step HOSVD are robust provided that the chosen ranks are neither too small (underfitting) nor too large (overfitting).
+This behavior is consistent with the bias–variance tradeoff formalized in Theorems 1 and 3.
+Moreover, in practice the choice of $r$ is often guided by considerations beyond minimizing mean-squared error, such as visualization, interpretability, or performance on downstream tasks.
+This perspective aligns with the widespread and long standing use of low-rank methods across diverse application domains, biology and genomics (Price et al., 2006), economics (Filmer and Pritchett, 2001), and computer vision (Turk and Pentland, 1991).
+We evaluate the bias-variance tradeoff on (i) a real 3D brain-MRI dataset with controlled additive noise (Group, 2002, T1-weighted volumes), and (ii) simulated data in both matrix and third–order tensor settings.
 4.1 Real data: 3D brain MRI with controlled noise
-We evaluate on a T1-weighted brain MRI volume drawn from the IXI dataset (IXI,, 2002).
-The selected volume is a 3D tensor of shape $256\times 256\times 150$.
-To probe low-rank behavior in a realistic preprocessing pipeline, the volume is prepared in two wavelet-smoothed variants using standard orthonormal families—(a) Daubechies-6 (db6) and (b) Symlet-8 (sym8) (Daubechies,, 1988, 1992).
-We implement separable 3D multilevel decompositions (wavedecn) with periodic boundary handling, and apply subband-wise soft-thresholding via BayesShrink (Chang et al.,, 2000; Donoho and Johnstone,, 1994).
-The literature supports wavelet shrinkage as an effective denoising/smoothing step for MR magnitude images (Nowak,, 1999; Wood and Johnson,, 1999).
-Noise model and ranks. For our prepared volume $X$ and each noise level $\lambda\in\{0.01,0.05,0.1,0.2\}$ we generate
+We evaluate on five T1-weighted brain MRI volumes drawn from the IXI dataset (Group, 2002). The selected volumes are third-order tensors of shape $256\times 256\times 150$.
+For each volume, we consider the original tensor and two wavelet-smoothed tensors obtained using orthonormal wavelet families Daubechies-6 (db6) and Symlet-8 (sym8).
+Wavelet smoothing is performed using separable 3D multilevel decompositions (wavedecn) followed by subband-wise soft-thresholding with BayesShrink (Donoho and Johnstone, 1994; Chang et al., 2000).
+This approach is well supported in the MRI literature as an effective smoothing strategy for MRI volumes (Nowak, 1999; Wood and Johnson, 1999).
+Noise model and ranks. For each noise level $\lambda\in\{0.2,0.3,0.4,0.5,0.6,0.7\}$ and each volume $X^{\ast}$, we generate
 $Y\;=\;X^{*}+Z,\qquad Z_{ijk}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathcal{N}(0,\kappa^{2}),\quad\kappa\;=\;\lambda\,\frac{\|X^{*}\|_{\mathrm{F}}}{\sqrt{p_{1}p_{2}p_{3}}}\,$,
-which keeps the per-entry SNR comparable across images and variants. We compute one-step HOSVD reconstructions (Algorithm 1) $\widetilde{X}$ at Tucker rank $(r,r,r)$ for $r\in\{50,65,80\}.$
+which keeps the per-entry SNR comparable across images and variants.
+We compute one-step HOSVD reconstructions (Algorithm 1) $\widetilde{X}$ at Tucker rank $(r,r,r)$ for $r\in\{35,50,60,65,70,80,90,95\}.$
 Findings.
-Figure 1 shows representative mid-sagittal slices from the reconstructions of the wavelet-smoothed variants of single volume.
-Qualitatively, the slices exhibit the expected progression: very small ranks $r$ lead to underfitting with over-smoothed, low-rank artifacts, while an intermediate range of $r$ preserves cortical detail and simultaneously suppresses background fluctuations.
-Figure 1: Mid-sagittal MRI slice of the reconstructed volume at selected ranks and noise levels $\lambda$ for the db6 variant (left), and for the sym8 variant (right).
+Figure 1 displays mid-sagittal slices from reconstructions of a selected volume, comparing results obtained from inputs smoothed with db6 and sym8 wavelets.
+At intermediate ranks, the images sharpen and cortical detail appears while background fluctuations remain suppressed.
+We provide selected quantitative results in Table 1, which show this expected bias-variance tradeoff progression.
+Figure 1: Mid-sagittal MRI slices reconstructed with Algorithm 1 at Tucker ranks $(r,r,r)$ across noise levels $\lambda$ for a selected volume.
+Inputs smoothed with db6 (left) and sym8 (right) wavelets.
+Table 1: Mean and standard deviation (SD) of the relative reconstruction error with 55 MRI volumnes. Results are shown across varying scaling parameters $\lambda$ and ranks $r$.
+$\lambda$ $r$ original db6 sym8
+0.4 50 0.1753 (0.0330) 0.1747 (0.0537) 0.1681 (0.0466)
+65 0.1553 (0.0268) 0.1561 (0.0435) 0.1501 (0.0373)
+80 0.1490 (0.0183) 0.1499 (0.0291) 0.1451 (0.0245)
+95 0.1550 (0.0102) 0.1558 (0.0159) 0.1525 (0.0128)
+0.5 50 0.1797 (0.0322) 0.1793 (0.0521) 0.1728 (0.0451)
+65 0.1655 (0.0248) 0.1668 (0.0400) 0.1610 (0.0341)
+80 0.1677 (0.0154) 0.1691 (0.0244) 0.1647 (0.0201)
+95 0.1840 (0.0073) 0.1847 (0.0116) 0.1819 (0.0089)
+0.6 50 0.1851 (0.0311) 0.1850 (0.0503) 0.1787 (0.0434)
+65 0.1776 (0.0224) 0.1793 (0.0362) 0.1738 (0.0307)
+80 0.1890 (0.0124) 0.1904 (0.0200) 0.1866 (0.0162)
+95 0.2150 (0.0049) 0.2154 (0.0085) 0.2132 (0.0062)
 4.2 Synthetic data: Tensors
-We generate third-order tensors $X^{*}\in\mathbb{R}^{p\times p\times p}$ in the Tucker form
-$X^{*}\;=\;\mathcal{G}\times_{1}U_{1}\times_{2}U_{2}\times_{3}U_{3},\qquad U_{1},U_{2},U_{3}\in\mathbb{R}^{p\times s},\ U_{1}^{\top}U_{1}=U_{2}^{\top}U_{2}=U_{3}^{\top}U_{3}=I_{s},$
-where $U_{1},U_{2},U_{3}$ are drawn independently with orthonormal columns.
-The core $\mathcal{G}\in\mathbb{R}^{s\times s\times s}$ is diagonal along its main fiber with entries $\{\gamma_{i}\}_{i=1}^{s}$ decaying exponentially, i.e.
-$\gamma_{i}=\beta^{i},\,\beta=0.8$.
-We observe $Y=X^{*}+Z$ with i.i.d.
+We generate third-order tensors $X^{\ast}\in\mathbb{R}^{p\times p\times p}$ in the Tucker form
+$X^{\ast}=\mathcal{G}\times_{1}U_{1}\times_{2}U_{2}\times_{3}U_{3},\qquad U_{1},U_{2},U_{3}\in\mathbb{R}^{p\times p},\ U_{1}^{\top}U_{1}=U_{2}^{\top}U_{2}=U_{3}^{\top}U_{3}=I_{p},$
+where $U_{1},U_{2},U_{3}$ are drawn independently with orthonormal columns. The core tensor $\mathcal{G}\in\mathbb{R}^{p\times p\times p}$ is defined entrywise by
+$\mathcal{G}_{ijk}=\gamma_{ijk}\beta^{i+j+k},\qquad\beta=0.8,$
+where the signs $\gamma_{ijk}$ are independent with $\mathbb{P}(\gamma_{ijk}=-1)=1/3,\mathbb{P}(\gamma_{ijk}=1)=2/3$.
+We observe $Y=X^{\ast}+Z$ with i.i.d.
 Gaussian noise $Z_{ijk}\sim\mathcal{N}(0,1)$ entry-wise.
 The SNR is controlled by scaling the signal such that
-$\|X^{*}\|_{\mathrm{F}}\;=\;\lambda\,\sqrt{p^{3}},\qquad\lambda\in\{10,50\}.$
+$\|X^{\ast}\|_{\mathrm{F}}=\lambda\sqrt{p^{3}},\qquad\lambda\in\{10,50\}.$
 The estimator $\widetilde{X}$ is obtained via the one-step HOSVD (Algorithm 1) at the Tucker rank $(r,r,r)$.
-Dimensions and evaluated ranks. We vary $p\in\{20,50,75,100\}$ and for each $(p,s)$, we evaluate two ranks $r$ as listed below:
-$p$ $s$ $r$ (target)
-20 15 10, 12
-50 25 10, 15
-75 20 10, 15
-100 80 30, 40
-Table 1: Synthetic tensor grid: dimensions and evaluated Tucker ranks.
-Each configuration $(p,s,\lambda,r)$ is repeated $R=50$ times with freshly drawn $(U_{1},U_{2},U_{3})$ and noise.
-Implementations use exact batched SVD on GPU via TensorLy (PyTorch backend) (Kossaifi et al.,, 2019).
-We report the sample mean and standard error of the relative Frobenius error
-$\mathrm{RelErr}(\widetilde{X};X^{*})\;=\;\frac{\|\widetilde{X}-X^{*}\|_{\mathrm{F}}}{\|X^{*}\|_{\mathrm{F}}}\,$,
+Dimensions and evaluated ranks. We vary $p\in\{20,50,75,100\}$ and for each $p$, we evaluate two ranks $r$ as outlined in Table 2.
+Each configuration $(p,\lambda,r)$ is repeated $R=50$ times with freshly drawn $(U_{1},U_{2},U_{3})$ and noise.
+Implementations use exact batched SVD on GPU via TensorLy (PyTorch backend) (Kossaifi et al., 2019).
+We report the sample mean and standard deviation of the relative Frobenius error
+$\mathrm{RelErr}(\widetilde{X};X^{\ast})=\frac{\|\widetilde{X}-X^{\ast}\|_{\mathrm{F}}}{\|X^{\ast}\|_{\mathrm{F}}}\,$,
 summarized on the right panel of Table 2.
@@
 Overall, one-step HOSVD is robust across the tested sizes and ranks, yielding accurate estimates on the synthetic tensors.
+Bias–variance curve.
+To empirically quantify the contributions of bias and variance to the error $\|\widetilde{X}-X^{\ast}\|_{\mathrm{F}}$, we consider out synthetic tensor model with $p=100$.
+For each candidate Tucker rank $r\in\{12,15,18,\ldots,84,87\}$, we run $50$ independent repetitions.
+The bias component is obtained as the error $\|\widetilde{X}-X^{\ast}\|_{\mathrm{F}}$, where $\widetilde{X}$ denotes the output of Algorithm 1 when applied directly to the noiseless tensor $X^{\ast}$.
+The variance component is obtained as
+$\max\left\{\sqrt{r}\left\|\mathcal{M}_{1}(X^{\ast})\cdot\left\{U_{2}^{(1)}\otimes U_{3}^{(1)}\right\}\right\|,\,\sqrt{r}\left\|\mathcal{M}_{2}(X^{\ast})\cdot\left\{U_{1}^{(1)}\otimes U_{3}^{(1)}\right\}\right\|,\,\sqrt{r}\left\|\mathcal{M}_{3}(X^{\ast})\cdot\left\{U_{1}^{(1)}\otimes U_{2}^{(1)}\right\}\right\|,\,\left\|X^{\ast}\times_{1}U_{1}^{(1)}\times_{2}U_{2}^{(1)}\times_{3}U_{3}^{(1)}\right\|_{\mathrm{F}}\right\}.$
+Figure 2 reports the resulting mean bias and variance curves as functions of the target rank: as the target rank increases, the bias term decreases while the variance term grows, leading to an intermediate rank that minimizes the overall error.
+Figure 2: Mean bias and variance components of the error as functions of the target rank in the synthetic experiment with $p=100$.
+The left panel represent the curve with scaling parameter $\lambda=10$ and the right panel $\lambda=50$.
 4.3 Synthetic data: Matrices
 We examine the matrix analogue under a controlled SNR design that mirrors the tensor experiments. For each configuration, we draw an $m\times n$ latent signal
-$X^{*}\;=\;U\,\Sigma\,V^{\top},\qquad U\in\mathbb{R}^{m\times n},\ \ V\in\mathbb{R}^{n\times n},\ \ U^{\top}U=V^{\top}V=I_{n},$
+$X^{\ast}=U\,\Sigma\,V^{\top},\qquad U\in\mathbb{R}^{m\times n},\ \ V\in\mathbb{R}^{n\times n},\ \ U^{\top}U=V^{\top}V=I_{n},$
 where $\Sigma=\mathrm{diag}(\sigma_{1},\ldots,\sigma_{n})$ has exponentially decaying singular values, i.e.
 $\sigma_{i}=\beta^{\,i}$ with $\beta=0.8$.
-To mirror the tensor ambient sizes while probing rectangularity, we set $m=5p$ and $n=s$ according to the tensor grid in Table 1.
-We observe $Y=X^{\star}+Z$, where $Z_{ij}\sim\mathcal{N}(0,1)$ is i.i.d.
+We choose $m$ and $n$ as outlined in Table 2.
+We observe $Y=X^{\ast}+Z$, where $Z_{ij}\sim\mathcal{N}(0,1)$ is i.i.d.
 Gaussian noise entry-wise.
 The SNR is controlled by scaling the signal such that
-$\|X^{*}\|_{\mathrm{F}}\;=\;\lambda\,\sqrt{mn},\qquad\lambda\in\{10,50\}.$
+$\|X^{\ast}\|_{\mathrm{F}}=\lambda\sqrt{mn},\qquad\lambda\in\{10,50\}.$
 This design varies SNR via $\lambda$ while keeping the per–entry noise scale identical across sizes.
@@
 The truncated SVD is robust over a wide range of selected ranks.
-Table 2: Results for matrix (left) and tensor (right).
+Table 2: Mean and standard deviation (SD) of the relative error (RelErr) for matrices (left) and tensors (right).
+Results are shown across varying scaling parameters $\lambda$, dimensions ($m\times n$ for matrices; $p\times p\times p$ for tensors), and ranks $r$.
+For each configuration, results are averaged over $R=50$ Monte Carlo simulations.
 Matrix
-$\lambda$ $m$ $n$ $r$ Mean (SE)
+$\lambda$ $m$ $n$ $r$ Mean (SD)
 10 100 15 10 0.1178 (0.00125)
@@
 Tensor
-$\lambda$ $p$ $s$ $r$ Mean (SE)
-10 20 15 10 0.1092 (0.00031)
-12 0.0776 (0.00058)
-50 25 10 0.1081 (0.00002)
-15 0.0402 (0.00012)
-75 20 10 0.1081 (0.00002)
-15 0.0353 (0.00004)
-100 80 30 0.0204 (0.00007)
+$\lambda$ $p$ $r$ Mean (SD)
+10 20 10 0.1420 (0.00350)
+12 0.0977 (0.00208)
+50 10 0.1370 (0.00347)
+15 0.0459 (0.00102)
+75 10 0.1367 (0.00348)
+15 0.0427 (0.00113)
+100 30 0.0203 (0.00008)
 40 0.0294 (0.00009)
-50 20 15 10 0.1018 (0.00001)
-12 0.0599 (0.00002)
-50 25 10 0.1073 (0.00000)
-15 0.0352 (0.00000)
-75 20 10 0.1067 (0.00000)
-15 0.0333 (0.00000)
-100 80 30 0.0039 (0.00001)
+50 20 10 0.1361 (0.00381)
+12 0.0840 (0.00207)
+50 10 0.1362 (0.00359)
+15 0.0410 (0.00109)
+75 10 0.1368 (0.00374)
+15 0.0412 (0.00098)
+100 30 0.0039 (0.00002)
 40 0.0058 (0.00002)
@@
 Supplementary Materials for “Bias–variance Tradeoff in Tensor Estimation”
-Without loss of generality, if $W\sim\mathrm{subGaussian}(0,\kappa^{2})$, we assume throughout, unless otherwise specified in the appendix, that $\mathbb{E}[W^{2}]=\kappa^{2}$.
-Appendix A Proof of Theorem 1
+Throughout this appendix we work with centered sub-Gaussian random variables with parameter $\kappa^{2}$.
+we assume, unless otherwise specified, that each such variable $X$ satisfies $\mathbb{E}X^{2}=\kappa^{2}$.
+Appendix A Proofs in Section 2
+Proof of Theorem 1.
+By the triangle inequality,
+$\|Y_{(r)}-X^{*}\|_{\mathrm{F}}\leq\|X_{(r)}^{*}-X^{*}\|_{\mathrm{F}}+\|Y_{(r)}-X_{(r)}^{*}\|_{\mathrm{F}}=\xi_{(r)}+\|Y_{(r)}-X_{(r)}^{*}\|_{\mathrm{F}}.$ (10)
+Since $\mathrm{rank}(Y_{(r)}-X_{(r)}^{*})\leq 2r$, we have
+$\|Y_{(r)}-X_{(r)}^{*}\|_{\mathrm{F}}^{2}=\sum_{k=1}^{2r}\sigma_{k}^{2}(Y_{(r)}-X_{(r)}^{*}).$ (11)
+Write
+$Y_{(r)}-X_{(r)}^{*}=(X^{*}-X_{(r)}^{*})+Z+(Y_{(r)}-Y).$
+It follows from Ky Fan’s Theorem (Lemma 18) that,
+$\sqrt{\sum_{i=1}^{2r}\sigma_{i}^{2}(Y_{(r)}-X_{(r)}^{*})}\leq\sqrt{\sum_{i=1}^{2r}\sigma_{i}^{2}(X^{*}-X_{(r)}^{*})}+\sqrt{\sum_{i=1}^{2r}\sigma_{i}^{2}(Z)}+\sqrt{\sum_{i=1}^{2r}\sigma_{i}^{2}(Y_{(r)}-Y)}.$
+Using the identity $\sigma_{i}(X^{*}-X^{*}_{(r)})=\sigma_{r+i}(X^{*})$ and likewise for $Y$, we obtain
+$\|Y_{(r)}-X^{*}_{(r)}\|_{\mathrm{F}}$ $\leq\sqrt{\sum_{i=1}^{2r}\sigma_{r+i}^{2}(X^{*})}+\sqrt{\sum_{i=1}^{2r}\sigma_{i}^{2}(Z)}+\sqrt{\sum_{i=1}^{2r}\sigma_{r+i}^{2}(Y)}.$ (12)
+To further simplify the third term, we apply Weyl’s inequality (Theorem 6) to obtain
+$\sigma_{r+i}(Y)\leq\sigma_{r+i}(X^{*})+\sigma_{1}(Z),$
+which gives
+$\sigma_{r+i}^{2}(Y)\leq 2\left\{\sigma_{r+i}^{2}(X^{*})+\sigma_{1}^{2}(Z)\right\}.$
+Substituting this into (12), we get
+$\|Y_{(r)}-X^{*}_{(r)}\|_{\mathrm{F}}$ $\leq\sqrt{\sum_{i=1}^{2r}\sigma_{r+i}^{2}(X^{*})}+\sqrt{\sum_{i=1}^{2r}\sigma_{i}^{2}(Z)}+\sqrt{2\sum_{i=1}^{2r}\sigma_{r+i}^{2}(X^{*})}+2\sqrt{r}\sigma_{1}(Z)$
+$\leq(1+\sqrt{2})\sqrt{\sum_{i=1}^{2r}\sigma_{r+i}^{2}(X^{*})}+({2+\sqrt{2}})\sqrt{r}\|Z\|$
+$=(1+\sqrt{2})\sqrt{\sum_{i=r+1}^{3r}\sigma_{i}^{2}(X^{*})}+({2+\sqrt{2}})\sqrt{r}\|Z\|,$
+where the second inequality follows from the fact that for any $k\geq 1$, we have $\sigma_{k}(Z)\leq\sigma_{1}(Z)=\|Z\|.$ Substituting the bound of $\|Y_{(r)}-X^{*}_{(r)}\|_{\mathrm{F}}$ into (10), we obtain
+$\|Y_{(r)}-X^{*}\|_{\mathrm{F}}\leq(2+\sqrt{2})\sqrt{\sum_{i=r+1}^{\min\{m,n\}}\sigma^{2}_{i}(X^{*})}+(2+\sqrt{2})\sqrt{r}\|Z\|.$
+Proof of Corollary 1.
+By Lemma 8, $\|Z\|\leq C_{1}\sqrt{m+n}$ with probability at least $1-\exp(-(m+n))$. The desired result follows immediately from Theorem 1. ∎
+Proof of Theorem 2.
+Without loss of generality, suppose that $m\leq n$ and $r\leq m/2$. Therefore it suffices to show that
+$\inf_{\widehat{X}}\sup_{X^{*}\in\mathcal{M}_{r}^{\xi}}\ \mathbb{E}\bigl\|\widehat{X}-X^{*}\bigr\|_{\mathrm{F}}^{2}\$$\geq\ c_{1}\Bigl\{\kappa^{2}nr+\xi^{2}\Bigr\}.$ (13)
+Let $r^{\prime}=\lfloor\xi^{2}/(n\kappa^{2})\rfloor$. Note that since $\xi^{2}\leq(m-r)n\kappa^{2}$, it follows that
+$r^{\prime}+r\leq m.$
+Step 1. Consider the set of matrices
+$\mathcal{B}=\{B\in\mathbb{R}^{m\times n}:\|B\|_{\infty}\leq\kappa^{2},B_{i,j}=0\text{ if }i>r+r^{\prime}\}.$
+In this step, we show that
+$\mathcal{B}\subset\mathcal{M}^{\xi}_{r}.$ (14)
+To see this, let $B\in\mathcal{B}$, and denote $B^{\prime}$ to be the matrix such that the first $r$ rows of $B^{\prime}$ equals the first $r$ rows of $B$, and the rest of (mr)(m-r) rows of $B^{\prime}$ are all 0.
+Then since $B^{\prime}$ has at most $r$ rows that are nonzero, $\operatorname*{rank}(B^{\prime})\leq r$ and
+$\|B-B^{\prime}\|_{\mathrm{F}}^{2}\leq r^{\prime}n\|B\|_{\infty}=r^{\prime}n\kappa^{2}\leq\xi^{2}.$
+Therefore $\mathcal{B}\subset\mathcal{M}^{\xi}_{r}$. So
+$\inf_{\widehat{X}}\sup_{X^{*}\in\mathcal{M}_{r}^{\xi}}\ \mathbb{E}\bigl\|\widehat{X}-X^{*}\bigr\|_{\mathrm{F}}^{2}\ \geq\inf_{\widehat{X}}\sup_{X^{*}\in\mathcal{B}}\ \mathbb{E}\bigl\|\widehat{X}-X^{*}\bigr\|_{\mathrm{F}}^{2}.$ (15)
+Step 2.
+Note that matrices in $\mathcal{B}$ are having at most $(r+r^{\prime})n$ nonzero entries, the set of matrices $\mathcal{B}$ can be viewed as vectors in $\mathbb{R}^{d}$, with $d=(r+r^{\prime})n$.
+Denote
+$\mathcal{V}=\{V\in\mathbb{R}^{d},\|V\|_{\infty}\leq\kappa\}.$
+Then
+$\inf_{\widehat{X}}\sup_{X^{*}\in\mathcal{B}}\ \mathbb{E}\bigl\|\widehat{X}-X^{*}\bigr\|_{\mathrm{F}}^{2}=\inf_{\widehat{V}}\sup_{V\in\mathcal{V}}\mathbb{E}\|\widehat{V}-V\|^{2}.$
+Since by Lemma 19, $\inf_{\widehat{V}}\sup_{V\in\mathcal{V}}\mathbb{E}\|\widehat{V}-V\|^{2}\geq c_{2}\kappa^{2}d,$ it follows that
+$\inf_{\widehat{X}}\sup_{X^{*}\in\mathcal{M}_{r}^{\xi}}\mathbb{E}\bigl\|\widehat{X}-X^{*}\bigr\|_{\mathrm{F}}^{2}\geq\inf_{\widehat{X}}\sup_{X^{*}\in\mathcal{B}}\ \mathbb{E}\bigl\|\widehat{X}-X^{*}\bigr\|_{\mathrm{F}}^{2}=\inf_{\widehat{V}}\sup_{V\in\mathcal{V}}\mathbb{E}\|\widehat{V}-V\|^{2}\geq c_{2}\kappa^{2}d=c_{2}\kappa^{2}(r+r^{\prime})n.$
+Step 3. There are two cases. (a) Suppose $\xi^{2}\geq n\kappa^{2}$, then $r^{\prime}=\lfloor\xi^{2}/(n\kappa^{2})\rfloor\geq 1$ and therefore
+$\inf_{\widehat{X}}\sup_{X^{*}\in\mathcal{M}_{r}^{\xi}}\mathbb{E}\bigl\|\widehat{X}-X^{*}\bigr\|_{\mathrm{F}}^{2}\geq c_{2}\kappa^{2}(r+r^{\prime})n\geq c_{2}rn+c_{3}\xi^{2}.$
+(b) Suppose $\xi^{2}<n\kappa^{2}$. Then $r^{\prime}=0$
+$\inf_{\widehat{X}}\sup_{X^{*}\in\mathcal{M}_{r}^{\xi}}\ \mathbb{E}\bigl\|\widehat{X}-X^{*}\bigr\|_{\mathrm{F}}^{2}\$$\geq\ c_{2}\kappa^{2}rn\geq\frac{c_{2}}{2}(\kappa^{2}rn+\xi^{2}),$ (16)
+where the first inequality follows from Step 1 and Step 2, the second inequality follows from $\xi^{2}<n\kappa^{2}$. ∎
+Proof of Corollary 2.
+Let $\widetilde{Z}=Y-X^{*}.$ By Remark 4.7.3 of Vershynin (2018), with probability at least $1-\exp(-cn)$ one has $\|\widetilde{Z}\|\;\leq\;C_{1}\kappa^{2}\Bigl[\sqrt{\frac{n}{N}}+\frac{n}{N}\Bigr].$ Applying Theorem 1 with $Z$ replaced by $\widetilde{Z}$ yields the desired bound.
+∎
+Appendix B Proofs in Section 3
 Proof of Theorem 3.
@@
 $\leq\left\|I_{p_{1}}-\mathcal{P}_{U_{1}^{(1)}}\right\|\left\|\mathcal{M}_{1}(X^{*})\cdot(U_{2}^{*}\otimes U_{3}^{*})_{\perp}\right\|_{\mathrm{F}}\left\|(U_{2}^{*}\otimes U_{3}^{*})_{\perp}^{\top}\cdot(U_{2}^{(0)}\otimes U_{3}^{(0)})\right\|$
-$=\left\|\mathcal{M}_{1}(X^{*})\cdot(U_{2}^{*}\otimes U_{3}^{*})_{\perp}\right\|_{\mathrm{F}}.$
+$\leq\left\|\mathcal{M}_{1}(X^{*})\cdot(U_{2}^{*}\otimes U_{3}^{*})_{\perp}\right\|_{\mathrm{F}}.$
 Note that
@@
 In addition,
-$\|B-B_{r_{1}}\|_{F}=$ $\|\mathcal{M}_{1}(X^{*})\cdot U_{2}^{(0)}\otimes U_{3}^{(0)}-\{\mathcal{M}_{1}(X^{*})\cdot U_{2}^{(0)}\otimes U_{3}^{(0)}\}_{r_{1}}\|_{\mathrm{F}}$
+$\|B-B_{r_{1}}\|_{\mathrm{F}}=$ $\|\mathcal{M}_{1}(X^{*})\cdot U_{2}^{(0)}\otimes U_{3}^{(0)}-\{\mathcal{M}_{1}(X^{*})\cdot U_{2}^{(0)}\otimes U_{3}^{(0)}\}_{r_{1}}\|_{\mathrm{F}}$
 $\leq$ $\|\mathcal{M}_{1}(X^{*})\cdot U_{2}^{(0)}\otimes U_{3}^{(0)}-\{\mathcal{M}_{1}(X^{*})\}_{r_{1}}\cdot U_{2}^{(0)}\otimes U_{3}^{(0)}\|_{\mathrm{F}}$
@@
 $\left\|\widetilde{X}-X^{*}\right\|_{\mathrm{F}}\leq C_{7}\kappa(\sqrt{p_{1}r_{1}}+\sqrt{p_{2}r_{2}}+\sqrt{p_{3}r_{3}}+\sqrt{r_{1}r_{2}r_{3}})+C_{7}\xi_{(r_{1},r_{2},r_{3})}.$
-Appendix B Deviation Bounds
-Throughout this appendix we work with centered sub-Gaussian random variables with parameter $\kappa^{2}$, and we assume (without loss of generality) that each such variable $X$ satisfies $\mathbb{E}X^{2}=\kappa^{2}$.
-Any other case can be handled with additional constants in the bounds.
-Lemma 6 (Theorem 4.4.3 in Vershynin, (2018)).
-Assume all the entries of $Z\in\mathbb{R}^{m\times n}$ are independent mean-zero sub-Gaussian random variables, i.e.
-$\|Z_{ij}\|_{\psi_{2}}=\sup_{q\geq 1}\mathbb{E}(|Z_{ij}|^{q})^{1/q}/q^{1/2}\leq\kappa.$
-Then there exist some universal constant $C>0$, such that
-$\|Z\|\leq C\kappa\left(\sqrt{m}+\sqrt{n}\right)$
-with probability at least $1-\exp(-(m+n)))$.
-Lemma 7.
-Suppose all the entries of $Z\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}}$ are independent mean-zero sub-Gaussian random variables, i.e.
-$\|Z_{ijk}\|_{\psi_{2}}=\sup_{q\geq 1}\mathbb{E}(|Z_{ijk}|^{q})^{1/q}/q^{1/2}\leq\kappa.$
-Then there exist some universal constants $C,c>0$, such that
-$\sup_{\begin{subarray}{c}A\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}},\\ \|A\|_{\mathrm{F}}\leq 1,A\in\mathcal{T}_{(r_{1},r_{2},r_{3})}\end{subarray}}\langle Z,A\rangle\leq C\kappa\left(r_{1}r_{2}r_{3}+\sum_{k=1}^{3}p_{k}r_{k}\right)^{1/2}$
-with probability at least $1-\exp(-c\sum_{k=1}^{3}p_{k}r_{k})$.
-Proof.
-This directly follows from Lemma E.5 in Han et al., (2022). ∎
-Lemma 8.
-Suppose all the entries of $Z\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}}$ are independent mean-zero sub-Gaussian random variables, i.e.
-$\|Z_{ijk}\|_{\psi_{2}}=\sup_{q\geq 1}\mathbb{E}(|Z_{ijk}|^{q})^{1/q}/q^{1/2}\leq\kappa.$
-Let $W_{2}\in\mathbb{O}^{p_{2}\times s_{2}}$ and $W_{3}\in\mathbb{O}^{p_{3}\times s_{3}}$ be non-random. Then there exists absolute positive constants $C_{1},C_{2}$ and $c$ such that
-$\mathbb{P}\bigg(\big\|\mathcal{M}_{1}(Z)\,(W_{2}\otimes W_{3})\big\|\geq C_{1}\kappa(\sqrt{p_{1}+s_{2}s_{3}})\bigg)\leq C_{2}\exp\left(-cp_{1}\right),$
-Proof.
-It suffices to observe that $W_{2}\otimes W_{3}\in\mathbb{O}^{p_{2}p_{3}\times r_{2}r_{3}}$. The desired result is a direct consequence of Lemma 10. ∎
-Lemma 9.
-Suppose all the entries of $Z\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}}$ are independent mean-zero sub-Gaussian random variables, i.e.
-$\|Z_{ijk}\|_{\psi_{2}}=\sup_{q\geq 1}\mathbb{E}(|Z_{ijk}|^{q})^{1/q}/q^{1/2}\leq\kappa.$
-Then there exists absolute positive constants $C_{1},C_{2}$ and $c$ such that
-$\mathbb{P}\bigg(\sup_{\begin{subarray}{c}V_{2}\in\mathbb{R}^{p_{2}\times r_{2}},\,\|V_{2}\|\leq 1\\ V_{3}\in\mathbb{R}^{p_{3}\times r_{3}},\,\|V_{3}\|\leq 1\end{subarray}}\big\|\mathcal{M}_{1}(Z)\,(V_{2}\otimes V_{3})\big\|\geq C_{1}\kappa\Big(\sqrt{p_{1}+r_{2}r_{3}+p_{2}r_{2}+p_{3}r_{3}}\Bigg)$
-$\leq C_{2}\exp\left(-c(p_{1}+p_{2}+p_{3})\right),$
-Proof.
-It follows from the assumption that $\left(\mathcal{M}_{1}(Z)\right)_{i,j}\overset{i.i.d}{\sim}\mathrm{subGaussian}(0,\kappa^{2})$.
-Step 1. For fixed $V_{2}\in\mathbb{R}^{p_{2}\times r_{2}}$ with $\|V_{2}\|\leq 1$ and $V_{3}\in\mathbb{R}^{p_{3}\times r_{3}}$ with $\|V_{3}\|\leq 1$, it follows that
-$\left\|V_{2}\otimes V_{3}\right\|=\left\|V_{2}\right\|\left\|V_{3}\right\|\leq 1,$
-We upper bound $\|\mathcal{M}_{1}(Z)\cdot(V_{2}\otimes V_{3})\|$ for any fixed $V_{2}$ and $V_{3}$. Since,
-$\mathcal{M}(Z)\in\mathbb{R}^{p_{1}\times p_{2}p_{3}},\left(\mathcal{M}(Z)\right)_{i,j}\overset{i.i.d}{\sim}\mathrm{subGaussian}(0,\kappa^{2}),\quad\text{and}\quad\left\|V_{2}\otimes V_{3}\right\|=\left\|V_{2}\right\|\left\|V_{3}\right\|=1.$
-It follows from Lemma 10 that
-$\mathbb{P}\Big(\left\|\mathcal{M}_{1}(Z)\cdot(V_{2}\otimes V_{3})\right\|>x\Big)\leq 2\exp\left(C(p_{1}+r_{2}r_{3})-cx^{2}\kappa^{-2}\right).$
-Step 2. Let $\mathcal{N}_{p_{2},r_{2}}(\epsilon)$ denote an $\epsilon$ net of the set
-$\{A\in\mathbb{R}^{p_{2}\times r_{2}}:\|A\|\leq 1\}$
-with respect to the operator norm $\|\cdot\|$. Similarly, let Let $\mathcal{N}_{p_{3},r_{3}}(\epsilon)$ denote an $\epsilon$ net of the set
-$\{B\in\mathbb{R}^{p_{3}\times r_{3}}:\|B\|\leq 1\}$
-with respect to the operator norm $\|\cdot\|$. Denote the random quantity
-$\psi=\sup_{\begin{subarray}{c}V_{2}\in\mathbb{R}^{p_{2}\times r_{2}},\,\|V_{2}\|\leq 1\\ V_{3}\in\mathbb{R}^{p_{3}\times r_{3}},\,\|V_{3}\|\leq 1\end{subarray}}\big\|\mathcal{M}_{1}(Z)\,(V_{2}\otimes V_{3})\big\|.$
-For any given $V_{2}\in\mathbb{R}^{p_{2}\times r_{2}}$ and $V_{3}\in\mathbb{R}^{q}$ with $\|V_{2}\|\leq 1$ and $\|V_{3}\|\leq 1$, let $\widetilde{V}_{2}\in\mathcal{N}_{p_{2},r_{2}}(1/4)$ and $\widetilde{V}_{3}\in\mathcal{N}_{p_{3},r_{3}}(1/4)$ be such that
-$\|V_{2}-\widetilde{V}_{2}\|\leq 1/4\quad\text{and}\quad\|V_{3}-\widetilde{V}_{3}\|\leq 1/4.$
+Proof of Corollary 3.
+It suffices to observe that for $k=1,2,3$
+$\|X^{*}\|_{\mathrm{F}}^{2}\geq r_{k}\sigma_{r_{k}}^{2}(\mathcal{M}_{k}(X^{*})).$
+The desired follows from the assumption that $r_{k}\leq\sqrt{p_{\min}}$ and (7). ∎
+Proof of Theorem 4.
+Step 1. We follow a similar strategy as in the proof of Theorem 2 to show
+$\inf_{\widehat{X}}\sup_{X^{\ast}\in\mathcal{T}_{(r_{1},r_{2},r_{3})}^{\xi}}\ \mathbb{E}\bigl\|\widehat{X}-X^{\ast}\bigr\|_{\mathrm{F}}^{2}\$$\geq\ c(\kappa^{2}r_{1}r_{2}r_{3}+\xi^{2}).$ (27)
+Let
+$r_{1}^{\prime}=\lfloor\xi^{2}/(r_{2}r_{3}\kappa^{2})\rfloor.$
+Consider the set of tensors
+$\mathcal{B}=\{B\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}}:\|B\|_{\infty}\leq\kappa^{2},B_{i,j,k}=0\text{ if }i>r_{1}+r_{1}^{\prime}\text{ or }j>r_{2}\text{ or }k>r_{3}\}.$
+In this step, we show that
+$\mathcal{B}\subset\mathcal{T}_{(r_{1},r_{2},r_{3})}^{\xi}.$ (28)
+To see this, let $B\in\mathcal{B}$, and denote $B^{\prime}$ to be the matrix such that
+$B^{\prime}_{i,j,k}=\begin{cases}B_{i,j,k}&\text{if }i\leq r_{1}\text{ and }j\leq r_{2}\text{ and }k\leq r_{3};\\ 0&\text{otherwise}.\end{cases}$
+Then $B^{\prime}\in\mathcal{T}_{(r_{1},r_{2},r_{3})}$ and
+$\|B-B^{\prime}\|_{\mathrm{F}}^{2}\leq r^{\prime}_{1}r_{2}r_{3}\|B\|_{\infty}=r^{\prime}_{1}r_{2}r_{3}\kappa^{2}\leq\xi^{2}.$
+Therefore $\mathcal{B}\subset\mathcal{T}_{(r_{1},r_{2},r_{3})}^{\xi}$. So
+$\inf_{\widehat{X}}\sup_{X^{*}\in\mathcal{T}_{(r_{1},r_{2},r_{3})}^{\xi}}\ \mathbb{E}\bigl\|\widehat{X}-X^{*}\bigr\|_{\mathrm{F}}^{2}\ \geq\inf_{\widehat{X}}\sup_{X^{*}\in\mathcal{B}}\ \mathbb{E}\bigl\|\widehat{X}-X^{*}\bigr\|_{\mathrm{F}}^{2}.$ (29)
+Step 1.1.
+Note that the tensors in $\mathcal{B}$ are having at most $(r_{1}+r_{1}^{\prime})r_{2}r_{3}$ nonzero entries, the set of matrices $\mathcal{B}$ can be viewed as vectors in $\mathbb{R}^{d}$, with $d=(r_{1}+r_{1}^{\prime})r_{2}r_{3}$.
+Denote
+$\mathcal{V}=\{V\in\mathbb{R}^{d},\|V\|_{\infty}\leq\kappa\}.$
 Then
-$\|\mathcal{M}_{1}(Z)V_{2}\otimes V_{3}\|\leq\|\mathcal{M}_{1}(Z)(V_{2}-\widetilde{V}_{2})\otimes V_{3}\|+\|\mathcal{M}_{1}(Z)\widetilde{V}_{2}\otimes(V_{3}-\widetilde{V}_{3})\|+\|\mathcal{M}_{1}(Z)\widetilde{V}_{2}\otimes\widetilde{V}_{3}\|.$
-Note that $\|V_{2}-\widetilde{V}_{2}\|\leq 1/4$. So
-$\|\mathcal{M}_{1}(Z)(V_{2}-\widetilde{V}_{2})\otimes V_{3}\|=\frac{1}{4}\|\mathcal{M}_{1}(Z)\{4(V_{2}-\widetilde{V}_{2})\}\otimes V_{3}\|\leq\frac{\psi}{4}.$
-Similarly
-$\|\mathcal{M}_{1}(Z)V_{2}\otimes(V_{3}-\widetilde{V}_{3})\|\leq\frac{\psi}{4}.$
-In addition,
-$\|\mathcal{M}_{1}(Z)\widetilde{V}_{2}\otimes\widetilde{V}_{3}\|\leq\sup_{V_{2}\in\mathcal{N}_{p_{2},r_{2}}(1/4),b\in\mathcal{N}_{p_{3},r_{3}}(1/4)}\|\mathcal{M}_{1}(Z)V_{2}\otimes V_{3}\|.$
-So for any $V_{2}$ and $V_{3}$,
-$\|\mathcal{M}_{1}(Z)V_{2}\otimes V_{3}\|\leq\frac{1}{2}\psi+\sup_{V_{2}\in\mathcal{N}_{p_{2},r_{2}}(1/4),b\in\mathcal{N}_{p_{3},r_{3}}(1/4)}\|\mathcal{M}_{1}(Z)V_{2}\otimes V_{3}\|.$
-Taking sup over all $V_{2}\in\{A\in\mathbb{R}^{p_{2}\times r_{2}}:\|A\|\leq 1\}$ and $V_{3}\in\{B\in\mathbb{R}^{p_{3}\times r_{3}}:\|B\|\leq 1\}$, it follows that
-$\psi\leq\frac{1}{2}\psi+\sup_{V_{2}\in\mathcal{N}_{p_{2},r_{2}}(1/4),b\in\mathcal{N}_{p_{3},r_{3}}(1/4)}\|\mathcal{M}_{1}(Z)V_{2}\otimes V_{3}\|,$
-or simply
-$\psi\leq 2\sup_{V_{2}\in\mathcal{N}_{p_{2},r_{2}}(1/4),b\in\mathcal{N}_{p_{3},r_{3}}(1/4)}\|\mathcal{M}_{1}(Z)V_{2}\otimes V_{3}\|.$
-Step 3.
-By Proposition 8 in Pajor, (1998), the cardinality of $\mathcal{N}_{p_{2},r_{2}}(\epsilon)$ is bounded $(\frac{C}{\epsilon})^{p_{2}r_{2}}$, and $\mathcal{N}_{p_{3},r_{3}}(\epsilon)$ is bounded $(\frac{C}{\epsilon})^{p_{3}r_{3}}$.
-Therefore
-$\mathbb{P}(\psi\geq 2t)\leq$ $\mathbb{P}\bigg(\sup_{V_{2}\in\mathcal{N}_{p_{2},r_{2}}(1/4),b\in\mathcal{N}_{p_{3},r_{3}}(1/4)}\|\mathcal{M}_{1}(Z)V_{2}\otimes V_{3}\|\geq t\bigg)$
-$\leq$ $C_{2}^{p_{2}r_{2}}C_{3}^{p_{3}r_{3}}\sup_{V_{2}\in\mathcal{N}_{p_{2},r_{2}}(1/4),b\in\mathcal{N}_{p_{3},r_{3}}(1/4)}\mathbb{P}\bigg(\|\mathcal{M}_{1}(Z)V_{2}\otimes V_{3}\|\geq t\bigg)$
-$\leq$ $2\exp\left(C(p_{1}+r_{2}r_{3}+p_{2}r_{2}+p_{3}r_{3})-ct^{2}\kappa^{-2}\right).$
-Here $C$ and $C_{3}$ are positive constants. The desired result follows by noting
-$\psi=\sup_{\begin{subarray}{c}V_{2}\in\mathbb{R}^{p_{2}\times r_{2}},\,\|V_{2}\|\leq 1\\ V_{3}\in\mathbb{R}^{p_{3}\times r_{3}},\,\|V_{3}\|\leq 1\end{subarray}}\big\|\mathcal{M}_{1}(Z)\,(V_{2}\otimes V_{3})\big\|.$
-Lemma 10.
-Suppose $Z\in\mathbb{R}^{n\times m}$, with $Z_{ij}\overset{i.i.d}{\sim}\mathrm{subGaussian}(0,\kappa^{2})$.
-Let $A\in\mathbb{R}^{p\times n}$ and $B\in\mathbb{R}^{m\times q}$ be non-random matrices.
-Then for any $t>0$
-$\mathbb{P}\left(\left\|A\,Z\,B\right\|>t\right)\leq C_{1}\,\exp\left(C(p+q)-\frac{c\,t^{2}}{\kappa^{2}\,\left\|A\right\|^{2}\,\left\|B\right\|^{2}}\right).$ (20)
-Proof.
-Step 1.
-Let $u\in\mathbb{R}^{n}$ and $v\in\mathbb{R}^{m}$ be non-random.
-Then that $u^{\top}\,Z\,v=\sum_{i=1}^{n}\sum_{j=1}^{n}u_{i}Z_{ij}v_{j}$.
-Since $Z_{ij}$ are i.i.d.
-sub-Gaussian with parameter $\kappa^{2}$, it follows that $u^{\top}\,Z\,v$ is sub-Gaussian with parameter $\kappa^{2}\|u\|_{2}^{2}\|v\|_{2}^{2}$.
-Consequently by Hoeffding’s inequality,
-$\mathbb{P}\left(\left|u^{\top}\,Z\,v\right|\,>t\,\right)\leq 2\,\exp\left(-\frac{c\,t^{2}}{\kappa^{2}\,\left\|u\right\|_{2}^{2}\,\left\|v\right\|_{2}^{2}}\right).$
-Step 2. Let $a\in\mathbb{R}^{p}$ and $b\in\mathbb{R}^{q}$ be non-random vectors such that $\|a\|_{2},\|b\|_{2}\leq 1$. By Step 1, it follows that
-$\mathbb{P}\left(\left|a^{\top}AZ\,Bb\right|\,>t\,\right)\leq 2\,\exp\left(-\frac{ct^{2}}{\kappa^{2}\,\left\|Aa\right\|_{2}^{2}\,\left\|Bb\right\|_{2}^{2}}\right)\leq 2\exp\left(-\frac{ct^{2}}{\kappa^{2}\,\left\|A\right\|^{2}\,\left\|B\right\|^{2}}\right).$
-Step 3.
-Let $\mathcal{N}_{p}(\epsilon)$ be an $\epsilon$-net of the unit ball in $\mathbb{R}^{p}$.
-It follows that for any $a\in\mathbb{R}^{p}$ with $\|a\|_{2}=1$, there exists $\widetilde{a}\in\mathcal{N}_{p}(\epsilon)$ such that
-$\|a-\widetilde{a}\|_{2}\leq\epsilon.$
-Similarly let $\mathcal{N}_{q}(\epsilon)$ be an $\epsilon$-net of the unit ball in $\mathbb{R}^{q}$.
-Denote the random quantity
-$\psi=\|AZB\|=\sup_{a\in\mathbb{R}^{p},b\in\mathbb{R}^{q}\|a\|_{2}=\|b\|_{2}=1}|a^{\top}AZBb|.$
-For any given $a\in\mathbb{R}^{p}$ and $b\in\mathbb{R}^{q}$, let $\widetilde{a}\in\mathcal{N}_{p}(1/4)$ and $\widetilde{b}\in\mathcal{N}_{q}(1/4)$ be such that
-$\|a-\widetilde{a}\|_{2}\leq 1/4\quad\text{and}\quad\|b-\widetilde{b}\|_{2}\leq 1/4.$
-Then
-[…]
-Note that $\|a-\widetilde{a}\|\leq\frac{1}{4}$. So
-$|(a-\widetilde{a})^{\top}AZBb|=\frac{1}{4}|\{4(a-\widetilde{a})^{\top}\}AZBb|\leq\frac{\psi}{4}.$
-Similarly
-$|\widetilde{a}^{\top}AZB(\widetilde{b}-b)|\leq\frac{\psi}{4}.$
-In addition,
-$|\widetilde{a}^{\top}AZB\widetilde{b}|\leq\sup_{a\in\mathcal{N}_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}|a^{\top}AZBb|.$
-So for any $a$ and $b$,
-$|a^{\top}AZBb|\leq\frac{1}{2}\psi+\sup_{a\in N_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}|a^{\top}AZBb|.$
-Taking sup over all unit vectors $a\in\mathbb{R}^{p}$ and $b\in\mathbb{R}^{q}$, it follows that
-$\psi\leq\frac{1}{2}\psi+\sup_{a\in\mathcal{N}_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}|a^{\top}AZBb|,$
-or simply
-$\psi\leq 2\sup_{a\in\mathcal{N}_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}|a^{\top}AZBb|.$
-Step 4.
-By Vershynin, (2018), the cardinality of $\mathcal{N}_{p}(\epsilon)$ is bounded by $(\frac{C}{\epsilon})^{p}$, and the cardinality of $\mathcal{N}_{q}(\epsilon)$ is bounded by $(\frac{C}{\epsilon})^{q}$.
-Therefore
-$\mathbb{P}(\psi\geq 2t)=\mathbb{P}\bigg(\sup_{a\in\mathcal{N}_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}|a^{\top}AZBb|\geq t\bigg)\leq$ ${C_{2}}^{p}{C_{2}}^{q}\sup_{a\in\mathcal{N}_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}\mathbb{P}(|a^{\top}AZBb|\geq t)$
-$\leq$ $2{C_{2}}^{p}{C_{2}}^{q}\exp\left(-\frac{ct^{2}}{\kappa^{2}\,\left\|A\right\|^{2}\,\left\|B\right\|^{2}}\right),$
-where $c,C$ are positive constants. The desired result follows by noting $\psi=\|AZB\|$. ∎
-Lemma 11.
-Suppose $Z\in\mathbb{R}^{n\times m}$, with $Z_{ij}\overset{i.i.d}{\sim}\mathrm{subGaussian}(0,\kappa^{2})$.
-Let $A\in\mathbb{R}^{m\times p}$ and $B\in\mathbb{R}^{m\times q}$ be non-random matrices.
-Then for any $t>0$
-$\mathbb{P}\left(\left\|A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B\right\|>t\right)\leq C_{1}\,\exp\left(C(p+q)-\min\left(\frac{t^{2}}{n\,\kappa^{4}\,\left\|B\right\|^{2}\,\left\|A\right\|^{2}},\,\frac{t}{\kappa^{2}\left\|B\right\|\,\left\|A\right\|}\right)\right),$ (21)
-where $C$ and $C_{1}$ are positive constants.
-Proof.
-For any non-random $u,v\in\mathbb{R}^{m}$ , it follows that
-$u^{\top}Z^{\top}Zv-n\kappa^{2}u^{\top}v=\sum_{j=1}^{n}(u^{\top}Z_{j})(v^{\top}Z_{j})-\mathbb{E}\{(u^{\top}Z_{j})(v^{\top}Z_{j})\},$
-where $Z_{j}$ is the $j$-th row of $Z$.
-Note that $(u^{\top}Z_{j})$ is sub-Gaussian with parameter $\kappa^{2}\|u\|_{2}^{2}$, and $(v^{\top}Z_{j})$ is sub-Gaussian with parameter $\kappa^{2}\|v\|_{2}^{2}$.
-Since $Z$ have i.i.d.
-entries, it follows that $\{(u^{\top}Z_{j})(v^{\top}Z_{j})\}_{j=1}^{n}$ are i.i.d.
-sub-exponential with parameter $\kappa^{4}\|u\|_{2}^{2}\|v\|_{2}^{2}$.
-So
-$\mathbb{P}\bigg(|u^{\top}Z^{\top}Zv-n\kappa^{2}u^{\top}v|\geq t\bigg)\leq 2\exp\bigg(-c\min\bigg\{\frac{t^{2}}{n\kappa^{4}\|u\|_{2}^{2}\|v\|_{2}^{2}},\frac{t}{\kappa^{2}\|u\|_{2}\|v\|_{2}}\bigg\}\bigg).$
-Step 1.
-Let $\mathcal{N}_{p}(\epsilon)$ be the $\epsilon$-net of the unit ball in $\mathbb{R}^{p}$.
-It follows that for any $a\in\mathbb{R}^{p}$ with $\|a\|_{2}=1$, there exists $\widetilde{a}\in\mathcal{N}_{p}(\epsilon)$ such that
-$\|a-\widetilde{a}\|_{2}\leq\epsilon.$
-Similarly let $\mathcal{N}_{q}(\epsilon)$ be the $\epsilon$-net of the unit ball in $\mathbb{R}^{q}$.
-Denote the random quantity
-$\psi=\|A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B\|=\sup_{a\in\mathbb{R}^{p},b\in\mathbb{R}^{q}\|a\|_{2}=\|b\|_{2}=1}\bigl|a^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b\bigr|.$
-For any given $a\in\mathbb{R}^{p}$ and $b\in\mathbb{R}^{q}$, let $\widetilde{a}\in\mathcal{N}_{p}(1/4)$ and $\widetilde{b}\in\mathcal{N}_{q}(1/4)$ be such that
-$\|a-\widetilde{a}\|_{2}\leq 1/4\quad\text{and}\quad\|b-\widetilde{b}\|_{2}\leq 1/4.$
-Then
-$|a^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b|\leq|(a-\widetilde{a})^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b|$
-[…]
-Note that $\|a-\widetilde{a}\|_{2}\leq\frac{1}{4}$. So
-$|(a-\widetilde{a})^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b|=\frac{1}{4}|\{4(a-\widetilde{a})^{\top}\}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b|\leq\frac{\psi}{4}.$
-Similarly
-$|\widetilde{a}^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)(\widetilde{b}-b)|\leq\frac{\psi}{4}.$
-In addition,
-$|\widetilde{a}^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)\widetilde{b}|\leq\sup_{a\in\mathcal{N}_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}|a^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b|.$
-So for any $a$ and $b$,
-$|a^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b|\leq\frac{1}{2}\psi+\sup_{a\in N_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}|a^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b|.$
-Taking sup over all unit vectors $a\in\mathbb{R}^{p}$ and $b\in\mathbb{R}^{q}$, it follows that
-$\psi\leq\frac{1}{2}\psi+\sup_{a\in\mathcal{N}_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}|a^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b|,$
-or simply
-$\psi\leq 2\sup_{a\in\mathcal{N}_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}|a^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b|.$
-Step 2.
-By Vershynin, (2018), the cardinality of $\mathcal{N}_{p}(\epsilon)$ is bounded by $(\frac{C}{\epsilon})^{p}$, and the cardinality of $\mathcal{N}_{q}(\epsilon)$ is bounded by $(\frac{C}{\epsilon})^{q}$, for a positive constant $C$.
-Therefore
-$\mathbb{P}(\psi\geq 2t)=$ $\mathbb{P}\bigg(\sup_{a\in\mathcal{N}_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}|a^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b|\geq t\bigg)$
-$\leq$ ${C_{2}}^{p}{C_{2}}^{q}\sup_{a\in\mathcal{N}_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}\mathbb{P}(|a^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b|\geq t)$
-$=$ ${C_{2}}^{p}{C_{2}}^{q}\sup_{a\in\mathcal{N}_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}\mathbb{P}(|(Aa)^{\top}Z^{\top}Z(Bb)-n\kappa^{2}(Aa)^{\top}(Bb)|\geq t)$
-$\leq$ $2{C_{2}}^{p}{C_{2}}^{q}\exp\bigg(-c\min\bigg\{\frac{t^{2}}{n\kappa^{4}\|Aa\|_{2}^{2}\|Bb\|_{2}^{2}},\frac{t}{\kappa^{2}\|Aa\|_{2}\|Bb\|_{2}}\bigg\}\bigg),$
-where $C_{2}$ is a positive constant. The desired result follows from the observation that $\|Aa\|_{2}\leq\|A\|\|a\|_{2}\leq\|A\|$, and $\|Bb\|_{2}\leq\|B\|\|b\|_{2}\leq\|B\|$. ∎
-Lemma 12.
-Suppose $Z\in\mathbb{R}^{m\times n}$ is a sub-Gaussian random matrix in the sense that for any $u\in\mathbb{R}^{m},v\in\mathbb{R}^{n}$, it holds that
-$\|u^{\top}Zv\|_{\psi_{2}}\leq\kappa\|u\|_{2}\|v\|_{2}.$
-Then with probability at least $1-\exp(-c(m+n))$, it holds that
-$\mathbb{P}\left(\left\|Z\right\|>t\right)\leq C_{1}\,\exp\left(C(m+n)-\frac{c\,t^{2}}{\kappa^{2}\,}\right),$
-where $c,C$ and $C_{1}$ are positive constants.
-Proof.
-By assumption,
-$\mathbb{P}\left(\left|u^{\top}\,Z\,v\right|\,>t\,\right)\leq 2\,\exp\left(-\frac{c\,t^{2}}{\kappa^{2}\,\left\|u\right\|_{2}^{2}\,\left\|v\right\|_{2}^{2}}\right).$
-The rest of the proof is similar and simpler than Lemma 10 and is omitted.
-B.1 SVD for Unbalanced Matrices
+$\inf_{\widehat{X}}\sup_{X^{*}\in\mathcal{B}}\ \mathbb{E}\bigl\|\widehat{X}-X^{*}\bigr\|_{\mathrm{F}}^{2}=\inf_{\widehat{V}}\sup_{V\in\mathcal{V}}\mathbb{E}\|\widehat{V}-V\|^{2}.$
+Since by Lemma 19, $\inf_{\widehat{V}}\sup_{V\in\mathcal{V}}\mathbb{E}\|\widehat{V}-V\|^{2}\geq c_{2}\kappa^{2}d,$ it follows that
+$\inf_{\widehat{X}}\sup_{X^{*}\in\mathcal{B}}\ \mathbb{E}\bigl\|\widehat{X}-X^{*}\bigr\|_{\mathrm{F}}^{2}=\inf_{\widehat{V}}\sup_{V\in\mathcal{V}}\mathbb{E}\|\widehat{V}-V\|^{2}\geq c_{2}\kappa^{2}d=c_{2}\kappa^{2}(r_{1}+r_{1}^{\prime})r_{2}r_{3}.$
+Step 1.2. To show (27), there are two cases. (a) Suppose $\xi^{2}\geq r_{2}r_{3}\kappa^{2}$, then $r_{1}^{\prime}=\lfloor\xi^{2}/(r_{2}r_{3}\kappa^{2})\rfloor\geq 1$ and therefore
+$\inf_{\widehat{X}}\sup_{X^{*}\in\mathcal{T}_{(r_{1},r_{2},r_{3})}^{\xi}}\mathbb{E}\bigl\|\widehat{X}-X^{*}\bigr\|_{\mathrm{F}}^{2}\geq c_{2}\kappa^{2}(r_{1}+r_{1}^{\prime})r_{2}r_{3}\geq c_{2}r_{1}r_{2}r_{3}+c_{3}\xi^{2}.$
+(b) Suppose $\xi^{2}<r_{2}r_{3}\kappa^{2}$. Then $r^{\prime}_{1}=0$
+$\inf_{\widehat{X}}\sup_{X^{*}\in\mathcal{T}_{(r_{1},r_{2},r_{3})}^{\xi}}\ \mathbb{E}\bigl\|\widehat{X}-X^{*}\bigr\|_{\mathrm{F}}^{2}\$$\geq\ c_{2}\kappa^{2}r_{1}r_{2}r_{3}\geq\frac{c_{2}}{2}(\kappa^{2}r_{1}r_{2}r_{3}+\xi^{2}),$
+where the first inequality follows from Step 1.1, the second inequality follows from $\xi^{2}<r_{2}r_{3}\kappa^{2}$.
+Step 2. Note that
+$\inf_{\widehat{X}}\ \sup_{X^{*}\in\mathcal{T}_{(r_{1},r_{2},r_{3})}^{\xi}}\ \mathbb{E}\bigl\|\widehat{X}-X^{*}\bigr\|_{\mathrm{F}}^{2}\geq\inf_{\widehat{X}}\ \sup_{X^{*}\in\mathcal{T}_{(r_{1},r_{2},r_{3})}}\ \mathbb{E}\bigl\|\widehat{X}-X^{*}\bigr\|_{\mathrm{F}}^{2}\ \geq\ c_{3}\kappa^{2}\sum_{k=1}^{3}p_{k}r_{k},$ (30)
+where the second inequality follows from Zhang and Xia (2018). The desired result directly follows from (27) and (30). ∎
+Appendix C SVD for Unbalanced Matrices
 Lemma 1.
@@
 $\begin{split}&\mathbb{P}\Big(\left\|M^{\top}V_{r}^{\top}\left(Z^{\top}Z-n\kappa^{2}I_{m}\right)V_{\perp}\right\|\geq x\Big)\\ &\leq 2\,\exp\left(Cm-c\min\left\{x^{2}\,\frac{\sigma_{r}^{2}(X)+n\kappa^{2}}{n\kappa^{4}},\,x\,\frac{\sqrt{\sigma_{r}^{2}(X)+n\kappa^{2}}}{\kappa^{2}}\right\}\right),\\ &\leq 2\,\exp\left(Cm-c\min\left\{\frac{x^{2}}{\kappa^{2}},\,x\,\frac{\sqrt{\sigma_{r}^{2}(X)+n\kappa^{2}}}{\kappa^{2}}\right\}\right),\end{split}$ (46)
-where the last inequality uses $\frac{\sigma_{R}^{2}(X)+n{\kappa}^{2}}{n\,\kappa^{2}}\geq 1$. Thus, combining (B.1), (35), (36) and (37) with $x=\kappa\sqrt{\frac{2C}{c}m\,}$, we have
+where the last inequality uses $\frac{\sigma_{R}^{2}(X)+n{\kappa}^{2}}{n\,\kappa^{2}}\geq 1$. Thus, combining (C), (44), (45) and (46) with $x=\kappa\sqrt{\frac{2C}{c}m\,}$, we have
 $\mathbb{P}\!\left(\left\|\mathcal{P}_{YV_{r}}YV_{\perp}\right\|^{2}\geq\frac{36C}{c}\,m\kappa^{2}\right)$ $\leq 6\exp\!\left(Cr-c\frac{\sigma_{r}^{2}(X)+n\kappa^{2}}{4\kappa^{2}}\right)+4\exp(-Cm)$ (47)
@@
 $\left\|\mathcal{P}_{YV_{r}}YV_{\perp}\right\|^{2}\leq\frac{8C}{c}\,m\kappa^{2}\Bigg\}.$
-It follows from (B.1), (33) and (38) that by the union bound,
+It follows from (C), (42) and (47) that by the union bound,
 $\mathbb{P}(\mathcal{E}^{c})$ $\leq 6\exp\left(Cr-c\min\left\{\frac{1}{36\kappa^{2}}\,\frac{\left(\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X)\right)^{2}}{\sigma^{2}_{r}(X)+n\kappa^{2}},\,\frac{\sigma_{r}^{2}(X)-\sigma_{r+1}^{2}(X)}{6\kappa^{2}}\right\}\right)$ (48)
@@
 Suppose the conditions of Theorem 3 hold, in particular the condition in Equation (6) for each mode-$k$ unfolding $\mathcal{M}_{k}(X^{*})$ with target rank $r_{k}$.
-Let $U_{k}^{(0)}$ be the matrix of top $r_{k}$ left singular vectors of the mode-$k$ unfolding $Y^{(k)}:=\mathcal{M}_{k}(Y)$.
+Let $U_{k}^{(0)}$ be the matrix of top $r_{k}$ left singular vectors of the mode-$k$ unfolding $Y^{(k)}=\mathcal{M}_{k}(Y)$.
 Then for each $k\in\{1,2,3\}$, with probability at least $1-C_{1}\exp(-C_{2}p_{k})$,
@@
 Proof.
-Apply Lemma 13 to $Y^{(k)}=\mathcal{M}_{k}(Y)\in\mathbb{R}^{p_{k}\times p_{\neg k}}$ with signal $X^{(k)}=\mathcal{M}_{k}(X^{*})$ and noise $Z^{(k)}=\mathcal{M}_{k}(Z)$, where $p_{\neg k}:=\prod_{j\neq k}p_{j}$.
-Use rank $r=r_{k}$ and identify $n:=p_{k}$ (rows) and $m:=p_{\neg k}$ (columns).
+Apply Lemma 1 to $Y^{(k)}=\mathcal{M}_{k}(Y)\in\mathbb{R}^{p_{k}\times p_{\neq k}}$ with signal $X^{(k)}=\mathcal{M}_{k}(X^{*})$ and noise $Z^{(k)}=\mathcal{M}_{k}(Z)$, where $p_{\neq k}=\prod_{j\neq k}p_{j}$.
+Use rank $r=r_{k}$ and identify $n=p_{k}$ (rows) and $m=p_{\neq k}$ (columns).
 The lemma (applied to left singular vectors, or equivalently to the transpose) gives
-$\bigl\|\sin\Theta(U_{k}^{(0)},U_{k}^{*})\bigr\|^{2}\;\leq\;C\left\{\frac{p_{k}\,\kappa^{2}}{\Delta_{k}^{2}}+\frac{\kappa^{4}\,p_{k}\,p_{\neg k}}{\Delta_{k}^{4}}\right\},$
-where $\Delta_{k}:=\sigma_{r_{k}}\!\bigl(\mathcal{M}_{k}(X^{*})\bigr)-\sigma_{r_{k}+1}\!\bigl(\mathcal{M}_{k}(X^{*})\bigr)$ is the mode-$k$ spectral gap.
+$\bigl\|\sin\Theta(U_{k}^{(0)},U_{k}^{*})\bigr\|^{2}\;\leq\;C\left\{\frac{p_{k}\,\kappa^{2}}{\Delta_{k}^{2}}+\frac{\kappa^{4}\,p_{k}\,p_{\neq k}}{\Delta_{k}^{4}}\right\},$
+where $\Delta_{k}=\sigma_{r_{k}}\!\bigl(\mathcal{M}_{k}(X^{*})\bigr)-\sigma_{r_{k}+1}\!\bigl(\mathcal{M}_{k}(X^{*})\bigr)$ is the mode-$k$ spectral gap.
 By the assumption (6) in Theorem 3 (applied to mode $k$),
-$\Delta_{k}^{2}\;\geq\;C_{\mathrm{gap}}\kappa^{2}\!\left(\sqrt{p_{k}\,p_{\neg k}\,r_{\max}}+r_{\max}\sum_{j=1}^{3}p_{j}\right).$
+$\Delta_{k}^{2}\;\geq\;C_{\mathrm{gap}}\kappa^{2}\!\left(\sqrt{p_{k}\,p_{\neq k}\,r_{\max}}+r_{\max}\sum_{j=1}^{3}p_{j}\right).$
 Choosing $C_{\mathrm{gap}}$ sufficiently large makes the right–hand side above at most $1/(4r_{\max})$, hence
@@
 The probability bound $1-C_{1}\exp(-C_{2}p_{k})$ matches the row dimension in the matrix lemma. ∎
-Appendix C Matrix Perturbation Bounds
+Appendix D Auxiliary results
+D.1 Deviations bounds
+Lemma 2 (Theorem 4.4.3 in Vershynin (2018)).
+Assume all the entries of $Z\in\mathbb{R}^{m\times n}$ are independent mean-zero sub-Gaussian random variables, i.e.
+$\|Z_{ij}\|_{\psi_{2}}=\sup_{q\geq 1}\mathbb{E}(|Z_{ij}|^{q})^{1/q}/q^{1/2}\leq\kappa.$
+Then there exist some universal constant $C>0$, such that
+$\|Z\|\leq C\kappa\left(\sqrt{m}+\sqrt{n}\right)$
+with probability at least $1-\exp(-(m+n)))$.
+Lemma 3.
+Suppose all the entries of $Z\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}}$ are independent mean-zero sub-Gaussian random variables, i.e.
+$\|Z_{ijk}\|_{\psi_{2}}=\sup_{q\geq 1}\mathbb{E}(|Z_{ijk}|^{q})^{1/q}/q^{1/2}\leq\kappa.$
+Then there exist some universal constants $C,c>0$, such that
+$\sup_{\begin{subarray}{c}A\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}},\\ \|A\|_{\mathrm{F}}\leq 1,A\in\mathcal{T}_{(r_{1},r_{2},r_{3})}\end{subarray}}\langle Z,A\rangle\leq C\kappa\left(r_{1}r_{2}r_{3}+\sum_{k=1}^{3}p_{k}r_{k}\right)^{1/2}$
+with probability at least $1-\exp(-c\sum_{k=1}^{3}p_{k}r_{k})$.
+Proof.
+This directly follows from Lemma E.5 in Han et al. (2022). ∎
+Lemma 4.
+Suppose all the entries of $Z\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}}$ are independent mean-zero sub-Gaussian random variables, i.e.
+$\|Z_{ijk}\|_{\psi_{2}}=\sup_{q\geq 1}\mathbb{E}(|Z_{ijk}|^{q})^{1/q}/q^{1/2}\leq\kappa.$
+Let $W_{2}\in\mathbb{O}^{p_{2}\times s_{2}}$ and $W_{3}\in\mathbb{O}^{p_{3}\times s_{3}}$ be non-random. Then there exists absolute positive constants $C_{1},C_{2}$ and $c$ such that
+$\mathbb{P}\bigg(\big\|\mathcal{M}_{1}(Z)\,(W_{2}\otimes W_{3})\big\|\geq C_{1}\kappa(\sqrt{p_{1}+s_{2}s_{3}})\bigg)\leq C_{2}\exp\left(-cp_{1}\right),$
+Proof.
+It suffices to observe that $W_{2}\otimes W_{3}\in\mathbb{O}^{p_{2}p_{3}\times r_{2}r_{3}}$. The desired result is a direct consequence of Lemma 6. ∎
+Lemma 5.
+Suppose all the entries of $Z\in\mathbb{R}^{p_{1}\times p_{2}\times p_{3}}$ are independent mean-zero sub-Gaussian random variables, i.e.
+$\|Z_{ijk}\|_{\psi_{2}}=\sup_{q\geq 1}\mathbb{E}(|Z_{ijk}|^{q})^{1/q}/q^{1/2}\leq\kappa.$
+Then there exists absolute positive constants $C_{1},C_{2}$ and $c$ such that
+$\mathbb{P}\bigg(\sup_{\begin{subarray}{c}V_{2}\in\mathbb{R}^{p_{2}\times r_{2}},\,\|V_{2}\|\leq 1\\ V_{3}\in\mathbb{R}^{p_{3}\times r_{3}},\,\|V_{3}\|\leq 1\end{subarray}}\big\|\mathcal{M}_{1}(Z)\,(V_{2}\otimes V_{3})\big\|\geq C_{1}\kappa\Big(\sqrt{p_{1}+r_{2}r_{3}+p_{2}r_{2}+p_{3}r_{3}}\Bigg)$
+$\leq C_{2}\exp\left(-c(p_{1}+p_{2}+p_{3})\right),$
+Proof.
+It follows from the assumption that $\left(\mathcal{M}_{1}(Z)\right)_{i,j}\overset{i.i.d}{\sim}\mathrm{subGaussian}(0,\kappa^{2})$.
+Step 1. For fixed $V_{2}\in\mathbb{R}^{p_{2}\times r_{2}}$ with $\|V_{2}\|\leq 1$ and $V_{3}\in\mathbb{R}^{p_{3}\times r_{3}}$ with $\|V_{3}\|\leq 1$, it follows that
+$\left\|V_{2}\otimes V_{3}\right\|=\left\|V_{2}\right\|\left\|V_{3}\right\|\leq 1,$
+We upper bound $\|\mathcal{M}_{1}(Z)\cdot(V_{2}\otimes V_{3})\|$ for any fixed $V_{2}$ and $V_{3}$. Since,
+$\mathcal{M}(Z)\in\mathbb{R}^{p_{1}\times p_{2}p_{3}},\left(\mathcal{M}(Z)\right)_{i,j}\overset{i.i.d}{\sim}\mathrm{subGaussian}(0,\kappa^{2}),\quad\text{and}\quad\left\|V_{2}\otimes V_{3}\right\|=\left\|V_{2}\right\|\left\|V_{3}\right\|=1.$
+It follows from Lemma 6 that
+$\mathbb{P}\Big(\left\|\mathcal{M}_{1}(Z)\cdot(V_{2}\otimes V_{3})\right\|>x\Big)\leq 2\exp\left(C(p_{1}+r_{2}r_{3})-cx^{2}\kappa^{-2}\right).$
+Step 2. Let $\mathcal{N}_{p_{2},r_{2}}(\epsilon)$ denote an $\epsilon$ net of the set
+$\{A\in\mathbb{R}^{p_{2}\times r_{2}}:\|A\|\leq 1\}$
+with respect to the operator norm $\|\cdot\|$. Similarly, let Let $\mathcal{N}_{p_{3},r_{3}}(\epsilon)$ denote an $\epsilon$ net of the set
+$\{B\in\mathbb{R}^{p_{3}\times r_{3}}:\|B\|\leq 1\}$
+with respect to the operator norm $\|\cdot\|$. Denote the random quantity
+$\psi=\sup_{\begin{subarray}{c}V_{2}\in\mathbb{R}^{p_{2}\times r_{2}},\,\|V_{2}\|\leq 1\\ V_{3}\in\mathbb{R}^{p_{3}\times r_{3}},\,\|V_{3}\|\leq 1\end{subarray}}\big\|\mathcal{M}_{1}(Z)\,(V_{2}\otimes V_{3})\big\|.$
+For any given $V_{2}\in\mathbb{R}^{p_{2}\times r_{2}}$ and $V_{3}\in\mathbb{R}^{q}$ with $\|V_{2}\|\leq 1$ and $\|V_{3}\|\leq 1$, let $\widetilde{V}_{2}\in\mathcal{N}_{p_{2},r_{2}}(1/4)$ and $\widetilde{V}_{3}\in\mathcal{N}_{p_{3},r_{3}}(1/4)$ be such that
+$\|V_{2}-\widetilde{V}_{2}\|\leq 1/4\quad\text{and}\quad\|V_{3}-\widetilde{V}_{3}\|\leq 1/4.$
+Then
+$\|\mathcal{M}_{1}(Z)V_{2}\otimes V_{3}\|\leq\|\mathcal{M}_{1}(Z)(V_{2}-\widetilde{V}_{2})\otimes V_{3}\|+\|\mathcal{M}_{1}(Z)\widetilde{V}_{2}\otimes(V_{3}-\widetilde{V}_{3})\|+\|\mathcal{M}_{1}(Z)\widetilde{V}_{2}\otimes\widetilde{V}_{3}\|.$
+Note that $\|V_{2}-\widetilde{V}_{2}\|\leq 1/4$. So
+$\|\mathcal{M}_{1}(Z)(V_{2}-\widetilde{V}_{2})\otimes V_{3}\|=\frac{1}{4}\|\mathcal{M}_{1}(Z)\{4(V_{2}-\widetilde{V}_{2})\}\otimes V_{3}\|\leq\frac{\psi}{4}.$
+Similarly
+$\|\mathcal{M}_{1}(Z)V_{2}\otimes(V_{3}-\widetilde{V}_{3})\|\leq\frac{\psi}{4}.$
+In addition,
+$\|\mathcal{M}_{1}(Z)\widetilde{V}_{2}\otimes\widetilde{V}_{3}\|\leq\sup_{V_{2}\in\mathcal{N}_{p_{2},r_{2}}(1/4),b\in\mathcal{N}_{p_{3},r_{3}}(1/4)}\|\mathcal{M}_{1}(Z)V_{2}\otimes V_{3}\|.$
+So for any $V_{2}$ and $V_{3}$,
+$\|\mathcal{M}_{1}(Z)V_{2}\otimes V_{3}\|\leq\frac{1}{2}\psi+\sup_{V_{2}\in\mathcal{N}_{p_{2},r_{2}}(1/4),b\in\mathcal{N}_{p_{3},r_{3}}(1/4)}\|\mathcal{M}_{1}(Z)V_{2}\otimes V_{3}\|.$
+Taking sup over all $V_{2}\in\{A\in\mathbb{R}^{p_{2}\times r_{2}}:\|A\|\leq 1\}$ and $V_{3}\in\{B\in\mathbb{R}^{p_{3}\times r_{3}}:\|B\|\leq 1\}$, it follows that
+$\psi\leq\frac{1}{2}\psi+\sup_{V_{2}\in\mathcal{N}_{p_{2},r_{2}}(1/4),b\in\mathcal{N}_{p_{3},r_{3}}(1/4)}\|\mathcal{M}_{1}(Z)V_{2}\otimes V_{3}\|,$
+or simply
+$\psi\leq 2\sup_{V_{2}\in\mathcal{N}_{p_{2},r_{2}}(1/4),b\in\mathcal{N}_{p_{3},r_{3}}(1/4)}\|\mathcal{M}_{1}(Z)V_{2}\otimes V_{3}\|.$
+Step 3.
+By Proposition 8 in Pajor (1998), the cardinality of $\mathcal{N}_{p_{2},r_{2}}(\epsilon)$ is bounded $(\frac{C}{\epsilon})^{p_{2}r_{2}}$, and $\mathcal{N}_{p_{3},r_{3}}(\epsilon)$ is bounded $(\frac{C}{\epsilon})^{p_{3}r_{3}}$.
+Therefore
+$\mathbb{P}(\psi\geq 2t)\leq$ $\mathbb{P}\bigg(\sup_{V_{2}\in\mathcal{N}_{p_{2},r_{2}}(1/4),b\in\mathcal{N}_{p_{3},r_{3}}(1/4)}\|\mathcal{M}_{1}(Z)V_{2}\otimes V_{3}\|\geq t\bigg)$
+$\leq$ $C_{2}^{p_{2}r_{2}}C_{3}^{p_{3}r_{3}}\sup_{V_{2}\in\mathcal{N}_{p_{2},r_{2}}(1/4),b\in\mathcal{N}_{p_{3},r_{3}}(1/4)}\mathbb{P}\bigg(\|\mathcal{M}_{1}(Z)V_{2}\otimes V_{3}\|\geq t\bigg)$
+$\leq$ $2\exp\left(C(p_{1}+r_{2}r_{3}+p_{2}r_{2}+p_{3}r_{3})-ct^{2}\kappa^{-2}\right).$
+Here $C$ and $C_{3}$ are positive constants. The desired result follows by noting
+$\psi=\sup_{\begin{subarray}{c}V_{2}\in\mathbb{R}^{p_{2}\times r_{2}},\,\|V_{2}\|\leq 1\\ V_{3}\in\mathbb{R}^{p_{3}\times r_{3}},\,\|V_{3}\|\leq 1\end{subarray}}\big\|\mathcal{M}_{1}(Z)\,(V_{2}\otimes V_{3})\big\|.$
+Lemma 6.
+Suppose $Z\in\mathbb{R}^{n\times m}$, with $Z_{ij}\overset{i.i.d}{\sim}\mathrm{subGaussian}(0,\kappa^{2})$.
+Let $A\in\mathbb{R}^{p\times n}$ and $B\in\mathbb{R}^{m\times q}$ be non-random matrices.
+Then for any $t>0$
+$\mathbb{P}\left(\left\|A\,Z\,B\right\|>t\right)\leq C_{1}\,\exp\left(C(p+q)-\frac{c\,t^{2}}{\kappa^{2}\,\left\|A\right\|^{2}\,\left\|B\right\|^{2}}\right).$ (52)
+Proof.
+Step 1.
+Let $u\in\mathbb{R}^{n}$ and $v\in\mathbb{R}^{m}$ be non-random.
+Then that $u^{\top}\,Z\,v=\sum_{i=1}^{n}\sum_{j=1}^{n}u_{i}Z_{ij}v_{j}$.
+Since $Z_{ij}$ are i.i.d.
+sub-Gaussian with parameter $\kappa^{2}$, it follows that $u^{\top}\,Z\,v$ is sub-Gaussian with parameter $\kappa^{2}\|u\|_{2}^{2}\|v\|_{2}^{2}$.
+Consequently by Hoeffding’s inequality,
+$\mathbb{P}\left(\left|u^{\top}\,Z\,v\right|\,>t\,\right)\leq 2\,\exp\left(-\frac{c\,t^{2}}{\kappa^{2}\,\left\|u\right\|_{2}^{2}\,\left\|v\right\|_{2}^{2}}\right).$
+Step 2. Let $a\in\mathbb{R}^{p}$ and $b\in\mathbb{R}^{q}$ be non-random vectors such that $\|a\|_{2},\|b\|_{2}\leq 1$. By Step 1, it follows that
+$\mathbb{P}\left(\left|a^{\top}AZ\,Bb\right|\,>t\,\right)\leq 2\,\exp\left(-\frac{ct^{2}}{\kappa^{2}\,\left\|Aa\right\|_{2}^{2}\,\left\|Bb\right\|_{2}^{2}}\right)\leq 2\exp\left(-\frac{ct^{2}}{\kappa^{2}\,\left\|A\right\|^{2}\,\left\|B\right\|^{2}}\right).$
+Step 3.
+Let $\mathcal{N}_{p}(\epsilon)$ be an $\epsilon$-net of the unit ball in $\mathbb{R}^{p}$.
+It follows that for any $a\in\mathbb{R}^{p}$ with $\|a\|_{2}=1$, there exists $\widetilde{a}\in\mathcal{N}_{p}(\epsilon)$ such that
+$\|a-\widetilde{a}\|_{2}\leq\epsilon.$
+Similarly let $\mathcal{N}_{q}(\epsilon)$ be an $\epsilon$-net of the unit ball in $\mathbb{R}^{q}$.
+Denote the random quantity
+$\psi=\|AZB\|=\sup_{a\in\mathbb{R}^{p},b\in\mathbb{R}^{q}\|a\|_{2}=\|b\|_{2}=1}|a^{\top}AZBb|.$
+For any given $a\in\mathbb{R}^{p}$ and $b\in\mathbb{R}^{q}$, let $\widetilde{a}\in\mathcal{N}_{p}(1/4)$ and $\widetilde{b}\in\mathcal{N}_{q}(1/4)$ be such that
+$\|a-\widetilde{a}\|_{2}\leq 1/4\quad\text{and}\quad\|b-\widetilde{b}\|_{2}\leq 1/4.$
+Then
+[…]
+Note that $\|a-\widetilde{a}\|\leq\frac{1}{4}$. So
+$|(a-\widetilde{a})^{\top}AZBb|=\frac{1}{4}|\{4(a-\widetilde{a})^{\top}\}AZBb|\leq\frac{\psi}{4}.$
+Similarly
+$|\widetilde{a}^{\top}AZB(\widetilde{b}-b)|\leq\frac{\psi}{4}.$
+In addition,
+$|\widetilde{a}^{\top}AZB\widetilde{b}|\leq\sup_{a\in\mathcal{N}_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}|a^{\top}AZBb|.$
+So for any $a$ and $b$,
+$|a^{\top}AZBb|\leq\frac{1}{2}\psi+\sup_{a\in N_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}|a^{\top}AZBb|.$
+Taking sup over all unit vectors $a\in\mathbb{R}^{p}$ and $b\in\mathbb{R}^{q}$, it follows that
+$\psi\leq\frac{1}{2}\psi+\sup_{a\in\mathcal{N}_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}|a^{\top}AZBb|,$
+or simply
+$\psi\leq 2\sup_{a\in\mathcal{N}_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}|a^{\top}AZBb|.$
+Step 4.
+By Vershynin (2018), the cardinality of $\mathcal{N}_{p}(\epsilon)$ is bounded by $(\frac{C}{\epsilon})^{p}$, and the cardinality of $\mathcal{N}_{q}(\epsilon)$ is bounded by $(\frac{C}{\epsilon})^{q}$.
+Therefore
+$\mathbb{P}(\psi\geq 2t)=\mathbb{P}\bigg(\sup_{a\in\mathcal{N}_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}|a^{\top}AZBb|\geq t\bigg)\leq$ ${C_{2}}^{p}{C_{2}}^{q}\sup_{a\in\mathcal{N}_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}\mathbb{P}(|a^{\top}AZBb|\geq t)$
+$\leq$ $2{C_{2}}^{p}{C_{2}}^{q}\exp\left(-\frac{ct^{2}}{\kappa^{2}\,\left\|A\right\|^{2}\,\left\|B\right\|^{2}}\right),$
+where $c,C$ are positive constants. The desired result follows by noting $\psi=\|AZB\|$. ∎
+Lemma 7.
+Suppose $Z\in\mathbb{R}^{n\times m}$, with $Z_{ij}\overset{i.i.d}{\sim}\mathrm{subGaussian}(0,\kappa^{2})$.
+Let $A\in\mathbb{R}^{m\times p}$ and $B\in\mathbb{R}^{m\times q}$ be non-random matrices.
+Then for any $t>0$
+$\mathbb{P}\left(\left\|A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B\right\|>t\right)\leq C_{1}\,\exp\left(C(p+q)-\min\left(\frac{t^{2}}{n\,\kappa^{4}\,\left\|B\right\|^{2}\,\left\|A\right\|^{2}},\,\frac{t}{\kappa^{2}\left\|B\right\|\,\left\|A\right\|}\right)\right),$ (53)
+where $C$ and $C_{1}$ are positive constants.
+Proof.
+For any non-random $u,v\in\mathbb{R}^{m}$ , it follows that
+$u^{\top}Z^{\top}Zv-n\kappa^{2}u^{\top}v=\sum_{j=1}^{n}(u^{\top}Z_{j})(v^{\top}Z_{j})-\mathbb{E}\{(u^{\top}Z_{j})(v^{\top}Z_{j})\},$
+where $Z_{j}$ is the $j$-th row of $Z$.
+Note that $(u^{\top}Z_{j})$ is sub-Gaussian with parameter $\kappa^{2}\|u\|_{2}^{2}$, and $(v^{\top}Z_{j})$ is sub-Gaussian with parameter $\kappa^{2}\|v\|_{2}^{2}$.
+Since $Z$ have i.i.d.
+entries, it follows that $\{(u^{\top}Z_{j})(v^{\top}Z_{j})\}_{j=1}^{n}$ are i.i.d.
+sub-exponential with parameter $\kappa^{4}\|u\|_{2}^{2}\|v\|_{2}^{2}$.
+So
+$\mathbb{P}\bigg(|u^{\top}Z^{\top}Zv-n\kappa^{2}u^{\top}v|\geq t\bigg)\leq 2\exp\bigg(-c\min\bigg\{\frac{t^{2}}{n\kappa^{4}\|u\|_{2}^{2}\|v\|_{2}^{2}},\frac{t}{\kappa^{2}\|u\|_{2}\|v\|_{2}}\bigg\}\bigg).$
+Step 1.
+Let $\mathcal{N}_{p}(\epsilon)$ be the $\epsilon$-net of the unit ball in $\mathbb{R}^{p}$.
+It follows that for any $a\in\mathbb{R}^{p}$ with $\|a\|_{2}=1$, there exists $\widetilde{a}\in\mathcal{N}_{p}(\epsilon)$ such that
+$\|a-\widetilde{a}\|_{2}\leq\epsilon.$
+Similarly let $\mathcal{N}_{q}(\epsilon)$ be the $\epsilon$-net of the unit ball in $\mathbb{R}^{q}$.
+Denote the random quantity
+$\psi=\|A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B\|=\sup_{a\in\mathbb{R}^{p},b\in\mathbb{R}^{q}\|a\|_{2}=\|b\|_{2}=1}\bigl|a^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b\bigr|.$
+For any given $a\in\mathbb{R}^{p}$ and $b\in\mathbb{R}^{q}$, let $\widetilde{a}\in\mathcal{N}_{p}(1/4)$ and $\widetilde{b}\in\mathcal{N}_{q}(1/4)$ be such that
+$\|a-\widetilde{a}\|_{2}\leq 1/4\quad\text{and}\quad\|b-\widetilde{b}\|_{2}\leq 1/4.$
+Then
+$|a^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b|\leq|(a-\widetilde{a})^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b|$
+[…]
+Note that $\|a-\widetilde{a}\|_{2}\leq\frac{1}{4}$. So
+$|(a-\widetilde{a})^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b|=\frac{1}{4}|\{4(a-\widetilde{a})^{\top}\}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b|\leq\frac{\psi}{4}.$
+Similarly
+$|\widetilde{a}^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)(\widetilde{b}-b)|\leq\frac{\psi}{4}.$
+In addition,
+$|\widetilde{a}^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)\widetilde{b}|\leq\sup_{a\in\mathcal{N}_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}|a^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b|.$
+So for any $a$ and $b$,
+$|a^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b|\leq\frac{1}{2}\psi+\sup_{a\in N_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}|a^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b|.$
+Taking sup over all unit vectors $a\in\mathbb{R}^{p}$ and $b\in\mathbb{R}^{q}$, it follows that
+$\psi\leq\frac{1}{2}\psi+\sup_{a\in\mathcal{N}_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}|a^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b|,$
+or simply
+$\psi\leq 2\sup_{a\in\mathcal{N}_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}|a^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b|.$
+Step 2.
+By Vershynin (2018), the cardinality of $\mathcal{N}_{p}(\epsilon)$ is bounded by $(\frac{C}{\epsilon})^{p}$, and the cardinality of $\mathcal{N}_{q}(\epsilon)$ is bounded by $(\frac{C}{\epsilon})^{q}$, for a positive constant $C$.
+Therefore
+$\mathbb{P}(\psi\geq 2t)=$ $\mathbb{P}\bigg(\sup_{a\in\mathcal{N}_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}|a^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b|\geq t\bigg)$
+$\leq$ ${C_{2}}^{p}{C_{2}}^{q}\sup_{a\in\mathcal{N}_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}\mathbb{P}(|a^{\top}(A^{\top}Z^{\top}ZB-n\kappa^{2}A^{\top}B)b|\geq t)$
+$=$ ${C_{2}}^{p}{C_{2}}^{q}\sup_{a\in\mathcal{N}_{p}(1/4),b\in\mathcal{N}_{q}(1/4)}\mathbb{P}(|(Aa)^{\top}Z^{\top}Z(Bb)-n\kappa^{2}(Aa)^{\top}(Bb)|\geq t)$
+$\leq$ $2{C_{2}}^{p}{C_{2}}^{q}\exp\bigg(-c\min\bigg\{\frac{t^{2}}{n\kappa^{4}\|Aa\|_{2}^{2}\|Bb\|_{2}^{2}},\frac{t}{\kappa^{2}\|Aa\|_{2}\|Bb\|_{2}}\bigg\}\bigg),$
+where $C_{2}$ is a positive constant. The desired result follows from the observation that $\|Aa\|_{2}\leq\|A\|\|a\|_{2}\leq\|A\|$, and $\|Bb\|_{2}\leq\|B\|\|b\|_{2}\leq\|B\|$. ∎
+Lemma 8.
+Suppose $Z\in\mathbb{R}^{m\times n}$ is a sub-Gaussian random matrix in the sense that for any $u\in\mathbb{R}^{m},v\in\mathbb{R}^{n}$, it holds that
+$\|u^{\top}Zv\|_{\psi_{2}}\leq\kappa\|u\|_{2}\|v\|_{2}.$
+Then with probability at least $1-\exp(-c(m+n))$, it holds that
+$\mathbb{P}\left(\left\|Z\right\|>t\right)\leq C_{1}\,\exp\left(C(m+n)-\frac{c\,t^{2}}{\kappa^{2}\,}\right),$
+where $c,C$ and $C_{1}$ are positive constants.
+Proof.
+By assumption,
+$\mathbb{P}\left(\left|u^{\top}\,Z\,v\right|\,>t\,\right)\leq 2\,\exp\left(-\frac{c\,t^{2}}{\kappa^{2}\,\left\|u\right\|_{2}^{2}\,\left\|v\right\|_{2}^{2}}\right).$
+The rest of the proof is similar and simpler than Lemma 6 and is omitted.
+D.2 Matrix perturbation bounds
 Lemma 9.
@@
 $=$ $3\left\|B_{(R)}-B\right\|_{\mathrm{F}}+2\min\left\{\sqrt{R}\|Z\|,\|Z\|_{\mathrm{F}}\right\},$
-where $B_{(R)}$ denote the rank-$R$ truncated SVD of $B$, this is, if $B=U\Sigma V^{\top}$ then $B_{(R)}:=U_{(R)}\Sigma_{(R)}V_{(R)}^{\top}$.
+where $B_{(R)}$ denote the rank-$R$ truncated SVD of $B$, this is, if $B=U\Sigma V^{\top}$ then $B_{(R)}=U_{(R)}\Sigma_{(R)}V_{(R)}^{\top}$.
 Proof.
@@
 $\leq$ $\min\left\{\sqrt{R}\|Z\|,\|Z\|_{\mathrm{F}}\right\}+\sqrt{\sum_{j=R+1}^{n}\sigma_{j}^{2}(B)},$ (56)
-where the first inequality follows from the triangle inequality, and second inequality follows from Weyl’s inequality (Weyl,, 1912), i.e.
+where the first inequality follows from the triangle inequality, and second inequality follows from Weyl’s inequality (Weyl, 1912), i.e.
 $|\sigma_{j}(A)-\sigma_{j}(B)|\leq\|A-B\|$ for all $1\leq j\leq n$, as well as the fact that
 $\left\|(\sigma_{R+1}(A)-\sigma_{R+1}(B),\dots,\sigma_{2R}(A)-\sigma_{2R}(B))^{\top}\right\|\leq\left\|\Sigma(A)-\Sigma(B)\right\|_{\mathrm{F}}\leq\left\|Z\right\|_{\mathrm{F}},$
-where the last inequality follows from Theorem 23. It then follows from (C),
+where the last inequality follows from Theorem 5. It then follows from (D.2),
 $\sqrt{\sum_{j=1}^{R}\sigma_{j}^{2}(\mathcal{P}_{\widehat{U}_{\perp}}B)}=\left\|(\sigma_{1}(\mathcal{P}_{\widehat{U}_{\perp}}(A-Z)),\dots,\sigma_{R}(\mathcal{P}_{\widehat{U}_{\perp}}(A-Z))^{\top}\right\|$
@@
 $\leq$ $2\min\left\{\sqrt{R}\|Z\|,\|Z\|_{\mathrm{F}}\right\}+\sqrt{\sum_{j=R+1}^{n}\sigma_{j}^{2}(B)},$
-where the first two inequalities follow from the same arguments as in (C). Consequently,
+where the first two inequalities follow from the same arguments as in (D.2). Consequently,
 $\left\|\mathcal{P}_{\widehat{U}_{\perp}}B\right\|_{\mathrm{F}}\leq 3\sqrt{\sum_{j=R+1}^{n}\sigma_{j}^{2}(B)}+2\min\left\{\sqrt{R}\|Z\|,\|Z\|_{\mathrm{F}}\right\}.$
-Lemma 21 (Proposition 1 of Cai and Zhang, (2018)).
+Lemma 15 (Proposition 1 of Cai and Zhang (2018)).
 Suppose $Y\in\mathbb{R}^{m\times n}$, $\widehat{V}=[\widehat{V}_{r}\ \widehat{V}_{\perp}]\in\mathbb{O}_{n}$ where $\widehat{V}_{r}\in\mathbb{O}_{n,r}$, $\widehat{V}_{\perp}\in\mathbb{O}_{n,n-r}$ correspond to the first $r$ and last (nr)(n-r) right singular vectors of $Y$ respectively.
@@
     $\|\widehat{V}\widehat{V}^{\top}-VV^{\top}\|_{\mathrm{F}}=\sqrt{2}\|\sin\Theta(\widehat{V},V)\|_{\mathrm{F}}.$
-Theorem 23 (Mirsky’s singular value inequality in Mirsky, (1960)).
+Theorem 5 (Mirsky’s singular value inequality in Mirsky (1960)).
 For any matrices $A,B\in\mathbb{R}^{m\times n}$, let $A=V_{1}\Sigma(A)W_{1}^{\top}$ and $B=V_{2}\Sigma(B)W_{2}^{\top}$ be the full SVDs of $A$ and $B$, respectively.
@@
 $\sigma_{i+j-1}(A+B)\leq\sigma_{i}(A)+\sigma_{j}(B).$
-Lemma 25 (Ky Fan-type Inequality for Sums of Matrices).
-Let $A,B\in\mathbb{R}^{m\times n}$ and denote their singular values (in nonincreasing order) by $\{\sigma_{i}(A)\}$ and $\{\sigma_{i}(B)\}$ respectively.
-In addition denote the singular values of $A+B$ as $\{\sigma_{i}(A+B)\}$.
-Then for any $1\leq k\leq\min\{m,n\}$, it holds that
-$\sqrt{\sum_{i=1}^{k}\sigma_{i}^{2}(A+B)}\leq\sqrt{\sum_{i=1}^{k}\sigma_{i}^{2}(A)}+\sqrt{\sum_{i=1}^{k}\sigma_{i}^{2}(B)}.$
-Proof.
-For a symmetric matrix $M\in\mathbb{R}^{n\times n}$, by Ky Fan’s maximum principle (see e.g. II.1.13 in Bhatia, (2013)), for any $1\leq k\leq n$,
-$\sum_{i=1}^{k}\lambda_{i}(M)=\sup_{P\in\mathbb{O}^{n\times k}}\operatorname*{tr}(P^{\top}MP).$
-Therefore
-$\sum_{i=1}^{k}\sigma_{i}^{2}(A)=\sum_{i=1}^{k}\lambda_{i}(A^{\top}A)=\sup_{P\in\mathbb{O}^{n\times k}}\operatorname*{tr}(P^{\top}A^{\top}AP)=\sup_{P\in\mathbb{O}^{n\times k}}\|AP\|_{\mathrm{F}}^{2},$
-and so
-$\sqrt{\sum_{i=1}^{k}\sigma_{i}^{2}(A)}=\sup_{P\in\mathbb{O}^{n\times k}}\|AP\|_{\mathrm{F}}.$
-Then
-$\sqrt{\sum_{i=1}^{k}\sigma_{i}^{2}\left(A+B\right)}=$ $\max_{U\in\mathbb{O}_{n,k}}\|(A+B)U\|_{\mathrm{F}}\leq\max_{U\in\mathbb{O}_{n,k}}\|AU\|_{\mathrm{F}}+\max_{U\in\mathbb{O}_{n,k}}\|BU\|_{\mathrm{F}}$
-$=\sqrt{\sum_{i=1}^{k}\sigma_{i}^{2}\left(A\right)}+\sqrt{\sum_{i=1}^{k}\sigma_{i}^{2}\left(B\right)}.$
 Let $\mathcal{T}_{(r_{1},r_{2},r_{3})}$ denote the class of tensor in $\mathbb{R}^{p_{1}\times p_{2}\times p_{3}}$ with tucker ranks at most $(r_{1},r_{2},r_{3})$. More precisely
@@
 $\sqrt{\sum_{j=r_{1}+1}^{\operatorname*{rank}(\mathcal{M}_{1}(X^{*}))}\sigma^{2}_{j}(\mathcal{M}_{1}(X^{*}))}\leq\xi_{(r_{1},r_{2},r_{3})}.$
+Lemma 18 (Ky Fan-type Inequality for Sums of Matrices).
+Let $A,B\in\mathbb{R}^{m\times n}$ and denote their singular values (in nonincreasing order) by $\{\sigma_{i}(A)\}$ and $\{\sigma_{i}(B)\}$ respectively.
+In addition denote the singular values of $A+B$ as $\{\sigma_{i}(A+B)\}$.
+Then for any $1\leq k\leq\min\{m,n\}$, it holds that
+$\sqrt{\sum_{i=1}^{k}\sigma_{i}^{2}(A+B)}\leq\sqrt{\sum_{i=1}^{k}\sigma_{i}^{2}(A)}+\sqrt{\sum_{i=1}^{k}\sigma_{i}^{2}(B)}.$
+Proof.
+For a symmetric matrix $M\in\mathbb{R}^{n\times n}$, by Ky Fan’s maximum principle (see e.g. II.1.13 in Bhatia (2013)), for any $1\leq k\leq n$,
+$\sum_{i=1}^{k}\lambda_{i}(M)=\sup_{P\in\mathbb{O}^{n\times k}}\operatorname*{tr}(P^{\top}MP).$
+Therefore
+$\sum_{i=1}^{k}\sigma_{i}^{2}(A)=\sum_{i=1}^{k}\lambda_{i}(A^{\top}A)=\sup_{P\in\mathbb{O}^{n\times k}}\operatorname*{tr}(P^{\top}A^{\top}AP)=\sup_{P\in\mathbb{O}^{n\times k}}\|AP\|_{\mathrm{F}}^{2},$
+and so
+$\sqrt{\sum_{i=1}^{k}\sigma_{i}^{2}(A)}=\sup_{P\in\mathbb{O}^{n\times k}}\|AP\|_{\mathrm{F}}.$
+Then
+$\sqrt{\sum_{i=1}^{k}\sigma_{i}^{2}\left(A+B\right)}=$ $\max_{U\in\mathbb{O}_{n,k}}\|(A+B)U\|_{\mathrm{F}}\leq\max_{U\in\mathbb{O}_{n,k}}\|AU\|_{\mathrm{F}}+\max_{U\in\mathbb{O}_{n,k}}\|BU\|_{\mathrm{F}}$
+$=\sqrt{\sum_{i=1}^{k}\sigma_{i}^{2}\left(A\right)}+\sqrt{\sum_{i=1}^{k}\sigma_{i}^{2}\left(B\right)}.$
+D.3 Minimax lower bound for vectors
+Lemma 19.
+Let $\kappa>0$ and $d\in\mathbb{Z}^{+}$, and denote
+$\mathcal{V}=\{V\in\mathbb{R}^{d},\|V\|_{\infty}\leq\kappa\}.$
+Consider the model $Y\sim\mathcal{N}(V,\kappa^{2}I_{d})$ where $V\in\mathcal{V}$. Then there exists a universal constant $c$ such that
+$\inf_{\widehat{V}}\sup_{V\in\mathcal{V}}\mathbb{E}\|\widehat{V}-V\|^{2}\geq c\kappa^{2}d,$
+where infimum is taken over all the estimators $^{\widehat{V}}$ based on $Y$.
+Proof.
+This is a standard minimax lower bound for bounded normal means; see for example Berry (1990) and Donoho et al. (1990). ∎